### v1

# Hybrid Monte Carlo for Failure Probability Estimation with Gaussian Process Surrogates

Abstract
We tackle the problem of quantifying failure probabilities for expensive computer experiments with stochastic inputs. The computational cost of evaluating the computer simulation prohibits direct Monte Carlo (MC) and necessitates a statistical surrogate model. Surrogate-informed importance sampling – which leverages the surrogate to identify suspected failures, fits a bias distribution to these locations, then calculates failure probabilities using a weighted average – is popular, but it is data hungry and can provide erroneous results when budgets are limited. Instead, we propose a hybrid MC scheme which first uses the uncertainty quantification (UQ) of a Gaussian process (GP) surrogate to identify areas of high classification uncertainty, then combines surrogate predictions in certain regions with true simulator evaluation in uncertain regions. We also develop a stopping criterion which informs the allocation of a fixed budget of simulator evaluations between surrogate training and failure probability estimation. Our method is agnostic to surrogate choice (as long as UQ is provided); we showcase functionality with both GPs and deep GPs. It is also agnostic to design choices; we deploy contour locating sequential designs throughout. With these tools, we are able to effectively estimate small failure probabilities with only hundreds of simulator evaluations. We validate our method on a variety of synthetic benchmarks before deploying it on an expensive computer experiment of fluid flow around an airfoil.
Keywords: computer experiment, deep Gaussian process, entropy, emulator, contour location, reliability, importance sampling
1 Introduction
Quantifying the probability of failure in complex dynamical systems is of great interest but is challenging due to the computational expense of the so-called “computer experiment.” Consider a black-box computer simulation, $f:\mathcal{X}\rightarrow\mathbb{R}$ with inputs $\mathcal{X}\subset\mathbb{R}^{d}$ governed by known measure $\mathbb{P}_{\mathbf{x}}$ (with density $p(\mathbf{x})$). The response is thresholded such that $f(\mathbf{x})>t$ indicates failure. The research objective is to estimate the failure probability,
[…]
We are particularly motivated by aeronautic applications where system failures may incur immense financial or environmental costs, or even loss of life. In aircraft design and development, computer simulation experiments are a natural, often mandatory, precursor to direct experimentation and deployment. For example, $f(\cdot)$ may represent the vibrations experienced by an aircraft which become unsafe above certain levels, or the fuel efficiency of an aircraft, which if too low, can be detrimental to the environment, obliterate fuel budgets, or overly restrict travel distances. Inputs $\mathcal{X}$ may encapsulate design parameters and operating conditions (such as wing shapes, aircraft design, flight speeds, and angles of attack). Yet we acknowledge a slew of other relevant applications where quantifying failure probabilites for complex simulations is critical, including structural reliability analysis using finite element simulations (e.g., Mia et al.,, 2017), Boltzmann modeling of chemical processes in fuel filtration (e.g., Belot et al.,, 2021), and thermodynamic flows for nuclear reactors (e.g., Batet et al.,, 2014).
A rudimentary approach to solve Eq. (1) involves brute force Monte Carlo (MC); simply sample an exhaustively large set of inputs, evaluate the computer simulation at each of these configurations, and report the proportion that resulted in failure. But this approach is prohibitively expensive if the computer simulation is costly (incurring long wall clock times), the probability of failure is rare (requiring enourmous sample sizes), or both! In this work, we propose a novel method for quantifying failure probabilities of complex computer simulations when the computation budget restricts evaluations of the simulator to several hundred, rendering traditional MC impractical.
In the face of expensive data, statistical models (called “surrogates” or “emulators”) are essential. Surrogates are trained on a limited set of simulator evaluations in order to provide predictions with appropriate uncertainty quantification (UQ) at unobserved inputs. Gaussian processes (GPs) are the canonical choice (Santner et al.,, 2003; Gramacy,, 2020), but deep Gaussian processes (DGPs; Damianou and Lawrence,, 2013) are becoming increasingly popular, particularly for non-stationary computer experiments (Sauer,, 2023). Notably, our contribution is agnostic to surrogate choice as long as effective UQ is provided, and is thus applicable to traditional GPs and Bayesian DGPs (Sauer et al.,, 2023b) and, potentially, Bayesian neural networks (MacKay,, 1995).
Early methods of failure probability estimation, namely the first and second order reliability methods (FORM/SORM; Haldar and Mahadevan,, 1995), leveraged crude linear approximations of the complex model and placed all failure probability estimation in the hands of the “most probable point” on the predicted contour. Advancements on FORM and SORM have been made with more flexible GP models (e.g., Su et al.,, 2017). A variety of works have instead leveraged GP surrogates directly with MC sampling, using the GP predictions (which are of neglible cost) in place of true simulator evaluation (e.g., Cheng and Lu,, 2020; Li et al.,, 2021; Lu et al.,, 2023). We will use this “surrogate MC” method as a benchmark in our later exercises. While it is cheap, it places too much stock in the accuracy of the surrogate and funnels surrogate inaccuracies directly into the failure probability estimate.
Importance sampling (IS; Srinivasan,, 2002; Tokdar and Kass,, 2010) offers an unbiased alternative. It hinges on the construction of a bias distribution that intentionally focuses density in the failure region. Inputs are sampled from the bias distribution (hopefully yielding many failures) and are then utilized in a weighted average calculation that corrects for the disparity between the bias density and the original input density. Identifying an effective bias distribution in the face of limited training data is a hefty task (Tabandeh et al.,, 2022). Adaptive methods abound (e.g., Oh and Berger,, 1992; Kurtz and Song,, 2013; Dalbey and Swiler,, 2014; Bugallo et al.,, 2017; Cheng et al.,, 2023). Peherstorfer et al., (2016) propose what they termed “multifidelity importance sampling,” although we prefer the term “surrogate-informed importance sampling” (SIIS) to avoid confusion with multifidelity surrogate modeling (e.g., Park et al.,, 2017). In SIIS, a surrogate is used to identify predicted failures, and a bias distribution is fit to these locations (more on this in Section 2.2). SIIS has been deployed with GP surrogates in a variety of subsequent works (e.g., Cole et al.,, 2023; Renganathan et al.,, 2023; Booth et al.,, 2024a; Renganathan,, 2024), and we view it as our primary competitor. Yet SIIS has several crucial drawbacks. We will delay a thorough discussion to Section 2.2, but the punchline is simply that IS has too much variability when bias distributions are suboptimal and data sizes are limited. The difficulty in training an effective bias distribution can render SIIS estimated failure probabilities ineffective, forming the motivation behind our current contribution.
We argue that utility and precision in practice are more important than theoretical unbiasedness. To this end, we embrace a hybrid Monte Carlo calculation that uses both surrogate predictions and true simulator evaluations in strategically partitioned regions of the input space. Li and Xiu, (2010) first proposed such a hybrid calculation, but their work was restricted to polynomial chaos (PC) models with strict accuracy assumptions in order to guarantee convergence. GP-based surrogates afford several key advancements over PCs. For one, GPs are flexible and data-driven; their accuracy continuously improves as training data sizes increase. Perhaps more importantly, GPs offer UQ through posterior variances. We leverage this UQ to prioritize expensive model evaluation at locations where the surrogate is least certain in its pass/fail prediction.
Regardless of how a surrogate is to be used (surrogate MC, SIIS, our hybrid MC, etc.), surrogate accuracy in identifying failure regions is of primary importance. Not all training designs are created equally. Space-filling designs (Joseph,, 2016) may miss failure regions altogether, particularly when training budgets are limited. Instead, it is common to train a surrogate on a small initial design then pair it with a relevant acquisition function to select subsequent inputs for evaluation through the simulator. This iterative process of acquisition and surrogate updating can strategically hone in on the contour, all while minimizing evaluations of the expensive simulator. So-called “contour location” (CL) naturally outperforms space-filling counterparts. Acquisition functions seek to balance exploitation and exploration by leveraging both predicted distance from the contour and posterior uncertainty. Popular variations fall under expected improvement frameworks (Bichon et al.,, 2008; Ranjan et al.,, 2008), stepwise uncertainty reduction frameworks (Bect et al.,, 2012; Chevalier et al.,, 2014; Azzimonti et al.,, 2021), or both (Duhamel et al.,, 2023). Others utilize pass/fail predictive entropy (Marques et al.,, 2018; Cole et al.,, 2023). Booth et al., (2024b) recently extended CL sequential designs to DGP surrogates. We will utilize state-of-the-art sequentially designed GP and DGP surrogates in our benchmark exercises, but our contribution is agnostic to these choices. We see our method’s “plug-and-play” nature as a distinct advantage.
Both surrogate training and failure probability estimation require a budget of expensive simulator evaluations. These two tasks are typically conducted in independent stages, with separate pre-determined budgets (as we will use in the first round of our synthetic benchmarks). Yet, in practice, it is common for a total budget to be set according to computational allowances, with no mention of how many evaluations to spend on the training phase versus the estimation phase. Some works have investigated stopping criteria for surrogate training in reliability settings (e.g., Echard et al.,, 2011; Hristov et al.,, 2019), but this is still an under-examined area of the literature. In our hybrid MC set-up, we have a unique advantage: the more we train our surrogate, the more certain it will be, requiring fewer high-cost evaluations for effective failure probability estimation. There is a clear give-and-take between the budgets of the training and estimation stages; increasing the budget of one decreases the budget required for the other, and vice versa. Leveraging this trade-off, we propose a method for allocating a fixed budget between surrogate training and hybrid Monte Carlo estimation.
This manuscript is laid out as follows. First, in Section 2, we review GP/DGP surrogates and competing methods of failure probability estimation. In Section 3 we present our proposed hybrid Monte Carlo procedure, including implementation details and a variety of synthetic benchmarking exercises. In Section 4 we propose our budget allocation scheme, continuing the established benchmarks. Finally, we deploy our method on a computer experiment of an RAE-8288 airfoil in Section 5 and conclude in Section 6.
2 Failure probability estimation
Denote single inputs as the row vector $\mathbf{x}$ of size $1\times d$ with response $y=f(\mathbf{x})$. Let $X_{n}$ represent the row-combined matrix of $n$-many inputs with corresponding response vector $\mathbf{y}_{n}$. Our objective is to quantify the failure probability from Eq. (1), namely $\mathrm{P}(f(\mathbf{x})>t)$ for $\mathbf{x}\sim p(\mathbf{x})$. Typical Monte Carlo with a budget of size $M$ provides the estimator
[…]
This estimator is ineffective unless a sufficient number of failures are observed, necessitating very large $M$ particularly for small $\alpha$. But our computer model $f(\mathbf{x})$ is expensive; we simply cannot evaluate it thousands/millions/billions of times. We will describe our sample efficient method for estimating failure probabilities in Section 3, after a quick review of preliminaries.
2.1 Gaussian process surrogates
A surrogate model $\hat{f}$, trained on a limited budget of runs, may provide predictions at unobserved inputs in place of the true $f$. To enable our proposed methodology the surrogate must also provide posterior predictive distributions. Gaussian processes (GPs) are the predominant choice; they assume a multivariate normal prior distribution over the response, $y\sim\mathcal{N}(\mu,\Sigma(X))$. The prior mean $\mu$ is often simplified to the zero vector (which we will use moving forward), and the prior covariance $\Sigma(X)$ is typically a function of inverse euclidean distances, e.g. $\Sigma(X)^{ij}=\Sigma(\mathbf{x}_{i},\mathbf{x}_{j})=k(||\mathbf{x}_{i}-\mathbf{x}_{j}||^{2})$ with kernel $k(\cdot)$. See Santner et al., (2003); Rasmussen et al., (2006); Gramacy, (2020) for thorough reviews. Conditioned on training data $\left\{X_{n},\mathbf{y}_{n}\right\}$, posterior predictions at singeton input $\mathbf{x}$ follow
$\hat{f}_{n}(\mathbf{x})\sim\mathcal{N}\left(\mu_{n}(\mathbf{x}),\sigma_{n}^{2}(\mathbf{x})\right)\quad\textrm{for}\quad\begin{array}[]{rl}\mu_{n}(\mathbf{x})&=\Sigma(\mathbf{x},X_{n})\Sigma(X_{n})^{-1}\mathbf{y}_{n}\\ \sigma_{n}^{2}(\mathbf{x})&=\Sigma(\mathbf{x})-\Sigma(\mathbf{x},X_{n})\Sigma(X_{n})^{-1}\Sigma(X_{n},\mathbf{x}),\end{array}$ (3)
where $\Sigma(\mathbf{x},X_{n})$ denotes the vector formed from applying the kernel between $\mathbf{x}$ and every row of $X_{n}$. Joint predictions are straightforward but not necessary for our contribution.
A natural alternative to true MC (Eq. 2) is “surrogate MC” which leverages surrogate predictions in place of the true simulator,
$\hat{\alpha}_{\textrm{SURR}}=\frac{1}{M}\sum_{i=1}^{M}\mathbb{I}_{\left\{\mu_{n}(\mathbf{x}_{i})>t\right\}}\quad\textrm{for}\quad\mathbf{x}_{i}\sim p(\mathbf{x}).$ (4)
If the surrogate is not accurate near the contour, it will misclassify inputs which will severely hinder the effectiveness of this estimator.
Figure 1: Left/Center: Herbie function with failure contour (red) and input distribution (black). Right: Entropy surface (yellow/high, purple/low) for a GP trained on a 100 point LHS (white circles).
Figure 2: Slice of Herbie function along $x_{2}=1.1$. GP was trained on 100 point LHS in 2 dimensions.
For example, consider the two-dimensional “Herbie” function (Lee et al.,, 2011) shown in the left and center panels of Figure 1 (we will revisit the right panel momentarily). We define a failure threshold of $t=1.065$ which demarcates four disjoint failure regions (red contours). We also define $p(\mathbf{x})$ using independent normal distributions centered at the origin (solid black contours). [^1] Full details of all simulated functions, thresholds, and distributions are provided in the Supplementary Material. We trained a GP surrogate using a random Latin hypercube sample (LHS; McKay et al.,, 2000) of size $n=100$ in the two dimensions. Figure 2 shows the GP posterior prediction along the slice $x_{2}=1.1$. The posterior mean (blue dashed) and 95% CI (blue shading) capture the true nonlinear surface (solid black) fairly well. Even so, there are inputs (highlighted by the yellow shading) which the surrogate would misclassify as “passes” instead of “failures.” Thus any estimate reliant solely on this surrogate would underestimate the failure probability.
Deep Gaussian processes.
In some situations, surrogate predictions may be improved through adaptations to the surrogate model itself. Traditional GPs are limited by the commonly assumed stationarity of the covariance kernel; they are not able to accurately model non-stationary surfaces characterized by regime shifts and stark transitions. Deep Gaussian processes (Damianou and Lawrence,, 2013) are a promising more flexible alternative. They leverage functional compositions of GP layers in which latent layers act as warped versions of the original inputs, allowing for greater flexibility. DGPs have consistently outperformed stationary GPs in modeling non-stationary computer experiments (e.g., Rajaram et al.,, 2021; Yazdi,, 2022; Ming et al.,, 2023), particularly when deployed in a Bayesian framework with full UQ (e.g., Sauer et al.,, 2023b, a). See Sauer, (2023) for a thorough review of DGP surrogates. Crucially, our method is agnostic to surrogate choice as long as UQ is provided. We will demonstrate functionality with both traditional GP and Bayesian DGP surrogates in Section 3.2.
Contour location.
To facilitate failure probability estimation, a surrogate’s sole task it to accurately classify passes and failures. Strategic sequential designs targeting the failure contour, i.e., “contour location,” will outperform space-filling counterparts in this regard. In CL, after initializing with a small space-filling design, subsequent training locations are selected iteratively through the optimization of an acquisition criterion. The computer simulation is evaluated at the selected input(s), the surrogate is updated, and the process is repeated until the budget is exhausted or a stopping criterion is met.
Identifying acquisition criteria which effectively target failure contours is a hot topic, particularly in conjunction with GP surrogates, (e.g., Bichon et al.,, 2008; Ranjan et al.,, 2008; Bect et al.,, 2012; Chevalier et al.,, 2014; Marques et al.,, 2018; Cole et al.,, 2023). Here, we highlight one in particular – classification entropy – defined as
$H(\mathbf{x})=-p_{\mathbf{x}}\log(p_{\mathbf{x}})-(1-p_{\mathbf{x}})\log(1-p_{\mathbf{x}}),$ (5)
where $p_{\mathbf{x}}$ represents the probability of failure at input $\mathbf{x}$. With a GP surrogate, this probability boils down to a Gaussian CDF computation:
$p_{\mathbf{x}}=\mathrm{P}\left(\hat{f}(\mathbf{x})>t\right)=1-\Phi\left(\frac{t-\mu_{n}(\mathbf{x})}{\sigma_{n}(\mathbf{x})}\right).$
Returning to the Herbie example, the entropy surface for the trained GP is shown in the right panel of Figure 1 and along the slice in Figure 2 (green dotted/dashed). Notice the high entropy regions mostly agree with the locations of the true failure contour (we could improve upon this with a CL design instead of an LHS). A reasonable acquisition could fall in any of these local optima. Booth et al., (2024b) extended these ideas, with some necessary adaptations, to enable contour location with Bayesian DGP surrogates. Henceforth, all GPs and DGPs we consider will be from sequential CL designs.
2.2 Importance sampling
Rather than side-stepping the prohibitive computation of Eq. (1) with surrogate evaluations, importance sampling (Tokdar and Kass,, 2010) avoids expensive evaluations by sampling from a different distribution altogether. Let $q(\mathbf{x})$ denote a “bias” density which shares domain with $p(\mathbf{x})$. The failure probability $\alpha$ may be equivalently represented as
[…]
Monte Carlo approximation of this integral provides the estimator
$\hat{\alpha}_{\mathrm{IS}}=\frac{1}{B}\sum_{i=1}^{B}w_{i}\;\mathbb{I}_{\left\{f(\mathbf{x}_{i})>t\right\}}\quad\textrm{where}\quad w_{i}=\frac{p(\mathbf{x}_{i})}{q(\mathbf{x}_{i})}\quad\textrm{for}\quad\mathbf{x}_{i}\sim q(\mathbf{x}).$ (6)
The “bias weights” $w_{i}$ account for the difference between the input density and the bias density. This estimator is theoretically unbiased and requires a smaller budget of simulator evaluations ($B\ll M$) if the bias distribution appropriately covers and targets the failure region (Srinivasan,, 2002). But estimation of an effective bias distribution is tricky. The ideal bias distribution is $q(\mathbf{x})^{*}=\frac{1}{\alpha}\mathbb{I}_{\left\{f(\mathbf{x})>t\right\}}p(\mathbf{x})$, but naturally neither $\alpha$ nor the indicator function are known a priori. For the sake of feasibility, $q(\mathbf{x})$ is traditionally constrained to a known distributional family, such as a Gaussian mixture model (GMM; Reynolds et al.,, 2009). Bias distribution training is then relegated to learning distributional parameters (in a GMM, these include the number of mixtures, mixture weights, and component means and variances), often accomplished through cross validation or information criteria maximization.
In “surrogate-informed importance sampling,” Peherstorfer et al., (2016) leverage a surrogate to identify predicted failures from a large sample of the input distribution, then train a GMM bias distribution to these predicted failures. If the surrogate effectively identifies the failure region, then these samples should mimic the ideal $q(\mathbf{x})^{*}$. With a trained bias distribution in-hand, samples may be drawn from $q(\mathbf{x})$ and deployed in Eq. (6) to provide an estimate of the failure probability.
While SIIS has been widely embraced for its unbiasedness, it has several key weaknesses. First, the bias distribution, which features heavily in the final estimate of Eq. (6), is a completely auxilary quantity. Hypothetically, two different bias distributions could produce the same samples $\left\{\mathbf{x}_{i}\right\}_{i=1}^{B}$, thus producing the same $\left\{f(\mathbf{x}_{i})\right\}_{i=1}^{B}$. But in the calculation of Eq. (6), these two bias distributions would contribute different weights, resulting in different final answers even though the training data from the computer simulator was exactly the same. O’Hagan, (1987) first pointed out this conundrum as a violation of the “Likelihood principle.” A second weakness is the restriction of $q(\mathbf{x})$ to distributional families like GMMs. Failure regions may be complex and may not fit nicely under these target distributions. Consequently, SIIS typically requires large $B$ in order to adequately cover the failure region. For context, Peherstorfer et al., (2016) entertained $B\in\left\{100,1000,10000,100000\right\}$ and found performance continued to improve as $B$ increased. Furthermore, there is no clear avenue for determining the budget allocation between surrogate training and importance sampling estimation. Finally, the SIIS method does not leverage the full potential of the surrogate model. Many expensive simulator evaluations go into training the surrogate, and in the end the surrogate is only used to identify predicted failures, with no mention of its UQ. We suspect this is a byproduct of Peherstorfer et al., using reduced order models, which do not provide the UQ of a GP, in their seminal work. Ultimately, we propose a flexible estimation procedure which avoids bias distributions altogether, leverages surrogate UQ, and enables strategic budget allocation.
3 Hybrid Monte Carlo with GPs
In this section, we introduce a failure probability estimation scheme which (i) relies only on the surrogate model and the simulator data, (ii) leverages surrogate UQ in the estimation phase, and (iii) can provide effective estimates with a budget of only several hundred. Specifically, we will upgrade the Monte Carlo estimate of Eq. (2) so it requires fewer evaluations of $f(\mathbf{x})$ with minimal reduction in accuracy.
Our approach (and those of our competitors) may be organized into two sequential phases: the training of the surrogate and the estimation of the failure probability. Let $N$ denote the number of simulator evaluations used to train the surrogate, and let $B$ denote the remaining budget of simulator evaluations available for estimation. For the duration of this section, we assume both $N$ and $B$ are pre-determined; in Section 4 we will consider how to allocate evaluations between them. Recall, we are targeting expensive simulations where the total budget, $N+B$, is constrained to be in the hundreds.
3.1 Methodology
Let $\mathcal{U}\subset\mathcal{X}$ represent some subset of the input domain (boundary inclusive). We may partition the integral of Eq. (1) into two components as follows, with the corresponding Monte Carlo estimate,
[…]
Now, replace expensive evaluation of $f(\mathbf{x}_{i})$ outside of $\mathcal{U}$ with the surrogate prediction $\mu_{N}(\mathbf{x}_{i})$, resulting in the following “hybrid Monte Carlo” estimator:
[…]
The hybrid nature of this estimator is in the use of both true simulator evaluations (left term) and surrogate evaluations (right term). Each $\mathbf{x}_{i}$ will fall under either the left indicator term or the right one, never both. The computational cost of this estimator is greatly reduced since evaluations of the surrogate are effectively free. The accuracy of this estimator is contingent upon the classification accuracy of the surrogate and the choice of $\mathcal{U}$. Here, we focus on the choice of $\mathcal{U}$, presuming we already have a trained GP-based surrogate in-hand. [Training a surrogate for contour location has been previously investigated, as we reviewed in Section 2.1.]
Rather than thinking of $\mathcal{U}$ as a continuum, we find it easier to conceptualize our contribution with a discrete set of samples, which form the basis of any Monte Carlo estimator. Let $X_{M}$ denote the row-combined matrix of $M$-many samples from $p(\mathbf{x})$. To deploy our estimator $\hat{\alpha}_{\textrm{HYB}}$, we need a decision rule for classifying each $\mathbf{x}_{i}$ into or out of $\mathcal{U}$. In a perfect world, $\mathcal{U}$ would capture all the samples which the surrogate would misclassify. We are not privy to this information, but we can leverage the surrogate’s UQ to identify the samples at which it is most uncertain. Recall, entropy (Eq. 5) quantifies the degree of uncertainty in the pass/fail prediction of a surrogate. Accordingly, our decision rule is strikingly simple – place the $B$-many samples from $X_{M}$ with the highest entropy into $\mathcal{U}$. Denote these as $X_{M}^{\mathcal{U}}$. Hopefully our choice of the letter $\mathcal{U}$ is now apparent - it denotes the inputs with the highest “$\mathcal{U}$ncertainty”. Revisiting Figure 2, notice the high entropy regions (green dotted/dashed) align well with the misclassification regions (yellow). The entropy criterion is effectively identifying regions of high classification uncertainty where we could most utilize expensive true simulator evaluations.
Li and Xiu, (2010) employed a hybrid estimator similar to that of Eq. (7), but they chose to allocate expensive simulator evaluations to samples predicted to be closest to the contour. We suspect this choice was motivated by their use of polynomial chaos expansion models as surrogates which do not facilitate UQ. Nevertheless, we will implement this allocation scheme with our hybrid Monte Carlo estimator as a competitor in our benchmark exercises later.
For a fixed $N$, $B$, and $M$, our method proceeds as follows. First, train the surrogate using $N$-many evaluations of the high fidelity simulator. Our method is agnostic to surrogate and design choice as long as the surrogate provides posterior predictive distributions for the calculation of entropy, but we prefer contour locating designs with GP or DGP surrogates. Second, sample a very large number of points $X_{M}\sim p(\mathbf{x})$ and use the surrogate to calculate the classification entropy for each sample following Eq. (5). Third, identify the $B$-many points in $X_{M}$ with the highest entropy and evaluate the simulator at these locations. Fourth, update the surrogate with these new observations of the simulator. Finally, estimate $\hat{\alpha}_{\textrm{HYB}}$ following Eq. (7) by combining the simulator evaluations at $X_{M}^{\mathcal{U}}$ with surrogate predictions at all other samples of $X_{M}$.
This fourth step – updating the surrogate given the simulator data at the high-entropy sampled locations – is significant. In the end, the surrogate is trained on $N+B$ observations, making full use of all the available data. Since the $X_{M}^{\mathcal{U}}$ points are likely near a failure contour, they could significanly improve the accuracy of the surrogate near the contour. The ability to update the surrogate and make full use of all simulator data is a unique advantage over SIIS, whose final estimate is contingent only upon the $B$-many points sampled from the bias distribution.
There are some noteworthy distinctions between the training and estimation phases of our method. The $N$ points used to initially train the surrogate are acquired anywhere in the domain $\mathcal{X}$, and assuming they stem from a contour locating design, they seek to balance both exploration of the response surface and exploitation of the suspected contour. In fact, the contour locating sequential design strategies we will use next (Cole et al.,, 2023; Booth et al.,, 2024b) intentionally circumvent acquisitions based on maximal entropy in order to avoid clustered acquisitions and to promote more exploration of the surface. Alternatively, the $B$ points used for the hybridization and final surrogate update must originate from $p(\mathbf{x})$, and they are solely exploitative. By nature, they will be clustered around the predicted contour. We contend that having a mix of data from these two realms is superior to either extreme. To demonstrate superiority over the first extreme, in which the entire budget is relegated to training the surrogate over $\mathcal{X}$ ignorant of $p(\mathbf{x})$, we will offer empirical comparisons in our simulated examples. To justify this claim over the second extreme, in which all training data must stem from $p(\mathbf{x})$, we revisit the Herbie example of Figure 1. With the input distribution centered at the origin, the probability of generating any samples outside $[-1,1]^{2}$ is very low. But a GP surrogate would benefit from some “anchor points” in these outer regions (Gramacy and Apley,, 2015). This distinction will be less prominent if $p(\mathbf{x})$ is uniformly distributed over $\mathcal{X}$, but the trade-off between exploring in the training phase and exploiting in the estimation phase will still be impactful.
Choosing M.
Larger $M$ will produce more effective sum approximations of the true integrals, but we need surrogate predictions (e.g., Eq. 3) for every sample in $X_{M}$. Although this computation is significantly cheaper than evaluation of the expensive simulator, it may still be significant for large $M$, particularly with a heftier surrogate like a Bayesian DGP. Simply put, we would like $M$ to be large enough without being too cumbersome.
To reason through this choice, let’s revisit the unbiased Monte Carlo estimate of Eq. (2). Notice the indicator $\mathbb{I}_{\left\{f(\mathbf{x})>t\right\}}$ is a Bernoulli random variable with expected value $\alpha$ and variance $\alpha(1-\alpha)$. Thus, the MC estimator has variance inversely proportional to $M$:
$\mathbb{V}[\hat{\alpha}_{\textrm{MC}}]=\frac{1}{M^{2}}\sum_{i=1}^{M}\mathbb{V}\left[\mathbb{I}_{\left\{f(\mathbf{x})>t\right\}}\right]=\frac{\alpha(1-\alpha)}{M}.$
Let $\sigma_{\alpha}=\sqrt{\frac{\alpha(1-\alpha)}{M}}$ denote the standard error of this estimator. If we knew $\alpha$, we could determine $M$ by thresholding this standard error. For example, to achieve a standard error less than 10% of $\alpha$ requires:
$\sigma_{\alpha}<\frac{\alpha}{10}\quad\longrightarrow\quad\frac{100(1-\alpha)}{\alpha}<M.$
To achieve this precision, $\alpha=10^{-4}$ requires $M$ nearly 1 million, $\alpha=10^{-5}$ requires $M$ nearly 10 million, and $\alpha=10^{-6}$ requires $M$ nearly 100 million.
In reality, when the true $\alpha$ is unknown, we recommend either choosing $M$ conservatively given expert insight into the anticipated failure probability or performing a pilot study where $M$ is incremented say by one million at a time. Surrogate predictions are gathered along this incremental process, with $\alpha$ estimated along the way to inform the selection of $M$.
3.2 Synthetic experiments
In this section we provide a variety of synthetic experiments in which the training and estimation budgets, $N$ and $B$ respectively, are pre-determined. We start with three functions: the 2d Herbie function (Lee et al.,, 2011) with a total budget of 160, the 3d Ishigami function (Ishigami and Homma,, 1990) with a total budget of 400, and the 6d Hartmann function (Picheny et al.,, 2013) with a total budget of 800. We split the total budget of evaluations evenly between training and estimation. Table 1 summarizes these settings, among others. Functions are observed without noise to mimic a deterministic computer simulation. We define failure thresholds and input distributions for each function, resulting in the failure probabilities reported in Table 1. We choose $M$ based on these $\alpha$ values to keep $\sigma_{\alpha}$ conservatively under 5%. All input distributions are variations/combinations of independent uniform and truncated normal distributions; further details of the functions, thresholds, and distributions are reserved for Supplementary Material.
Function True $\alpha$ Surrogate $d$ $n_{0}$ $N$ $B$ $M$
Herbie $7.533\times 10^{-5}$ GP 2 20 80 80 $3.5\times 10^{7}$
Ishigami $1.904\times 10^{-4}$ GP 3 50 200 200 $1.5\times 10^{7}$
Hartmann $1.001\times 10^{-5}$ GP 6 100 400 400 $1.0\times 10^{8}$
Plateau $8.530\times 10^{-4}$ DGP 4 30 120 120 $3.5\times 10^{7}$
Table 1: Simulaton settings. True $\alpha$ values are estimated from 10 billion samples.
We begin with a random Latin hypercube sample (McKay et al.,, 2000) of size $n_{0}$ (chosen to be 25% of the total training budget, see Table 1). We train a GP surrogate using the Scikit-learn package in python (Pedregosa et al.,, 2011). We then conduct a sequential design following the entropy contour location scheme of Cole et al., (2023), until we reach a final design of size $N$. Ultimately, we deploy the trained surrogate, in conjunction with a large sample $X_{M}\sim p(\mathbf{x})$, to estimate $\alpha$ with each of the following methods:
  • hybMC ENT: Our hybrid MC estimate (Eq. 7) where $\mathcal{U}$ is chosen to contain the $B$ points from $X_{M}$ with highest entropy. [For now, ignore *hybMC ENT, we will return to it in Section 4.]
  • GP MC $N$: Surrogate MC estimate (Eq. 4) using the surrogate trained on $N$ points.
  • GP MC $N+B$: Surrogate MC estimate (Eq. 4) using a surrogate trained on $N+B$ points.
  • hybMC CLOSE: Inspired by Li and Xiu, (2010), the hybrid MC estimate (Eq. 7) where $\mathcal{U}$ is chosen to contain the $B$ points from $X_{M}$ which are predicted to be closest to the failure threshold (i.e., the samples with the smallest $|\mu_{N}(\mathbf{x}_{i})-t|$).
  • SIIS: Importance sampling estimate (Eq. 6; Peherstorfer et al.,, 2016) with a GMM bias distribution, also fit with the Scikit-learn package. The bias distribution is trained on the predicted failures, i.e., $\left\{\mathbf{x}_{i}\in X_{M}\;|\;\mu_{N}(\mathbf{x}_{i})>t\right\}$.
  • SIIS UCB: Same as SIIS, but the bias distribution is trained on the samples whose upper 90% confidence bound exceeds the threshold, i.e., $\left\{\mathbf{x}_{i}\in X_{M}\;|\;\mu_{N}(\mathbf{x}_{i})+1.645\sigma_{N}(\mathbf{x}_{i})>t\right\}$.
Note, the first four methods are variations of Monte Carlo, and the last two are variations of importance sampling. Crucially, each method uses the exact same trained surrogate (besides GP MC $N+B$, which is trained further) and the exact same $X_{M}$ sample. Differences in performance are thus attributed entirely to the estimation method. The GP MC $N$ method is the only one which does not utilize the entire budget; it serves as a snapshot of the surrogate’s accuracy before it is combined with a hybrid calculation or importance sampling. We repeat each exercise 30 times with re-randomized starting designs.
To demonstrate functionality with a DGP, we additionally consider the 4d plateau function (Booth et al.,, 2024b). This function is known to be non-stationary; it is characterized by flat regions with a steep sloping drop between them. To address the complexity in the response surface, we use a Bayesian DGP surrogate (Sauer et al.,, 2023b) fit with the deepgp R-package (Booth,, 2023). We train each DGP using the contour location scheme of Booth et al., (2024b), which involves proposing triangulation candidates (Gramacy et al.,, 2022) and selecting the candidate on the Pareto front of entropy and uncertainty. To keep all of the code for this example in R for easy integration with the DGP, we leverage the mclust R-package (Scrucca et al.,, 2016) to train the GMM bias distribution. Given the increased complexity of the surrogate, we perform 20 re-randomized repetitions. Reproducible code for all experiments, including both the python and R implementations, is available in our public git repository. [^2] https://bitbucket.org/boothlab/failprob/
Results are shown in Figure 3, with true failure probability indicated by the red dashed line. The grey shaded region highlights the interval $\{\alpha\pm 2\sigma_{\alpha}\}$ where $\sigma_{\alpha}$ is calculated from the provided $M$. This region provides important context – it captures the intrinsic variability that would accompany any MC estimator. In other words, if we used the true simulator for all $\{\mathbf{x}_{i}\}_{i=1}^{M}$ samples (Eq. 2), we would obtain a boxplot roughly spanning the grey interval. We thus consider any results within this region to be effective estimates. We could tighten the interval (and the boxplots of the MC methods) by increasing $M$ without requiring any additional simulator evaluations, another advantage MC holds over IS.
Figure 3: Estimated failure probabilities shown over true failure probability (red dashed) ±$\pm 2$ standard errors (grey shading). Boxplots show 30 repetitions (20 for Plateau).
Our hybMC ENT method (teal) performed well across the board. Unsurprisingly, it was able to improve upon the accuracy of the original surrogate (purple) by leveraging the additional budget of $B$ evaluations. More importantly, for all but the simplest Herbie function, it outperformed the surrogate MC estimates which made use of the full budget (blue). Recall, the (D)GP MC methods acquire points from $\mathcal{X}$, while the hybMC method acquires points from $p(\mathbf{x})$ (after the original training phase). The fact that hybMC is able to outperform the fully-trained surrogate indicates that at some point, it is better to prioritize expensive simulator evaluations for uncertain samples from $p(\mathbf{x})$ than to continue training the surrogate.
The strategy of Li and Xiu, (2010) in allotting simulator evaluations based on predicted distance to the contour (hybMC CLOSE, orange) did not significantly improve upon the accuracy of the surrogate alone. This idea fails to leverage the UQ of the surrogate. It is possible for the surrogate to have higher uncertainty in its pass/fail prediction for a point further from the contour. For example, refer back to Figure 2. The entropy surrounding $x_{1}=1$ is far greater than that surrounding $x_{1}=0.5$, even though the latter of these is predicted to be closer to the contour (blue dashed line). Reserving true simulator evaluation for the $x_{1}=1$ location, where the surrogate would misclassify a failure, is preferable.
The SIIS methods (green and pink) had the poorest performance across the board. They particularly struggled in the simpler settings (Herbie and Ishigami), underperforming the simple GP MC $N$ method which only used half the amount of simulator evaluations. We suspect the small budgets are to blame here; IS just needs more data, particularly when the GMM bias density is far from the ideal. Consequently, these IS variations had their best performance with the Hartmann function where 400 samples were budgeted for estimation. There appears to be very little benefit to the more conservative SIIS UCB (pink) over the traditional SIIS (green).
4 Allocating the simulation budget
Let $T=N+B$ represent the total budget of simulator evaluations, which is typically determined by the computational cost of the simulator. For expensive simulators which require long compute times, $T$ may be only several hundred. We presume $T$ is fixed, and we now address how to split it between $N$ and $B$. In our previous studies, we evenly split $T$ between training the surrogate and estimating $\alpha$, i.e., $N=B=\frac{1}{2}T$. But this decision was admittedly arbitrary. In this section, we propose a scheme for allocating $T$ between $N$ and $B$, specifically for use within our hybrid Monte Carlo estimation scheme. Since the sum of $N$ and $B$ is fixed, this choice boils down to selecting $N$; $B=T-N$ then follows directly. Ideally, we would like to halt training of the surrogate once learning has plateaued, saving the rest of the budget for the hybrid Monte Carlo estimate. If we stop training too early, the surrogate might be too inaccurate for the hybrid MC estimate to overcome. If we continue training for too long, we will not have any samples left for the hybrid estimation procedure. As we saw in Figure 3, in all but the simplest case, a MC estimate without any hybridization is suboptimal.
Figure 4: Left: Surrogate MC failure probabilty estimate for the Ishigami function, updated every 10th acquisition with true $\alpha$ in dashed red. Center/Right: Difference in successive estimates with ±$\pm\hat{\sigma}_{\alpha}$ shown in dashed green. The blue square marks the second update within ±$\pm\hat{\sigma}_{\alpha}$.
To monitor the surrogate’s learning, we propose tracking the progress of the surrogate MC estimate, $\hat{\alpha}_{\textrm{SURR}}$ (Eq. 4), which we will now denote as simply $\hat{\alpha}$. This estimator relies only on the trained surrogate and does not require additional simulator evaluations. As a demonstration, the left panel of Figure 4 reports the progress in $\hat{\alpha}$ for the 3d Ishigami function following the setup described in Section 3.2. To avoid excessive calculations, we evaluate the estimate after every 10th acquisision. It is apparent that the surrogate’s learning of the failure contour saturates well before it reaches $N=200$, which we used in our previous experiments. The center panel of Figure 4 shows the relative difference between each successive $\hat{\alpha}$. As expected, the relative differences converge to zero as the surrogate hones in on the true contour, but they never settle at exactly zero (this is more apparent in the right panel which is zoomed in).
How small of an update then is small enough to indicate learning has plateaued? The significance of an update to $\hat{\alpha}$ is only interpretable within the context of $\alpha$ and $M$. Small differences might be inconsequential for large $\alpha$ with small $M$, but could be significant for small $\alpha$ with large $M$. To this end, we propose the use of the standard error $\sigma_{\alpha}$ to quantify the relative magnitude of each update. This standard error intrinsically accounts for the magnitude of $\alpha$ and $M$. Since the true $\alpha$ is unknown, we estimate $\hat{\sigma}_{\alpha}=\sqrt{\frac{\hat{\alpha}(1-\hat{\alpha})}{M}}$ using the current surrogate MC estimate. The green dashed lines in the center and right panels of Figure 5 show ±$\pm\hat{\sigma}_{\alpha}$ for the estimated failure probabilities. At the start, updates to $\hat{\alpha}$ are far beyond one standard error. But after a while, the updates sit comfortably within these bounds. We posit that, for a fixed sample $X_{M}$, any update to $\hat{\alpha}$ within ±$\pm\hat{\sigma}$ is “small enough.” Accordingly, we halt surrogate training once we have observed two successive updates within ±$\pm\hat{\sigma}$.
Our budgeting scheme is summarized as follows. Initialize a surrogate, begin a contour locating sequential design, and collect one sample $X_{M}$. At the start, it is likely that there will be no failures observed and that $\hat{\alpha}$ will be equal to zero (as is the case in Figure 4). To safeguard against halting surrogate training prematurely, we build in a “common sense” check by setting a minimum number of failures that must be observed before the contour location could be halted. If the contour location design is succeeding, it will be placing some points in the failure region. We recommend a default of 10 failures, but suggest increasing this for higher dimensional problems. In our synthetic experiments, we used a minimum of 10 failures for the Herbie, Ishigami, and plateau functions but bumped this up to 40 for the larger dimension of the Hartmann function. Once the surrogate has surpassed these minimum requirements, estimate $\hat{\alpha}$ after every 10th acquisition. Calculate $\hat{\sigma}_{\alpha}$ and check if $|\hat{\alpha}^{\textrm{(current)}}-\hat{\alpha}^{\textrm{(previous)}}|<\hat{\sigma}_{\alpha}$. Once this condition has been satisfied twice in a row, halt the sequential design and proceed to the hybrid Monte Carlo estimation following Section 3.1.
This stopping criterion requires the same $X_{M}$ sample be used in every $\hat{\alpha}$ calculation. If a new $X_{M}$ is generated for each estimate, there will naturally be more variability among the $\hat{\alpha}$’s, and the ±$\pm\hat{\sigma}_{\alpha}$ threshold would need to be increased. Removing this extraneous source of variability allows successive estimates to pinpoint differences stemming solely from the surrogate.
Figure 5: Distribution of the chosen $N$ for each repetition of the synthetic examples from Figure 3. Red star indicates the value that was used previously.
Revisiting synthetic experiments.
We implemented this stopping criteria with each of the four functions described in Section 3.2. We used the same total budgets ($T=\{160,400,800,240\}$, respectively), but chose $N$ based on this decision rule. Once $N$ was determined, we set $B=T-N$. All other simulation settings were the same. Figure 5 shows the distribution of chosen $N$ values. The red star indicates the value used previously. Most repetitions halted surrogate training before the preset value, allotting more runs for estimation, but some proceeded longer. We view this variability as a strength – some random initializations need more data to effectively locate the contour while some get lucky from the start. Allowing the surrogate to dictate its own budget mitigates some of the risks of having an unlucky starting design. The resulting failure probability estimates are shown as the *hybMC ENT method (yellow) in Figure 3. The original hybMC ENT methods already performed well, and we were able to match this performance with the budgeted *hybMC ENT method. We acknowledge the benefit of the budgeting method is not in beating the un-budgeted performance on synthetic problems, but in informing budget allocations in real-world scenarios when there is no clear choice for $N$ and $B$.
5 RAE-2822 airfoil computer experiment
The RAE-2822 airfoil computer experiment simulates the flow of air around an aircraft wing. The simulation suite is solved via SU2, and the software is publicly available (Economon et al.,, 2016). The experiment accepts seven inputs: four shape parameters ($S_{i}$ for $i=1,\dots,4$) and 3 environmental parameters, namely angle of attack, Reynolds number, and Mach number. We are interested in the lift over drag ($L/D$) ratio. $L/D$ values that are too low represent inefficient flight conditions. We define a failure threshold at $t=3$ such that $L/D<3$ indicates failure. [Note the change from $f(\mathbf{x})>t$ to $f(\mathbf{x})<t$ does not affect the implementation of our method.] This computer experiment requires nearly 30 minutes of compute time with serial execution, but reductions are possible with parallel execution.
Our objective is to estimate the probability of failure given the following input distribution:
$S_{i}$ $\sim\mathrm{Unif}(-1e-4,1e-4)\;\;\textrm{for}\;\;i=1,\dots,4$
Angle of Attack $\sim\mathcal{N}\left(\mu=5,\sigma=1\right)\;\;\textrm{truncated to}\;\;[0,10]$
Reynolds number $\sim\mathcal{N}\left(\mu=1e7,\sigma=1e6\right)\;\;\textrm{truncated to}\;\;[5e6,1.5e7]$
Mach number $\sim\mathcal{N}\left(\mu=0.8,\sigma=0.02\right)\;\;\textrm{truncated to}\;\;[0.7,0.9]$
We fix our total budget at $T=500$ simulator evaluations. Booth et al., (2024b) found that DGP surrogates outperformed GP surrogates for contour location of this experiment. We follow the settings therein, starting with an LHS sample of size $n_{0}=100$ and fitting a Bayesian DGP surrogate with the deepgp package. We proceed with contour location following Booth et al., (2024b), stopping after every 10th acquisition to estimate $\hat{\alpha}$ and check whether it is with in the ±$\pm\hat{\sigma}_{\alpha}$ bounds. Given the input dimension, we require atleast 40 observed failures before we consider halting surrogate training. Since the true failure probability is unknown, we repeat this entire exercise with a newly randomized starting design for some additional context. Figure 6 reports the progress in $\hat{\alpha}$ over the course of each design.
Figure 6: Two surrogate MC failure probabilty estimates for the airfoil computer experiment, updated every 10th acquisition with final estimates in dashed red. Grey shading indicates $\hat{\alpha}\pm\hat{\sigma}_{\alpha}$ for the final estimates.
Our stopping criterion halted the sequential designs at $N=180$ and $N=160$. For each repetition, we collect a separate random sample of size $M=2.5\times 10^{6}$, then identify the remaining points with the highest entropy, and evaluate the true simulator at these locations. After updating the DGP surrogate, we obtain the final hybrid Monte Carlo estimates of $\hat{\alpha}_{\textrm{HYB}}=\{0.0003368,0.0003432\}$, respectively. These values are indicated by the red dashed lines in Figure 6. The grey shading shows $\hat{\alpha}\pm\hat{\sigma}_{\alpha}$, calculated from these final hybrid estimates. Although the two estimates vary slightly, they are both within the bounds of reasonable variability for our chosen $M$. Additionally, notice how the final hybrid estimate of the second exercise (right panel) brought the failure probability lower than what the surrogate had estimated, closer to the estimate of our first attempt (left panel). We take this as an indication that the hybrid estimator is effectively adjusting for surrogate inaccuracies.
6 Discussion
We presented a hybrid Monte Carlo method for estimating failure probabilities of expensive computer experiments. Our hybrid method leverages the uncertainty quantification provided by a surrogate to reserve expensive simulator evaluations for samples with high classification uncertainty. We limited our experiments to GP and DGP surrogates, but our proposed estimator is applicable to any surrogate which provides posterior predictive distributions (required for calculation of entropy). Extensions to other surrogates, such a treed Gaussian processes (Gramacy and Lee,, 2008) or Bayesian additive regression trees (Chipman et al.,, 2010), is straightforward.
Our hybrid MC estimator is computationally efficient in the number of evaluations of the expensive simulator, but it requires many predictive evaluations from the surrogate. Posterior predictions must be gathered at all $M$ sample points. While surrogate predictions are very cheap, collecting millions of them is not necessarily trivial. Our experiments were limited to failure probabilities above $1\times 10^{-5}$ for this very reason. We could tackle smaller failure probabilities, but it would require larger $M$, which would eventually become impractical. Our approach may simply not be suited to very rare failure probabilities.
The efficacy of any surrogate-informed failure probability estimation relies on the surrogate’s ability to accurately classify failures. Although our hybrid estimator is generally applicable, we strongly recommend the use of contour locating sequential designs to train the surrogate. We proposed a budgeting scheme which informs allocation of simulator evaluations between surrogate training and failure probability estimation based on progress in the surrogate Monte Carlo estimate. If a budget is severly restricted, the budget could be exhausted training the surrogate before our proposed stopping criterion is met. While this is not ideal, we contend that fully-training a surrogate and using a surrogate MC estimate is the best option in this scenario. Particularly in light of the data hungry nature of IS-based methods, which perform poorly in small data settings.
When the input distribution is not uniform over $\mathcal{X}$, it could be beneficial to leverage information about $p(\mathbf{x})$ in the sequential training of the surrogate (Abdelmalek-Lomenech et al.,, 2024). There is arguably no need to train a surrogate in a region where inputs will not plausibly occur, such as in the outer corners of the Herbie function surface (Figure 1, center). Incorporating this information in the contour location stage could improve failure probability estimation downstream. But in the case of rare probabilities, it may be hard to draw the line between regions that are plausible and ones that are not. This avenue is currently underexplored.
SUPPLEMENTARY MATERIAL
Here we provide the details of the functions, thresholds, and input distributions used in the benchmark exercises of Section 3.2. In each case, failures are defined as $f(\mathbf{x})>t$.
Herbie.
The 2d Herbie function (Lee et al.,, 2011) is defined in over $[-2,2]^{2}$ as
$f(\mathbf{x})=\prod_{i=1}^{2}\mathrm{exp}\left(-(x_{i}-1)^{2}\right)+\mathrm{exp}\left(-0.8*(x_{i}+1)^{2}\right)-0.05*\sin(8*(x_{i}+1)).$
We set the failure threshold at $t=1.065$ and define the input distribution as $x_{i}\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathcal{N}(0,0.36)$ for $i=1,2$, truncated to $[-2,2]$.
Ishigami.
The 3d Ishigami function (Ishigami and Homma,, 1990) is defined over $[-\pi,\pi]^{3}$ as
$f(\mathbf{x})=\sin(x_{1})+5*\sin(x_{2})^{2}+0.1*x_{3}^{4}*\sin(x_{1}).$
We set the failure threshold at $t=10.244$ and define the input distribution as
$x_{1}$ $\sim\mathcal{N}(-1,1)\;\;\textrm{truncated to}\;\;[-\pi,\pi]$
$x_{2}$ $\sim\mathcal{N}(1.5,1.5)\;\;\textrm{truncated to}\;\;[-\pi,\pi]$
$x_{3}$ $\sim\mathrm{Uniform}(-pi,pi).$
Hartmann.
The 6d Hartmann function (Picheny et al.,, 2013) is defined over $[0,1]^{6}$ as
$f(\mathbf{x})=-\sum_{i=1}^{4}\alpha_{i}\mathrm{exp}\left(-\sum_{j=1}^{6}A_{ij}(x_{j}-P_{ij})^{2}\right)\;\;\textrm{where}$
$\alpha^{\top}=\begin{bmatrix}1\\ 1.2\\ 3\\ 3.2\end{bmatrix}\quad\quad A^{\top}=\begin{bmatrix}10&.05&3&17\\ 3&10&3.5&8\\ 17&17&1.7&.05\\ 3.5&.1&10&10\\ 7&8&17&.1\\ 8&14&8&14\end{bmatrix}\quad\textrm{and}\quad P^{\top}=\begin{bmatrix}.1312&.2329&.2348&.4047\\ .1696&.4135&.1451&.8828\\ .5569&.8307&.3522&.8732\\ .0124&.3736&.2883&.5743\\ .8283&.1004&.3047&.1091\\ .5886&.9991&.6650&.0381\end{bmatrix}.$
We set the failure threshold at $t=2.46$ and define the input distribution as $x_{i}\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathcal{N}(0.5,0.1)$ for $i=1,\dots,6$, truncated to $[0,1]$.
Plateau.
The Plateau function (Booth et al.,, 2024b) is defined in arbitrary dimension over $[0,1]^{d}$ as
$f(\mathbf{x})=2*\Phi\left[\sqrt{2}\left(-4-3\sum_{i=1}^{3}(4x_{i}-2)\right)\right]-1$
where $\Phi$ is the standard Gaussian CDF. We use $d=4$ and set the failure threshold at $t=0$. We define the input distribution as $x_{i}\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathcal{N}(0.5,0.085)$ for $i=1,\dots,4$, truncated to $[0,1]$.

### v2

@@
-Hybrid Monte Carlo for Failure Probability Estimation with Gaussian Process Surrogates
+Two-stage Design for Failure Probability Estimation with Gaussian Process Surrogates
 Abstract
-We tackle the problem of quantifying failure probabilities for expensive computer experiments with stochastic inputs.
-The computational cost of evaluating the computer simulation prohibits direct Monte Carlo (MC) and necessitates a statistical surrogate model.
-Surrogate-informed importance sampling – which leverages the surrogate to identify suspected failures, fits a bias distribution to these locations, then calculates failure probabilities using a weighted average – is popular, but it is data hungry and can provide erroneous results when budgets are limited.
-Instead, we propose a hybrid MC scheme which first uses the uncertainty quantification (UQ) of a Gaussian process (GP) surrogate to identify areas of high classification uncertainty, then combines surrogate predictions in certain regions with true simulator evaluation in uncertain regions.
-We also develop a stopping criterion which informs the allocation of a fixed budget of simulator evaluations between surrogate training and failure probability estimation.
-Our method is agnostic to surrogate choice (as long as UQ is provided); we showcase functionality with both GPs and deep GPs.
-It is also agnostic to design choices; we deploy contour locating sequential designs throughout.
-With these tools, we are able to effectively estimate small failure probabilities with only hundreds of simulator evaluations.
-We validate our method on a variety of synthetic benchmarks before deploying it on an expensive computer experiment of fluid flow around an airfoil.
-Keywords: computer experiment, deep Gaussian process, entropy, emulator, contour location, reliability, importance sampling
+We tackle the problem of quantifying failure probabilities for expensive deterministic computer experiments with stochastic inputs.
+The computational cost of the computer simulation prohibits direct Monte Carlo (MC) and necessitates a surrogate model, turning the problem into a two-stage enterprise (surrogate training followed by probability estimation).
+Limited budgets create a design problem: how should expensive evaluations be allocated between and within the training and estimation stages?
+One may use the entire evaluation budget to sequentially train the surrogate through contour location (CL), with failure probabilities then estimated solely from the surrogate (we call it “surrogate MC”).
+But extended CL offers diminishing returns, and surrogate MC relies too stringently on surrogate accuracy.
+Alternatively, a partially trained surrogate may inform importance sampling, but this can provide erroneous results when budgets are limited.
+Instead we propose a two-stage design: starting with sequential CL, halting CL once learning has plateaued, then greedily allocating the remaining budget to MC samples with high classification entropy.
+Ultimately, we employ a “hybrid MC” estimator which leverages the trained surrogate in conjunction with the true responses observed in this second stage.
+Our unique two-stage design strikes an appropriate balance between exploring and exploiting, and outperforms alternatives, including both of the aforementioned approaches, on a variety of benchmark exercises.
+With these tools, we are able to effectively estimate small failure probabilities with only hundreds of simulator evaluations, showcasing functionality with both shallow and deep Gaussian process surrogates, and deploying our method on a simulation of fluid flow around an airfoil.
+Keywords: computer experiment, contour location, deep Gaussian process, emulator, entropy, importance sampling, Monte Carlo
 1 Introduction
-Quantifying the probability of failure in complex dynamical systems is of great interest but is challenging due to the computational expense of the so-called “computer experiment.” Consider a black-box computer simulation, $f:\mathcal{X}\rightarrow\mathbb{R}$ with inputs $\mathcal{X}\subset\mathbb{R}^{d}$ governed by known measure $\mathbb{P}_{\mathbf{x}}$ (with density $p(\mathbf{x})$).
+Quantifying the probability of failure in complex dynamic systems is of great interest but is challenging due to the computational expense of the so-called “computer experiment.” Consider a deterministic black-box computer simulation, $f:\mathcal{X}\rightarrow\mathbb{R}$ with inputs $\mathcal{X}\subset\mathbb{R}^{d}$ governed by known measure $\mathbb{P}_{\mathbf{x}}$ (with density $p(\mathbf{x})$).
 The response is thresholded such that $f(\mathbf{x})>t$ indicates failure.
@@
 We are particularly motivated by aeronautic applications where system failures may incur immense financial or environmental costs, or even loss of life.
-In aircraft design and development, computer simulation experiments are a natural, often mandatory, precursor to direct experimentation and deployment.
+In aircraft design and development, computer simulation experiments are a natural, often mandatory, precursor to direct experimentation and prototyping.
 For example, $f(\cdot)$ may represent the vibrations experienced by an aircraft which become unsafe above certain levels, or the fuel efficiency of an aircraft, which if too low, can be detrimental to the environment, obliterate fuel budgets, or overly restrict travel distances.
-Inputs $\mathcal{X}$ may encapsulate design parameters and operating conditions (such as wing shapes, aircraft design, flight speeds, and angles of attack).
-Yet we acknowledge a slew of other relevant applications where quantifying failure probabilites for complex simulations is critical, including structural reliability analysis using finite element simulations (e.g., Mia et al.,, 2017), Boltzmann modeling of chemical processes in fuel filtration (e.g., Belot et al.,, 2021), and thermodynamic flows for nuclear reactors (e.g., Batet et al.,, 2014).
+Inputs $\mathcal{X}$ may encapsulate design parameters and operating conditions (such as wing shapes, flight speeds, and angles of attack).
+Yet we acknowledge a slew of other relevant applications where quantifying failure probabilities for complex simulations is critical, including structural reliability analysis using finite element simulations (e.g., Mia et al.,, 2017), Boltzmann modeling of chemical processes in fuel filtration (e.g., Belot et al.,, 2021), and thermodynamic flows for nuclear reactors (e.g., Batet et al.,, 2014).
 A rudimentary approach to solve Eq.
-(1) involves brute force Monte Carlo (MC); simply sample an exhaustively large set of inputs, evaluate the computer simulation at each of these configurations, and report the proportion that resulted in failure.
-But this approach is prohibitively expensive if the computer simulation is costly (incurring long wall clock times), the probability of failure is rare (requiring enourmous sample sizes), or both!
-In this work, we propose a novel method for quantifying failure probabilities of complex computer simulations when the computation budget restricts evaluations of the simulator to several hundred, rendering traditional MC impractical.
-In the face of expensive data, statistical models (called “surrogates” or “emulators”) are essential.
+(1) involves brute force Monte Carlo (MC): simply sample an exhaustively large set of inputs, evaluate the computer simulation at each of these configurations, and report the proportion that resulted in failure.
+But this approach is prohibitively expensive if the computer simulation is costly (incurring long wall clock times), the probability of failure is rare (requiring enormous sample sizes), or both!
+In this work, we propose a new strategy for quantifying failure probabilities of complex computer simulations when the computation budget restricts evaluations of the simulator to only a few hundred, rendering traditional MC impractical.
+In the face of expensive data, statistical models ($\hat{f}$, called “surrogates” or “emulators”) are essential.
 Surrogates are trained on a limited set of simulator evaluations in order to provide predictions with appropriate uncertainty quantification (UQ) at unobserved inputs.
-Gaussian processes (GPs) are the canonical choice (Santner et al.,, 2003; Gramacy,, 2020), but deep Gaussian processes (DGPs; Damianou and Lawrence,, 2013) are becoming increasingly popular, particularly for non-stationary computer experiments (Sauer,, 2023).
+Gaussian processes (GPs) are the canonical choice (Santner et al.,, 2018; Gramacy,, 2020), but deep Gaussian processes (DGPs; Damianou and Lawrence,, 2013) are becoming increasingly popular, particularly for nonstationary computer experiments (Sauer,, 2023).
 Notably, our contribution is agnostic to surrogate choice as long as effective UQ is provided, and is thus applicable to traditional GPs and Bayesian DGPs (Sauer et al.,, 2023b) and, potentially, Bayesian neural networks (MacKay,, 1995).
-Early methods of failure probability estimation, namely the first and second order reliability methods (FORM/SORM; Haldar and Mahadevan,, 1995), leveraged crude linear approximations of the complex model and placed all failure probability estimation in the hands of the “most probable point” on the predicted contour.
-Advancements on FORM and SORM have been made with more flexible GP models (e.g., Su et al.,, 2017).
-A variety of works have instead leveraged GP surrogates directly with MC sampling, using the GP predictions (which are of neglible cost) in place of true simulator evaluation (e.g., Cheng and Lu,, 2020; Li et al.,, 2021; Lu et al.,, 2023).
-We will use this “surrogate MC” method as a benchmark in our later exercises.
-While it is cheap, it places too much stock in the accuracy of the surrogate and funnels surrogate inaccuracies directly into the failure probability estimate.
-Importance sampling (IS; Srinivasan,, 2002; Tokdar and Kass,, 2010) offers an unbiased alternative.
-It hinges on the construction of a bias distribution that intentionally focuses density in the failure region.
-Inputs are sampled from the bias distribution (hopefully yielding many failures) and are then utilized in a weighted average calculation that corrects for the disparity between the bias density and the original input density.
-Identifying an effective bias distribution in the face of limited training data is a hefty task (Tabandeh et al.,, 2022).
-Adaptive methods abound (e.g., Oh and Berger,, 1992; Kurtz and Song,, 2013; Dalbey and Swiler,, 2014; Bugallo et al.,, 2017; Cheng et al.,, 2023).
-Peherstorfer et al., (2016) propose what they termed “multifidelity importance sampling,” although we prefer the term “surrogate-informed importance sampling” (SIIS) to avoid confusion with multifidelity surrogate modeling (e.g., Park et al.,, 2017).
-In SIIS, a surrogate is used to identify predicted failures, and a bias distribution is fit to these locations (more on this in Section 2.2).
-SIIS has been deployed with GP surrogates in a variety of subsequent works (e.g., Cole et al.,, 2023; Renganathan et al.,, 2023; Booth et al.,, 2024a; Renganathan,, 2024), and we view it as our primary competitor.
-Yet SIIS has several crucial drawbacks.
-We will delay a thorough discussion to Section 2.2, but the punchline is simply that IS has too much variability when bias distributions are suboptimal and data sizes are limited.
-The difficulty in training an effective bias distribution can render SIIS estimated failure probabilities ineffective, forming the motivation behind our current contribution.
-We argue that utility and precision in practice are more important than theoretical unbiasedness.
-To this end, we embrace a hybrid Monte Carlo calculation that uses both surrogate predictions and true simulator evaluations in strategically partitioned regions of the input space.
-Li and Xiu, (2010) first proposed such a hybrid calculation, but their work was restricted to polynomial chaos (PC) models with strict accuracy assumptions in order to guarantee convergence.
-GP-based surrogates afford several key advancements over PCs.
-For one, GPs are flexible and data-driven; their accuracy continuously improves as training data sizes increase.
-Perhaps more importantly, GPs offer UQ through posterior variances.
-We leverage this UQ to prioritize expensive model evaluation at locations where the surrogate is least certain in its pass/fail prediction.
-Regardless of how a surrogate is to be used (surrogate MC, SIIS, our hybrid MC, etc.), surrogate accuracy in identifying failure regions is of primary importance.
+Choosing a surrogate model, like a GP or DGP, is only a first step.
+The utility of a surrogate boils down to two things: how it is trained and how it is used.
+For our objective (1), surrogate training should prioritize accurate classification of failures and surrogate deployment should focus on effective estimation of $\alpha$.
+The prevailing question is then: how can a limited budget of expensive simulator evaluations be optimally allocated between and within the training and estimation stages?
+We briefly digress to acknowledge some early methods of failure probability estimation, namely the first and second order reliability methods (FORM/SORM; Haldar and Mahadevan,, 1995), which leveraged simplistic linear approximations of the complex model and placed all failure probability estimation in the hands of the “most probable point” (MPP).
+Although advancements to FORM and SORM have been made with more flexible GP models (e.g., Zhang et al.,, 2015; Su et al.,, 2017), sole reliance on a MPP is too restricting for our target applications.
+To solve Eq.
+(1), surrogate accuracy in classifying failures is of primary importance.
 Not all training designs are created equally.
@@
 Popular variations fall under expected improvement frameworks (Bichon et al.,, 2008; Ranjan et al.,, 2008), stepwise uncertainty reduction frameworks (Bect et al.,, 2012; Chevalier et al.,, 2014; Azzimonti et al.,, 2021), or both (Duhamel et al.,, 2023).
-Others utilize pass/fail predictive entropy (Marques et al.,, 2018; Cole et al.,, 2023).
+Others utilize pass/fail classification entropy (Marques et al.,, 2018; Cole et al.,, 2023).
 Booth et al., (2024b) recently extended CL sequential designs to DGP surrogates.
-We will utilize state-of-the-art sequentially designed GP and DGP surrogates in our benchmark exercises, but our contribution is agnostic to these choices.
-We see our method’s “plug-and-play” nature as a distinct advantage.
-Both surrogate training and failure probability estimation require a budget of expensive simulator evaluations.
-These two tasks are typically conducted in independent stages, with separate pre-determined budgets (as we will use in the first round of our synthetic benchmarks).
-Yet, in practice, it is common for a total budget to be set according to computational allowances, with no mention of how many evaluations to spend on the training phase versus the estimation phase.
-Some works have investigated stopping criteria for surrogate training in reliability settings (e.g., Echard et al.,, 2011; Hristov et al.,, 2019), but this is still an under-examined area of the literature.
-In our hybrid MC set-up, we have a unique advantage: the more we train our surrogate, the more certain it will be, requiring fewer high-cost evaluations for effective failure probability estimation.
-There is a clear give-and-take between the budgets of the training and estimation stages; increasing the budget of one decreases the budget required for the other, and vice versa.
-Leveraging this trade-off, we propose a method for allocating a fixed budget between surrogate training and hybrid Monte Carlo estimation.
+For small failure regions, contour location may be replaced with subset simulation (Au and Beck,, 2001) which initially targets a conservative threshold ($t^{\star}<t$) and iteratively refines it until the true threshold is found ($t^{\star}\rightarrow t$) (Dubourg et al.,, 2011; Huang et al.,, 2016; Hristov et al.,, 2019).
+Yet these approaches require Markov chain Monte Carlo (MCMC) sampling of cascading input distributions, which we intentionally avoid here.
+A well trained surrogate may be leveraged directly within an MC estimator, with surrogate predictions (which are of neglible cost) in place of true simulator evaluations.
+One could reasonably use the entire budget of simulator evaluations for contour location, then deploy the trained surrogate in such a “surrogate MC” estimator.
+Although this is common (e.g., Zhu and Du,, 2016; Cheng and Lu,, 2020; Li et al.,, 2021; Lu et al.,, 2023), we find that surrogate learning of failure contours offers quickly diminishing returns.
+Furthermore, surrogate MC estimators place too much stock in the accuracy of the surrogate and funnel surrogate inaccuracies directly into the failure probability estimate.
+Another common approach splits the budget of simulator evaluations into two disjoint stages: some evaluations are used to train the surrogate and some are used for importance sampling (IS; Srinivasan,, 2002; Tokdar and Kass,, 2010).
+IS hinges on the construction of a bias distribution that intentionally focuses density in the failure region.
+Inputs are sampled from the bias distribution and are then utilized in a weighted average calculation that corrects for the disparity between the bias density and the original input density.
+Identifying an effective bias distribution in the face of limited training data is a hefty task (Tabandeh et al.,, 2022).
+Adaptive methods abound (e.g., Oh and Berger,, 1992; Kurtz and Song,, 2013; Dalbey and Swiler,, 2014; Bugallo et al.,, 2017; Cheng et al.,, 2023).
+Peherstorfer et al., (2016) propose what they termed “multifidelity importance sampling,” although we prefer the term “surrogate-informed importance sampling” (SIIS) to avoid confusion with multifidelity surrogate modeling (e.g., Park et al.,, 2017).
+In SIIS, a trained surrogate is used to identify predicted failures, and a bias distribution is fit to these locations (more on this in Section 2.3).
+SIIS has been deployed with GP surrogates in a variety of subsequent works (e.g., Cole et al.,, 2023; Renganathan et al.,, 2023; Booth et al.,, 2024a; Renganathan,, 2024), and we view it as our primary competitor.
+Yet SIIS has several crucial drawbacks.
+We will delay a thorough discussion to Section 2.3, but the highlights are: (i) SIIS relies too heavily on an arbitrary bias distribution, (ii) it largely ignores the carefully trained surrogate, and (iii) it offers no guidance for allocating the budget between surrogate training and IS estimation.
+We prefer to avoid a bias distribution altogether, instead focusing on how we could optimally leverage a surrogate and an MC estimator in tandem.
+The answer lies in a two-stage design.
+First, we leverage sequential contour location to pin down the failure contour.
+At the start, it is essential to target inputs which provide the most insight into the entire failure contour (getting the most “bang for our buck” from early observations).
+As training progresses, successive acquisitions will offer less additional insight, indicated by apparent convergence of the surrogate MC estimator.
+Although we can not be sure of theoretical convergence with our limited budget, we propose a decision rule to recognize practical convergence (when surrogate learning from CL has saturated).
+We then spend the remaining simulation budget precisely on the Monte Carlo samples with the highest classification entropy, ultimately training the surrogate on observations from both stages.
+Our final failure probability estimate is a “hybrid MC” estimator which combines the observations from this second stage with surrogate predictions at unobserved samples.
 This manuscript is laid out as follows.
-First, in Section 2, we review GP/DGP surrogates and competing methods of failure probability estimation.
-In Section 3 we present our proposed hybrid Monte Carlo procedure, including implementation details and a variety of synthetic benchmarking exercises.
-In Section 4 we propose our budget allocation scheme, continuing the established benchmarks.
-Finally, we deploy our method on a computer experiment of an RAE-8288 airfoil in Section 5 and conclude in Section 6.
-2 Failure probability estimation
+Section 2, reviews GP/DGP surrogates, contour location, and importance sampling.
+Section 3.2 presents our two-stage design and failure probability estimator.
+Section 4 presents a variety of synthetic benchmarks.
+Finally, we deploy our method on a computer experiment of an RAE-2822 airfoil in Section 5 and conclude in Section 6.
+2 Preliminaries
 Denote single inputs as the row vector $\mathbf{x}$ of size $1\times d$ with response $y=f(\mathbf{x})$.
@@
 This estimator is ineffective unless a sufficient number of failures are observed, necessitating very large $M$ particularly for small $\alpha$.
-But our computer model $f(\mathbf{x})$ is expensive; we simply cannot evaluate it thousands/millions/billions of times.
-We will describe our sample efficient method for estimating failure probabilities in Section 3, after a quick review of preliminaries.
+With an expensive computer model, we simply cannot evaluate $f(\mathbf{x})$ thousands/millions/billions of times.
+Throughout, let $N$ denote the total budget of expensive simulator evaluations alotted.
+We will describe our sample efficient method for estimating failure probabilities in Section 3.2, targeting $N$ in the hundreds, after a quick review of preliminaries.
 2.1 Gaussian process surrogates
-A surrogate model $\hat{f}$, trained on a limited budget of runs, may provide predictions at unobserved inputs in place of the true $f$.
-To enable our proposed methodology the surrogate must also provide posterior predictive distributions.
-Gaussian processes (GPs) are the predominant choice; they assume a multivariate normal prior distribution over the response, $y\sim\mathcal{N}(\mu,\Sigma(X))$.
-The prior mean $\mu$ is often simplified to the zero vector (which we will use moving forward), and the prior covariance $\Sigma(X)$ is typically a function of inverse euclidean distances, e.g.
-$\Sigma(X)^{ij}=\Sigma(\mathbf{x}_{i},\mathbf{x}_{j})=k(||\mathbf{x}_{i}-\mathbf{x}_{j}||^{2})$ with kernel $k(\cdot)$.
+A surrogate model $\hat{f}$, trained on a limited budget of simulations, may provide predictions at unobserved inputs in place of the true $f$.
+To enable our proposed methodology, the surrogate must also provide posterior predictive distributions.
+Gaussian processes (GPs) are the predominant choice; they assume a multivariate normal distribution over the response.
+For any $n$ observed locations, our GP prior assumption is $\mathbf{y}_{n}\sim\mathcal{N}(\mu,\Sigma(X_{n}))$.
+The prior mean $\mu$ is often simplified to the zero vector after centering (which we will use moving forward), and the prior covariance $\Sigma(X_{n})$ is typically a function of inverse euclidean distances, e.g., $\Sigma(X_{n})^{ij}=\Sigma(\mathbf{x}_{i},\mathbf{x}_{j})=k(||\mathbf{x}_{i}-\mathbf{x}_{j}||^{2})$ with kernel $k(\cdot)$.
 See Santner et al., (2018); Rasmussen et al., (2006); Gramacy, (2020) for thorough reviews.
-Conditioned on training data $\left\{X_{n},\mathbf{y}_{n}\right\}$, posterior predictions at singeton input $\mathbf{x}$ follow
-$\hat{f}_{n}(\mathbf{x})\sim\mathcal{N}\left(\mu_{n}(\mathbf{x}),\sigma_{n}^{2}(\mathbf{x})\right)\quad\textrm{for}\quad\begin{array}[]{rl}\mu_{n}(\mathbf{x})&=\Sigma(\mathbf{x},X_{n})\Sigma(X_{n})^{-1}\mathbf{y}_{n}\\ \sigma_{n}^{2}(\mathbf{x})&=\Sigma(\mathbf{x})-\Sigma(\mathbf{x},X_{n})\Sigma(X_{n})^{-1}\Sigma(X_{n},\mathbf{x}),\end{array}$ (3)
+Conditioned on training data $\left\{X_{n},\mathbf{y}_{n}\right\}$, posterior predictions at singleton input $\mathbf{x}$ follow
+$\hat{f}_{n}(\mathbf{x})\sim\mathcal{N}\left(\mu_{n}(\mathbf{x}),\sigma_{n}^{2}(\mathbf{x})\right)\quad\textrm{for}\quad\begin{array}[]{rl}\mu_{n}(\mathbf{x})&=\Sigma(\mathbf{x},X_{n})\Sigma(X_{n})^{-1}\mathbf{y}_{n}\\ \sigma_{n}^{2}(\mathbf{x})&=\Sigma(\mathbf{x},\mathbf{x})-\Sigma(\mathbf{x},X_{n})\Sigma(X_{n})^{-1}\Sigma(\mathbf{x},X_{n})^{\top},\end{array}$ (3)
 where $\Sigma(\mathbf{x},X_{n})$ denotes the vector formed from applying the kernel between $\mathbf{x}$ and every row of $X_{n}$.
 Joint predictions are straightforward but not necessary for our contribution.
-A natural alternative to true MC (Eq. 2) is “surrogate MC” which leverages surrogate predictions in place of the true simulator,
-$\hat{\alpha}_{\textrm{SURR}}=\frac{1}{M}\sum_{i=1}^{M}\mathbb{I}_{\left\{\mu_{n}(\mathbf{x}_{i})>t\right\}}\quad\textrm{for}\quad\mathbf{x}_{i}\sim p(\mathbf{x}).$ (4)
-If the surrogate is not accurate near the contour, it will misclassify inputs which will severely hinder the effectiveness of this estimator.
-Figure 1: Left/Center: Herbie function with failure contour (red) and input distribution (black). Right: Entropy surface (yellow/high, purple/low) for a GP trained on a 100 point LHS (white circles).
-Figure 2: Slice of Herbie function along $x_{2}=1.1$. GP was trained on 100 point LHS in 2 dimensions.
+A natural alternative to true MC (2) is “surrogate MC” which leverages surrogate predictions in place of the true simulator,
+[…]
+Here we use the posterior mean of our GP in place of $f$. If the surrogate is not accurate near the contour, it will misclassify inputs which will severely hinder the effectiveness of this estimator.
+Figure 1: Left/Center: Herbie function with failure contour (red) and input distribution (white). Right: Entropy surface (yellow/high, purple/low) for a GP trained on a 100-point LHS (white circles).
+Figure 2: Slice of Herbie function along $x_{2}=1.1$. GP was trained on 100-point LHS in 2 dimensions.
 For example, consider the two-dimensional “Herbie” function (Lee et al.,, 2011) shown in the left and center panels of Figure 1 (we will revisit the right panel momentarily).
 We define a failure threshold of $t=1.065$ which demarcates four disjoint failure regions (red contours).
-We also define $p(\mathbf{x})$ using independent normal distributions centered at the origin (solid black contours).
+We also define $p(\mathbf{x})$ using independent normal distributions centered at the origin (white contours).
 [^1] Full details of all simulated functions, thresholds, and distributions are provided in the Supplementary Material.
-We trained a GP surrogate using a random Latin hypercube sample (LHS; McKay et al.,, 2000) of size $n=100$ in the two dimensions.
+We trained a GP surrogate using a random Latin hypercube sample (LHS; McKay et al.,, 2000) of size $n=100$ in two dimensions.
 Figure 2 shows the GP posterior prediction along the slice $x_{2}=1.1$.
-The posterior mean (blue dashed) and 95% CI (blue shading) capture the true nonlinear surface (solid black) fairly well.
+The posterior mean (dashed blue) and 95% confidence interval (blue shading) capture the true nonlinear surface (solid black) fairly well.
 Even so, there are inputs (highlighted by the yellow shading) which the surrogate would misclassify as “passes” instead of “failures.” Thus any estimate reliant solely on this surrogate would underestimate the failure probability.
@@
 In some situations, surrogate predictions may be improved through adaptations to the surrogate model itself.
-Traditional GPs are limited by the commonly assumed stationarity of the covariance kernel; they are not able to accurately model non-stationary surfaces characterized by regime shifts and stark transitions.
+Traditional GPs are limited by the commonly assumed stationarity of the covariance kernel; they are not able to accurately model nonstationary surfaces characterized by regime shifts and stark transitions.
 Deep Gaussian processes (Damianou and Lawrence,, 2013) are a promising more flexible alternative.
 They leverage functional compositions of GP layers in which latent layers act as warped versions of the original inputs, allowing for greater flexibility.
-DGPs have consistently outperformed stationary GPs in modeling non-stationary computer experiments (e.g., Rajaram et al.,, 2021; Yazdi,, 2022; Ming et al.,, 2023), particularly when deployed in a Bayesian framework with full UQ (e.g., Sauer et al.,, 2023b, a).
+DGPs have consistently outperformed stationary GPs in modeling nonstationary computer experiments (e.g., Rajaram et al.,, 2021; Yazdi,, 2022; Ming et al.,, 2023), particularly when deployed in a Bayesian framework using MCMC sampling to achieve full posterior integration of latent layers (e.g., Sauer et al.,, 2023b, a).
 See Sauer, (2023) for a thorough review of DGP surrogates.
 Crucially, our method is agnostic to surrogate choice as long as UQ is provided.
-We will demonstrate functionality with both traditional GP and Bayesian DGP surrogates in Section 3.2.
-Contour location.
-To facilitate failure probability estimation, a surrogate’s sole task it to accurately classify passes and failures.
+We will demonstrate functionality with both traditional GP and Bayesian DGP surrogates in Section 4.
+2.2 Contour location
+To facilitate failure probability estimation (4), a surrogate must accurately classify passes and failures.
 Strategic sequential designs targeting the failure contour, i.e., “contour location,” will outperform space-filling counterparts in this regard.
@@
 where $p_{\mathbf{x}}$ represents the probability of failure at input $\mathbf{x}$. With a GP surrogate, this probability boils down to a Gaussian CDF computation:
-$p_{\mathbf{x}}=\mathrm{P}\left(\hat{f}(\mathbf{x})>t\right)=1-\Phi\left(\frac{t-\mu_{n}(\mathbf{x})}{\sigma_{n}(\mathbf{x})}\right).$
+$p_{\mathbf{x}}=\mathrm{P}\left(\hat{f}_{n}(\mathbf{x})>t\right)=1-\Phi\left(\frac{t-\mu_{n}(\mathbf{x})}{\sigma_{n}(\mathbf{x})}\right).$
 Returning to the Herbie example, the entropy surface for the trained GP is shown in the right panel of Figure 1 and along the slice in Figure 2 (green dotted/dashed).
@@
 A reasonable acquisition could fall in any of these local optima.
-Booth et al., (2024b) extended these ideas, with some necessary adaptations, to enable contour location with Bayesian DGP surrogates.
-Henceforth, all GPs and DGPs we consider will be from sequential CL designs.
+Naturally, implementations of entropy as an aqcuisition function vary on the finer details.
+For GP surrogates we will leverage the optimization scheme of Cole et al., (2023), which first deploys a small LHS, identifies the LHS sample with the highest entropy, then runs a numerical optimization of entropy initialized at that point.
+By avoiding multi-starts, Cole et al., target local optima in the entropy surface.
+For DGP surrogates, we will follow Booth et al., (2024b).
+To avoid cumbersome numerical optimizations with the heftier MCMC-based DGP, Booth et al., use triangulation candidates (Gramacy et al.,, 2022), then select the candidate on the Pareto front of entropy and uncertainty.
+Henceforth, all GPs and DGPs we consider will be trained with these two contour location methods, respectively.
+Figure 3: Various designs for estimating the failure probability of the Herbie function with a GP surrogate and a total budget of $N=150$.
+All methods start with the same CL design of size $n=100$ (black circles), leaving a remaining budget of $B=50$ (blue triangles) for either continued surrogate training or importance sampling.
+Left: Entire budget is spent on continued contour location, with acquisitions exploring the entire space.
+Center: Remaining budget is sampled from a bias distribution for importance sampling (black circles omitted since they do not feature in the SIIS estimate).
+Right: Remaining budget is used for MC samples with highest entropy, solely exploiting near the predicted contour.
+Comparison of failure probability estimates obtained from these is provided later in Section 4.
+Although there is no clear distinction in performace between the left/right designs for this simple illustrative example, the continued exploration of the “All CL” design will hinder performance on larger, harder problems (Figure 6).
+Continuing the Herbie function example, the left panel of Figure 3 shows the design points chosen from entropy-based contour location with a GP surrogate (Cole et al.,, 2023), indicated by both the black circles and blue triangles (the distinction between the two will become clear when we discuss the other panels).
+Notice, the chosen points attempt to balance exploration (points spread out over the entire domain) and exploitation (points hovering around the red failure contour).
+The resulting surrogate may inform an $\hat{\alpha}_{\textrm{SURR}}$ estimate (4); we will include this as a competitor in our later exercises.
 2.3 Importance sampling
@@
 Monte Carlo approximation of this integral provides the estimator
-$\hat{\alpha}_{\mathrm{IS}}=\frac{1}{B}\sum_{i=1}^{B}w_{i}\;\mathbb{I}_{\left\{f(\mathbf{x}_{i})>t\right\}}\quad\textrm{where}\quad w_{i}=\frac{p(\mathbf{x}_{i})}{q(\mathbf{x}_{i})}\quad\textrm{for}\quad\mathbf{x}_{i}\sim q(\mathbf{x}).$ (6)
+[…]
 The “bias weights” $w_{i}$ account for the difference between the input density and the bias density.
@@
 But estimation of an effective bias distribution is tricky.
-The ideal bias distribution is $q(\mathbf{x})^{*}=\frac{1}{\alpha}\mathbb{I}_{\left\{f(\mathbf{x})>t\right\}}p(\mathbf{x})$, but naturally neither $\alpha$ nor the indicator function are known a priori.
+The ideal bias distribution is $q(\mathbf{x})^{*}=\frac{1}{\alpha}\mathbbm{1}_{\left\{f(\mathbf{x})>t\right\}}p(\mathbf{x})$, but neither $\alpha$ nor the indicator function are known a priori.
 For the sake of feasibility, $q(\mathbf{x})$ is traditionally constrained to a known distributional family, such as a Gaussian mixture model (GMM; Reynolds et al.,, 2009).
 Bias distribution training is then relegated to learning distributional parameters (in a GMM, these include the number of mixtures, mixture weights, and component means and variances), often accomplished through cross validation or information criteria maximization.
-In “surrogate-informed importance sampling,” Peherstorfer et al., (2016) leverage a surrogate to identify predicted failures from a large sample of the input distribution, then train a GMM bias distribution to these predicted failures.
+In “surrogate-informed importance sampling,” Peherstorfer et al., (2016) leverage a surrogate to identify predicted failures from a large Monte Carlo sample of the input distribution, then train a GMM bias distribution to these predicted failures.
 If the surrogate effectively identifies the failure region, then these samples should mimic the ideal $q(\mathbf{x})^{*}$.
-With a trained bias distribution in-hand, samples may be drawn from $q(\mathbf{x})$ and deployed in Eq.
+With a trained bias distribution in-hand, samples may be drawn from $q(\mathbf{x})$, evaluated through the expensive simulator, and deployed in Eq.
 (6) to provide an estimate of the failure probability.
+To illustrate, consider our running Herbie function example.
+Using the GP surrogate trained on the 100-point contour locating design from Figure 3 (black circles) and a large MC sample from the input distribution, we trained a GMM bias distribution to the MC samples which the surrogate predicted to be failures.
+The blue triangles in the center panel show 50 points sampled from this bias distribution.
+These samples may be evaluated through the true model to estimate $\hat{\alpha}_{\textrm{IS}}$ (6), thereby exhausting our entire budget of 150 evaluations.
+While the blue triangles target the failure region, they do not provide enough coverage to ensure a precise estimate.
+In our later experiments (Section 4), SIIS performance is highly variable.
 While SIIS has been widely embraced for its unbiasedness, it has several key weaknesses.
@@
 Furthermore, there is no clear avenue for determining the budget allocation between surrogate training and importance sampling estimation.
-Finally, the SIIS method does not leverage the full potential of the surrogate model.
-Many expensive simulator evaluations go into training the surrogate, and in the end the surrogate is only used to identify predicted failures, with no mention of its UQ.
-We suspect this is a byproduct of Peherstorfer et al., using reduced order models, which do not provide the UQ of a GP, in their seminal work.
-Ultimately, we propose a flexible estimation procedure which avoids bias distributions altogether, leverages surrogate UQ, and enables strategic budget allocation.
-3 Hybrid Monte Carlo with GPs
-In this section, we introduce a failure probability estimation scheme which (i) relies only on the surrogate model and the simulator data, (ii) leverages surrogate UQ in the estimation phase, and (iii) can provide effective estimates with a budget of only several hundred.
-Specifically, we will upgrade the Monte Carlo estimate of Eq.
-(2) so it requires fewer evaluations of $f(\mathbf{x})$ with minimal reduction in accuracy.
-Our approach (and those of our competitors) may be organized into two sequential phases: the training of the surrogate and the estimation of the failure probability.
-Let $N$ denote the number of simulator evaluations used to train the surrogate, and let $B$ denote the remaining budget of simulator evaluations available for estimation.
-For the duration of this section, we assume both $N$ and $B$ are pre-determined; in Section 4 we will consider how to allocate evaluations between them.
-Recall, we are targeting expensive simulations where the total budget, $N+B$, is constrained to be in the hundreds.
-3.1 Methodology
+[Throughout, we will borrow from our proposed method to inform budget allocation for the SIIS competitor, which is how we chose to separate the budget of Figure 3 into 100 for training and 50 for estimation.] Finally, the SIIS method does not leverage the full potential of the surrogate model.
+Many expensive simulator evaluations go into training the surrogate and, in the end, the surrogate is only used to identify predicted failures, with no mention of its UQ.
+[Notice, as portrayed in the center panel of Figure 3, the black circles are abandoned after the bias distribution is trained.]
+3 Two-stage design for failure probability estimation
+All in all, we are unsatisfied with the two state-of-the-art surrogate-based methods for estimating $\alpha$ which we just described: surrogate MC and SIIS.
+While the drawbacks to SIIS are apparent, the limitations of surrogate MC deserve more attention.
+To investigate further, we conducted the same CL sequential design of the Herbie function (Figure 3), this time storing the surrogate MC estimate $\hat{\alpha}_{\textrm{SURR}}$ (4), which we now denote as simply $\hat{\alpha}$, after every 10th acquisition.
+Progress is shown in the left panel of Figure 4.
+While there is significant learning occurring with early acquisitions, it quickly plateaus.
+By the time the surrogate has been trained on 100 points, it knows just as much about $\alpha$ as it does when it has been trained on the full budget of $N=150$ points.
+In this case, there is bound to be a more effective way to use the remaining simulation budget.
+We propose a solution: first developing a stopping criterion to decide when to halt contour location, then proposing a hybrid Monte Carlo estimator that greedily leverages the remaining simulation budget precisely for failure probability estimation.
+Figure 4: Left: Surrogate MC failure probability estimate for the Herbie function, updated after every 10th contour locating acquisition.
+True $\alpha$ in dashed red.
+Right: Difference in successive estimates with ±$\pm\hat{\sigma}_{\alpha}$ shown in dashed green.
+Blue square marks the second update within ±$\pm\hat{\sigma}_{\alpha}$.
+3.1 A stopping criterion for contour location
+Let $n$ denote the budget used on contour location and $B$ denote the remaining budget such that the total budget is $N=n+B$.
+We seek an objective, automatable decision rule for selecting $n$.
+The right panel of Figure 4 shows the relative difference between each successive $\hat{\alpha}$ estimate from the left panel.
+As expected, the relative differences converge to zero as the surrogate hones in on the true contour, but they never settle at exactly zero.
+How small of an update then is small enough to indicate learning has plateaued?
+The significance of an update to $\hat{\alpha}$ is only interpretable within the context of $\alpha$, the true failure probability, and $M$, the size of the MC sample used.
+Small differences might be inconsequential for large $\alpha$ with small $M$, but could be significant for small $\alpha$ with large $M$.
+To this end, we propose the use of the standard error $\sigma_{\alpha}$ to quantify the relative magnitude of each update.
+Notice in the estimator of Eq.
+(2), the indicator $\mathbbm{1}_{\left\{f(\mathbf{x})>t\right\}}$ is a Bernoulli random variable with expected value $\alpha$ and variance $\alpha(1-\alpha)$.
+Thus, the MC estimator has variance inversely proportional to $M$:
+$\mathbb{V}[\hat{\alpha}_{\textrm{MC}}]=\frac{1}{M^{2}}\sum_{i=1}^{M}\mathbb{V}\left[\mathbbm{1}_{\left\{f(\mathbf{x})>t\right\}}\right]=\frac{\alpha(1-\alpha)}{M}.$
+Let $\sigma_{\alpha}=\sqrt{\frac{\alpha(1-\alpha)}{M}}$ denote the standard error of this estimator.
+This standard error intrinsically accounts for the magnitude of $\alpha$ and $M$.
+Since the true $\alpha$ is unknown, we estimate $\hat{\sigma}_{\alpha}=\sqrt{\frac{\hat{\alpha}(1-\hat{\alpha})}{M}}$ using the current surrogate MC estimate.
+The green dashed lines in the right panel of Figure 4 show ±$\pm\hat{\sigma}_{\alpha}$, updated along with $\hat{\alpha}$.
+At the start, updates to $\hat{\alpha}$ are far beyond one standard error.
+But after a while, the updates sit comfortably within these bounds.
+Let $X_{M}$ denote the row-combined matrix of $M$-many samples from $p(\mathbf{x})$.
+We posit that, for fixed $X_{M}$, any update to $\hat{\alpha}$ within ±$\pm\hat{\sigma}$ is “small enough.” Accordingly, we propose halting CL once we have observed two successive updates within ±$\pm\hat{\sigma}$.
+The blue square in Figure 4 marks this point, stopping CL for this example at $n=100$, effectively identifying a point where surrogate learning has saturated and leaving a remaining budget of $B=50$ (this informed the budget allocation of Figure 3).
+Note, this stopping criterion requires the same $X_{M}$ be used in every $\hat{\alpha}$ calculation.
+If a new $X_{M}$ is generated for each estimate, there will be additional variability among the $\hat{\alpha}$’s, and the proposed ±$\pm\hat{\sigma}_{\alpha}$ threshold would need to be increased.
+Removing this extraneous source of variability allows successive estimates to pinpoint differences stemming solely from the surrogate.
+Our budgeting scheme is summarized as follows.
+Initialize a surrogate, begin a contour locating sequential design, and collect one sample $X_{M}$.
+At the start, it is likely that there will be no failures observed and that $\hat{\alpha}$ will be equal to zero (as seen in Figure 4).
+To safeguard against halting surrogate training prematurely, we build in a “common sense” check by setting a minimum number of failures that must be observed before the contour location could be halted.
+If the contour location design is succeeding, it will be placing some points in the failure region.
+We recommend a default of 10 failures, but suggest increasing this for more complex, higher dimensional problems.
+Once the design has surpassed these minimum requirements, estimate $\hat{\alpha}$ after every 10th acquisition.
+Calculate $\hat{\sigma}_{\alpha}$ and check if $|\hat{\alpha}^{\textrm{(current)}}-\hat{\alpha}^{\textrm{(previous)}}|<\hat{\sigma}_{\alpha}$.
+When this condition has been satisfied twice in a row, halt the sequential design and proceed to the next stage (the topic of Section 3.2).
+Choosing M.
+In practice, it is important to select a MC sample size that is large enough without being too cumbersome.
+Larger $M$ will produce more effective sum approximations of the true integrals, but we need surrogate predictions (e.g., Eq.
+3) for every sample in $X_{M}$.
+Although this computation is significantly cheaper than evaluation of the expensive simulator, it may still be significant for extremely large $M$, particularly with a heftier surrogate like a Bayesian DGP.
+If we knew $\alpha$, we could determine $M$ by thresholding the standard error $\sigma_{\alpha}$.
+For example, to achieve a standard error less than 10% of $\alpha$ requires:
+$\sigma_{\alpha}<\frac{\alpha}{10}\quad\longrightarrow\quad\frac{100(1-\alpha)}{\alpha}<M.$
+To achieve this precision, $\alpha=10^{-4}$ requires $M$ nearly 1 million, $\alpha=10^{-5}$ requires $M$ nearly 10 million, and $\alpha=10^{-6}$ requires $M$ nearly 100 million.
+In practice, with unknown $\alpha$, we recommend either choosing $M$ conservatively given expert insight into the anticipated failure probability or performing a pilot study where $M$ is incremented say by one million at a time.
+Gather surrogate predictions along this incremental process, and use $\hat{\alpha}$ to inform $\hat{\sigma}_{\alpha}$.
+3.2 Hybrid Monte Carlo for failure probability estimation
+Once $n$ is determined, how should we spend the remaining budget of simulator evaluations ($B=N-n$), in conjunction with our trained surrogate, to most effectively estimate $\alpha$?
+We propose an upgrade to the surrogate MC estimator of Eq.
+(4).
 Let $\mathcal{U}\subset\mathcal{X}$ represent some subset of the input domain (boundary inclusive).
-We may partition the integral of Eq.
-(1) into two components as follows, with the corresponding Monte Carlo estimate,
+Then, Eq.
+(1) may be partitioned as
 […]
+inspiring an equivalent formulation of the surrogate MC estimator,
 […]
-Now, replace expensive evaluation of $f(\mathbf{x}_{i})$ outside of $\mathcal{U}$ with the surrogate prediction $\mu_{N}(\mathbf{x}_{i})$, resulting in the following “hybrid Monte Carlo” estimator:
+Now for $\mathbf{x}_{i}\in\mathcal{U}$, upgrade surrogate prediction $\mu_{n}(\mathbf{x}_{i})$ to expensive evaluation of the true $f(\mathbf{x}_{i})$, resulting in the following “hybrid MC” estimator:
 […]
 The hybrid nature of this estimator is in the use of both true simulator evaluations (left term) and surrogate evaluations (right term).
-Each $\mathbf{x}_{i}$ will fall under either the left indicator term or the right one, never both.
-The computational cost of this estimator is greatly reduced since evaluations of the surrogate are effectively free.
 The accuracy of this estimator is contingent upon the classification accuracy of the surrogate and the choice of $\mathcal{U}$.
-Here, we focus on the choice of $\mathcal{U}$, presuming we already have a trained GP-based surrogate in-hand.
-[Training a surrogate for contour location has been previously investigated, as we reviewed in Section 2.1.]
-Rather than thinking of $\mathcal{U}$ as a continuum, we find it easier to conceptualize our contribution with a discrete set of samples, which form the basis of any Monte Carlo estimator.
-Let $X_{M}$ denote the row-combined matrix of $M$-many samples from $p(\mathbf{x})$.
-To deploy our estimator $\hat{\alpha}_{\textrm{HYB}}$, we need a decision rule for classifying each $\mathbf{x}_{i}$ into or out of $\mathcal{U}$.
+As long as we restrict the size of $\mathcal{U}$ so only $B$-many samples from $X_{M}$ fall within it, i.e., $\sum_{i=1}^{M}\mathbbm{1}_{\left\{\mathbf{x}_{i}\in\mathcal{U}\right\}}=B$, then we will meet our budgetary constraints.
+Rather than thinking of $\mathcal{U}$ as a continuum, we find it easier to conceptualize in the context of the discrete set $X_{M}$.
 In a perfect world, $\mathcal{U}$ would capture all the samples which the surrogate would misclassify.
-We are not privy to this information, but we can leverage the surrogate’s UQ to identify the samples at which it is most uncertain.
-Recall, entropy (Eq.
-5) quantifies the degree of uncertainty in the pass/fail prediction of a surrogate.
-Accordingly, our decision rule is strikingly simple – place the $B$-many samples from $X_{M}$ with the highest entropy into $\mathcal{U}$.
-Denote these as $X_{M}^{\mathcal{U}}$.
-Hopefully our choice of the letter $\mathcal{U}$ is now apparent - it denotes the inputs with the highest “$\mathcal{U}$ncertainty”.
-Revisiting Figure 2, notice the high entropy regions (green dotted/dashed) align well with the misclassification regions (yellow).
-The entropy criterion is effectively identifying regions of high classification uncertainty where we could most utilize expensive true simulator evaluations.
-Li and Xiu, (2010) employed a hybrid estimator similar to that of Eq.
-(7), but they chose to allocate expensive simulator evaluations to samples predicted to be closest to the contour.
-We suspect this choice was motivated by their use of polynomial chaos expansion models as surrogates which do not facilitate UQ.
-Nevertheless, we will implement this allocation scheme with our hybrid Monte Carlo estimator as a competitor in our benchmark exercises later.
-For a fixed $N$, $B$, and $M$, our method proceeds as follows.
-First, train the surrogate using $N$-many evaluations of the high fidelity simulator.
-Our method is agnostic to surrogate and design choice as long as the surrogate provides posterior predictive distributions for the calculation of entropy, but we prefer contour locating designs with GP or DGP surrogates.
-Second, sample a very large number of points $X_{M}\sim p(\mathbf{x})$ and use the surrogate to calculate the classification entropy for each sample following Eq.
+We are not privy to this information, and we are under a limited budget, but we can leverage the surrogate’s UQ to identify the $B$-many samples at which it is most uncertain.
+Recall, classification entropy (5) quantifies the degree of uncertainty in the pass/fail prediction of a surrogate.
+Consequently, our proposed allocation of samples to $\mathcal{U}$ is strikingly simple – place the $B$-many samples from $X_{M}$ with the highest entropy into $\mathcal{U}$.
+Denote these as $X_{M}^{B}$.
+Revisiting Figure 2, notice the high entropy regions (green dotted/dashed) align well with the misclassification regions (yellow shading).
+The entropy criterion is effectively identifying regions of high classification uncertainty where we could most utilize expensive true simulator evaluations to correct inaccurate surrogate predictions.
+Ultimately, after evaluating $f(\mathbf{x})$ for $\mathbf{x}\in X_{M}^{B}$, we update the surrogate with these additional observations, thereby updating $\mu_{n}(\mathbf{x})$, before returning $\hat{\alpha}_{\textrm{HYB}}$ (7).
+Since the $X_{M}^{B}$ points are likely near a failure contour, they could significantly improve the accuracy of the surrogate predictions for other samples near the contour.
+For reference, consider the right panel of Figure 3.
+The black circles showing the contour locating design for the GP surrogate are unchanged, but here the additional budget of 50 evaluations is spent on the MC samples from $p(\mathbf{x})$ with the highest classification entropy (blue triangles).
+Compared to the left panel which continues to explore the entire domain, these highest entropy samples exclusively exploit the predicted failure contour.
+Compared to the bias samples portrayed in the middle panel, where the original surrogate design is abandoned, our two-stage design is able to leverage all simulator evaluations in its final estimate.
+3.3 Summary and related work
+We have presented a comprehensive strategy for estimating Eq.
+(1) from a budget of size $N$.
+It is summarized as follows.
+Start with a very large Monte Carlo sample from $p(\mathbf{x})$, denoted $X_{M}$.
+Train a surrogate model through a contour locating sequential design, monitoring the progress of the $\hat{\alpha}_{\textrm{SURR}}$ estimate for given $X_{M}$, and halting contour location once two consecutive updates are within ±$\pm\hat{\sigma}_{\alpha}$.
+Denote the design points chosen in this stage as $X_{n}$.
+Then, use the trained surrogate to identify the highest entropy samples, $X_{M}^{B}$, evaluate the simulator at these inputs, and update the surrogate with these new observations.
+Finally, return $\hat{\alpha}_{\textrm{HYB}}$ following Eq.
 (7).
-Third, identify the $B$-many points in $X_{M}$ with the highest entropy and evaluate the simulator at these locations.
-Fourth, update the surrogate with these new observations of the simulator.
-Finally, estimate $\hat{\alpha}_{\textrm{HYB}}$ following Eq.
-(7) by combining the simulator evaluations at $X_{M}^{\mathcal{U}}$ with surrogate predictions at all other samples of $X_{M}$.
-This fourth step – updating the surrogate given the simulator data at the high-entropy sampled locations – is significant.
-In the end, the surrogate is trained on $N+B$ observations, making full use of all the available data.
-Since the $X_{M}^{\mathcal{U}}$ points are likely near a failure contour, they could significanly improve the accuracy of the surrogate near the contour.
-The ability to update the surrogate and make full use of all simulator data is a unique advantage over SIIS, whose final estimate is contingent only upon the $B$-many points sampled from the bias distribution.
-There are some noteworthy distinctions between the training and estimation phases of our method.
-The $N$ points used to initially train the surrogate are acquired anywhere in the domain $\mathcal{X}$, and assuming they stem from a contour locating design, they seek to balance both exploration of the response surface and exploitation of the suspected contour.
-In fact, the contour locating sequential design strategies we will use next (Cole et al.,, 2023; Booth et al.,, 2024b) intentionally circumvent acquisitions based on maximal entropy in order to avoid clustered acquisitions and to promote more exploration of the surface.
-Alternatively, the $B$ points used for the hybridization and final surrogate update must originate from $p(\mathbf{x})$, and they are solely exploitative.
+Given the deterministic nature of $f(\mathbf{x})$, effective surrogates should interpolate observed points.
+In this light, the hybrid MC estimator (7) is akin to a surrogate MC estimator (4) when the surrogate has been trained on all $\mathbf{x}_{i}\in\mathcal{U}$.
+This connection permits us to frame our contribution as a two-stage design.
+In the first stage, $X_{n}$ is chosen over $\mathcal{X}$ based on contour location.
+In the second stage, $X_{M}^{B}$ is chosen over $\mathbbm{P}_{\mathbf{x}}$ based on maximum entropy.
+The final design of size $N$ is then $X_{N}=[X_{n}^{\top},{X_{M}^{B}}^{\top}]^{\top}$.
+There are some noteworthy distinctions between these two design stages.
+To illustrate, refer back to the black circles ($X_{n}$) and blue triangles ($X_{M}^{B}$) in the right panel of Figure 3.
+In the first stage, the $n$ points selected for contour location may be acquired anywhere in the domain $\mathcal{X}$, seeking to balance both exploration of the response surface and exploitation of the suspected contour.
+In fact, the contour locating sequential design strategies we consider in Section 4 (Cole et al.,, 2023; Booth et al.,, 2024b) intentionally circumvent acquisitions based on maximal entropy in order to avoid clustered acquisitions and to promote more exploration of the surface.
+Alternatively, the $B$ points chosen in the second stage must originate from $p(\mathbf{x})$ and are solely exploitative.
 By nature, they will be clustered around the predicted contour.
 We contend that having a mix of data from these two realms is superior to either extreme.
-To demonstrate superiority over the first extreme, in which the entire budget is relegated to training the surrogate over $\mathcal{X}$ ignorant of $p(\mathbf{x})$, we will offer empirical comparisons in our simulated examples.
-To justify this claim over the second extreme, in which all training data must stem from $p(\mathbf{x})$, we revisit the Herbie example of Figure 1.
+Some of the works we referenced in Section 1 have investigated stopping criteria for surrogate training in reliability settings.
+The most similar to our proposed criterion is that of Zhu and Du, (2016) who recommend a convergence criterion akin to $\frac{\hat{\sigma}_{\alpha}}{\hat{\alpha}}\leq\epsilon$, where $\epsilon$ is a user specified threshold.
+While similar to our stopping criterion, the choice of $\epsilon$ here is make-or-break, and little guidance is provided.
+If $\epsilon$ is too high, contour location will be halted too soon; if too low, contour location will continue past the point of utility.
+Our proposed stopping criterion avoids this choice and may be easily automated without problem-specific tuning.
+In our synthetic experiments, we find that it consistently results in appropriate termination with budget to spare.
+Hybrid MC estimators (similar to Eq.
+7) have been entertained in other settings.
+Li and Xiu, (2010) use polynomial chaos models for a hybrid MC estimator with samples delegated to $\mathcal{U}$ based on predicted proximity to the contour, i.e., $|\mu_{n}(\mathbf{x}_{i})-t|$.
+This work, however, does not consider surrogate design and does not leverage surrogate UQ.
+Fuhrländer and Schöps, (2020) use hybrid MC, but they only consider training the surrogate on samples from $p(\mathbf{x})$.
+Echard et al., (2011) similarly train only on a subset of $X_{M}$.
+We attempted to recreate the approach of Fuhrländer and Schöps, (2020) in our setting: (i) sample large $X_{M}$, (ii) train an initial surrogate on a random selection of these samples, (iii) use the surrogate with a metric such as entropy to select additional batches from $X_{M}$ for evaluation through the simulator, (iv) update the surrogate, and (v) repeat steps (iii)-(iv) until the budget is exhausted.
+Yet we were unable to obtain even reasonable results because initial surrogate fits were very poor, and surrogate-informed acquisitions were unable to recover.
+Random selections from $X_{M}$ were overwhelmingly in high probability regions, not low probability failure regions.
+Hristov et al., (2019) noticed this conundrum and stated it well: “using a surrogate that is built on frequent events to predict and analyze rare events” can be futile.
+To visualize this conundrum in a simple setting, we revisit the Herbie example of Figure 1.
 With the input distribution centered at the origin, the probability of generating any samples outside $[-1,1]^{2}$ is very low.
@@
 This distinction will be less prominent if $p(\mathbf{x})$ is uniformly distributed over $\mathcal{X}$, but the trade-off between exploring in the training phase and exploiting in the estimation phase will still be impactful.
-Choosing M.
-Larger $M$ will produce more effective sum approximations of the true integrals, but we need surrogate predictions (e.g., Eq.
-3) for every sample in $X_{M}$.
-Although this computation is significantly cheaper than evaluation of the expensive simulator, it may still be significant for large $M$, particularly with a heftier surrogate like a Bayesian DGP.
-Simply put, we would like $M$ to be large enough without being too cumbersome.
-To reason through this choice, let’s revisit the unbiased Monte Carlo estimate of Eq.
-(2).
-Notice the indicator $\mathbb{I}_{\left\{f(\mathbf{x})>t\right\}}$ is a Bernoulli random variable with expected value $\alpha$ and variance $\alpha(1-\alpha)$.
-Thus, the MC estimator has variance inversely proportional to $M$:
-$\mathbb{V}[\hat{\alpha}_{\textrm{MC}}]=\frac{1}{M^{2}}\sum_{i=1}^{M}\mathbb{V}\left[\mathbb{I}_{\left\{f(\mathbf{x})>t\right\}}\right]=\frac{\alpha(1-\alpha)}{M}.$
-Let $\sigma_{\alpha}=\sqrt{\frac{\alpha(1-\alpha)}{M}}$ denote the standard error of this estimator.
-If we knew $\alpha$, we could determine $M$ by thresholding this standard error.
-For example, to achieve a standard error less than 10% of $\alpha$ requires:
-$\sigma_{\alpha}<\frac{\alpha}{10}\quad\longrightarrow\quad\frac{100(1-\alpha)}{\alpha}<M.$
-To achieve this precision, $\alpha=10^{-4}$ requires $M$ nearly 1 million, $\alpha=10^{-5}$ requires $M$ nearly 10 million, and $\alpha=10^{-6}$ requires $M$ nearly 100 million.
-In reality, when the true $\alpha$ is unknown, we recommend either choosing $M$ conservatively given expert insight into the anticipated failure probability or performing a pilot study where $M$ is incremented say by one million at a time.
-Surrogate predictions are gathered along this incremental process, with $\alpha$ estimated along the way to inform the selection of $M$.
-3.2 Synthetic experiments
-In this section we provide a variety of synthetic experiments in which the training and estimation budgets, $N$ and $B$ respectively, are pre-determined.
-We start with three functions: the 2d Herbie function (Lee et al.,, 2011) with a total budget of 160, the 3d Ishigami function (Ishigami and Homma,, 1990) with a total budget of 400, and the 6d Hartmann function (Picheny et al.,, 2013) with a total budget of 800.
-We split the total budget of evaluations evenly between training and estimation.
-Table 1 summarizes these settings, among others.
+Even with uniform $p(\mathbf{x})$, training a surrogate only on maximal entropy points is suboptimal (Cole et al.,, 2023; Booth et al.,, 2024b).
+4 Synthetic experiments
+In this section we benchmark our method against state-of-the-art competitors on a variety of synthetic experiments.
+We consider four functions: the 2d Herbie function (Lee et al.,, 2011) with a total budget of 150, the 3d Ishigami function (Ishigami and Homma,, 1990) with a total budget of 300, the 6d Hartmann function (Picheny et al.,, 2013) with a total budget of 600, and the 4d Plateau function (Booth et al.,, 2024b) with a total budget of 200.
+The first three of these are plausibly stationary, so we use a standard GP surrogate, fit with the Scikit-learn package in python (Pedregosa et al.,, 2011), and the entropy contour location scheme of Cole et al., (2023).
+The Plateau function is known to be nonstationary; it is characterized by flat regions with a steep sloping drop between them.
+To address the nonstationarity, we use a Bayesian DGP surrogate (Sauer et al.,, 2023b), fit with the deepgp package in R (Booth,, 2023), with the contour location scheme of Booth et al., (2024b).
+With the right surrogate, the failure contour of the Plateau function can be well estimated from a rather small amount of data.
+Function True $\alpha$ $d$ $n_{0}$ $N$ $M$ Surrogate CL Scheme
+Herbie $7.533\times 10^{-5}$ 2 20 150 $3.5\times 10^{7}$ GP Cole et al., (2023)
+Ishigami $1.904\times 10^{-4}$ 3 50 300 $1.5\times 10^{7}$ GP Cole et al., (2023)
+Hartmann $1.001\times 10^{-5}$ 6 100 600 $1.0\times 10^{8}$ GP Cole et al., (2023)
+Plateau $4.308\times 10^{-4}$ 4 30 200 $3.5\times 10^{6}$ DGP Booth et al., (2024b)
+Table 1: Simulation settings. True $\alpha$ values are estimated from 10 billion samples.
+Table 1 summarizes the simulation settings.
 Functions are observed without noise to mimic a deterministic computer simulation.
 We define failure thresholds and input distributions for each function, resulting in the failure probabilities reported in Table 1.
-We choose $M$ based on these $\alpha$ values to keep $\sigma_{\alpha}$ conservatively under 5%.
 All input distributions are variations/combinations of independent uniform and truncated normal distributions; further details of the functions, thresholds, and distributions are reserved for Supplementary Material.
-Function True $\alpha$ Surrogate $d$ $n_{0}$ $N$ $B$ $M$
-Herbie $7.533\times 10^{-5}$ GP 2 20 80 80 $3.5\times 10^{7}$
-Ishigami $1.904\times 10^{-4}$ GP 3 50 200 200 $1.5\times 10^{7}$
-Hartmann $1.001\times 10^{-5}$ GP 6 100 400 400 $1.0\times 10^{8}$
-Plateau $8.530\times 10^{-4}$ DGP 4 30 120 120 $3.5\times 10^{7}$
-Table 1: Simulaton settings. True $\alpha$ values are estimated from 10 billion samples.
-We begin with a random Latin hypercube sample (McKay et al.,, 2000) of size $n_{0}$ (chosen to be 25% of the total training budget, see Table 1).
-We train a GP surrogate using the Scikit-learn package in python (Pedregosa et al.,, 2011).
-We then conduct a sequential design following the entropy contour location scheme of Cole et al., (2023), until we reach a final design of size $N$.
-Ultimately, we deploy the trained surrogate, in conjunction with a large sample $X_{M}\sim p(\mathbf{x})$, to estimate $\alpha$ with each of the following methods:
-  • hybMC ENT: Our hybrid MC estimate (Eq.
-7) where $\mathcal{U}$ is chosen to contain the $B$ points from $X_{M}$ with highest entropy.
-[For now, ignore *hybMC ENT, we will return to it in Section 4.]
-  • GP MC $N$: Surrogate MC estimate (Eq. 4) using the surrogate trained on $N$ points.
-  • GP MC $N+B$: Surrogate MC estimate (Eq. 4) using a surrogate trained on $N+B$ points.
-  • hybMC CLOSE: Inspired by Li and Xiu, (2010), the hybrid MC estimate (Eq.
-7) where $\mathcal{U}$ is chosen to contain the $B$ points from $X_{M}$ which are predicted to be closest to the failure threshold (i.e., the samples with the smallest $|\mu_{N}(\mathbf{x}_{i})-t|$).
-  • SIIS: Importance sampling estimate (Eq.
-6; Peherstorfer et al.,, 2016) with a GMM bias distribution, also fit with the Scikit-learn package.
+We choose $M$ for each $\alpha$ to keep $\sigma_{\alpha}$ conservatively under 5%.
+Figure 5: Distribution of the chosen $n$ across 30 repetitions.
+First, we deploy our stopping criterion (Section 3.1) for each of the four scenarios, starting with an initial LHS sample of size $n_{0}$ (Table 1) and repeating for 30 re-randomized Monte Carlo repetitions.
+To avoid premature termination, we require a minimum of 10 observed failures and $n\geq 2n_{0}$ (atleast as many contour locating acquisitions as original space-filling observations).
+Figure 5 shows the distribution of chosen $n$ values.
+Notice, all trials terminated before the allotted $N$, reserving some evaluations for additional estimation, but there was considerable variability across trials.
+We view this variability as a strength – some random initializations need more data to effectively locate the contour while some get lucky from the start.
+Allowing the surrogate to dictate its own budget mitigates some of the risks of having an unlucky starting design.
+For each trained surrogate with chosen $n$ (which varies across repetitions), we collect a large Monte Carlo sample ($X_{M}$, randomized across repetitions), then deploy the following estimation methods:
+  • All CL: Ignoring the chosen $n$, continue contour location until the entire budget is exhausted and return the surrogate MC estimate (4).
+  • Hybrid Entropy: Our two-stage approach using the hybrid MC estimate (7) where $\mathcal{U}$ is chosen to contain the $B$ points from $X_{M}$ with highest entropy.
+  • Hybrid Proximity: Same as “Hybrid Entropy,” but $\mathcal{U}$ is chosen to contain the $B$ points from $X_{M}$ which are predicted to be closest to the failure threshold (i.e., the samples with the smallest $|\mu_{n}(\mathbf{x}_{i})-t|$).
+This is inspired by the work of Li and Xiu, (2010).
+  • SIIS: Surrogate-informed importance sampling (6) with a GMM bias distribution (Peherstorfer et al.,, 2016).
 The bias distribution is trained on the predicted failures, i.e., $\left\{\mathbf{x}_{i}\in X_{M}\;|\;\mu_{n}(\mathbf{x}_{i})>t\right\}$.
-  • SIIS UCB: Same as SIIS, but the bias distribution is trained on the samples whose upper 90% confidence bound exceeds the threshold, i.e., $\left\{\mathbf{x}_{i}\in X_{M}\;|\;\mu_{N}(\mathbf{x}_{i})+1.645\sigma_{N}(\mathbf{x}_{i})>t\right\}$.
-Note, the first four methods are variations of Monte Carlo, and the last two are variations of importance sampling.
-Crucially, each method uses the exact same trained surrogate (besides GP MC $N+B$, which is trained further) and the exact same $X_{M}$ sample.
-Differences in performance are thus attributed entirely to the estimation method.
-The GP MC $N$ method is the only one which does not utilize the entire budget; it serves as a snapshot of the surrogate’s accuracy before it is combined with a hybrid calculation or importance sampling.
-We repeat each exercise 30 times with re-randomized starting designs.
-To demonstrate functionality with a DGP, we additionally consider the 4d plateau function (Booth et al.,, 2024b).
-This function is known to be non-stationary; it is characterized by flat regions with a steep sloping drop between them.
-To address the complexity in the response surface, we use a Bayesian DGP surrogate (Sauer et al.,, 2023b) fit with the deepgp R-package (Booth,, 2023).
-We train each DGP using the contour location scheme of Booth et al., (2024b), which involves proposing triangulation candidates (Gramacy et al.,, 2022) and selecting the candidate on the Pareto front of entropy and uncertainty.
-To keep all of the code for this example in R for easy integration with the DGP, we leverage the mclust R-package (Scrucca et al.,, 2016) to train the GMM bias distribution.
-Given the increased complexity of the surrogate, we perform 20 re-randomized repetitions.
+For our GP surrogates implemented in python, we use the scikit-learn package (Pedregosa et al.,, 2011) to train GMM bias distributions.
+For our DGP surrogates implemented in R, we use the mclust package (Scrucca et al.,, 2016).
+  • SIIS UCB: Same as “SIIS,” but the bias distribution is trained on the samples whose upper 95% confidence bound exceeds the threshold, i.e., $\left\{\mathbf{x}_{i}\in X_{M}\;|\;\mu_{n}(\mathbf{x}_{i})+1.645\sigma_{n}(\mathbf{x}_{i})>t\right\}$.
+Each method uses the same $X_{M}$ and the same surrogate trained up to point $n$.
+The distinction is in what is done with the remaining $B$ samples.
+As implemented, the latter three are to some degree an internal comparison since they use our proposed stopping criterion to determine the budget allocation between surrogate training and estimation (Figure 5).
+Without borrowing from our proposed approach, there would be no guidance on the allocation of $n$ and $B$ for these methods.
+The “Hybrid Proximity” approach is truly an internal comparison as it borrows everything from our proposed two-stage design aside from the specific allocation of samples to $\mathcal{U}$.
+It is included to show the impact of the entropy selection criterion, which uses surrogate UQ, over a selection solely based on the predicted mean.
 Reproducible code for all experiments, including both the python and R implementations, is available in our public git repository.
 [^2] https://bitbucket.org/boothlab/failprob/
-Results are shown in Figure 3, with true failure probability indicated by the red dashed line.
-The grey shaded region highlights the interval $\{\alpha\pm 2\sigma_{\alpha}\}$ where $\sigma_{\alpha}$ is calculated from the provided $M$.
+Figure 6: Estimated failure probabilities shown over true failure probability (red dashed) ±$\pm 2$ standard errors (grey shading). Boxplots show 30 repetitions.
+Results are shown in Figure 6.
+Red dashed lines mark the true failure probabilities, and grey shaded regions highlight $\{\alpha\pm 2\sigma_{\alpha}\}$ where $\sigma_{\alpha}$ is calculated from the true $\alpha$ and chosen $M$.
 This region provides important context – it captures the intrinsic variability that would accompany any MC estimator.
-In other words, if we used the true simulator for all $\{\mathbf{x}_{i}\}_{i=1}^{M}$ samples (Eq.
-2), we would obtain a boxplot roughly spanning the grey interval.
+In other words, even if we used the true simulator for all $\{\mathbf{x}_{i}\}_{i=1}^{M}$ samples (2), we would obtain a boxplot roughly spanning the grey interval.
 We thus consider any results within this region to be effective estimates.
 We could tighten the interval (and the boxplots of the MC methods) by increasing $M$ without requiring any additional simulator evaluations, another advantage MC holds over IS.
-Figure 3: Estimated failure probabilities shown over true failure probability (red dashed) ±$\pm 2$ standard errors (grey shading). Boxplots show 30 repetitions (20 for Plateau).
-Our hybMC ENT method (teal) performed well across the board.
-Unsurprisingly, it was able to improve upon the accuracy of the original surrogate (purple) by leveraging the additional budget of $B$ evaluations.
-More importantly, for all but the simplest Herbie function, it outperformed the surrogate MC estimates which made use of the full budget (blue).
-Recall, the (D)GP MC methods acquire points from $\mathcal{X}$, while the hybMC method acquires points from $p(\mathbf{x})$ (after the original training phase).
-The fact that hybMC is able to outperform the fully-trained surrogate indicates that at some point, it is better to prioritize expensive simulator evaluations for uncertain samples from $p(\mathbf{x})$ than to continue training the surrogate.
-The strategy of Li and Xiu, (2010) in allotting simulator evaluations based on predicted distance to the contour (hybMC CLOSE, orange) did not significantly improve upon the accuracy of the surrogate alone.
-This idea fails to leverage the UQ of the surrogate.
-It is possible for the surrogate to have higher uncertainty in its pass/fail prediction for a point further from the contour.
-For example, refer back to Figure 2.
-The entropy surrounding $x_{1}=1$ is far greater than that surrounding $x_{1}=0.5$, even though the latter of these is predicted to be closer to the contour (blue dashed line).
-Reserving true simulator evaluation for the $x_{1}=1$ location, where the surrogate would misclassify a failure, is preferable.
-The SIIS methods (green and pink) had the poorest performance across the board.
-They particularly struggled in the simpler settings (Herbie and Ishigami), underperforming the simple GP MC $N$ method which only used half the amount of simulator evaluations.
-We suspect the small budgets are to blame here; IS just needs more data, particularly when the GMM bias density is far from the ideal.
-Consequently, these IS variations had their best performance with the Hartmann function where 400 samples were budgeted for estimation.
-There appears to be very little benefit to the more conservative SIIS UCB (pink) over the traditional SIIS (green).
-4 Allocating the simulation budget
-Let $T=N+B$ represent the total budget of simulator evaluations, which is typically determined by the computational cost of the simulator.
-For expensive simulators which require long compute times, $T$ may be only several hundred.
-We presume $T$ is fixed, and we now address how to split it between $N$ and $B$.
-In our previous studies, we evenly split $T$ between training the surrogate and estimating $\alpha$, i.e., $N=B=\frac{1}{2}T$.
-But this decision was admittedly arbitrary.
-In this section, we propose a scheme for allocating $T$ between $N$ and $B$, specifically for use within our hybrid Monte Carlo estimation scheme.
-Since the sum of $N$ and $B$ is fixed, this choice boils down to selecting $N$; $B=T-N$ then follows directly.
-Ideally, we would like to halt training of the surrogate once learning has plateaued, saving the rest of the budget for the hybrid Monte Carlo estimate.
-If we stop training too early, the surrogate might be too inaccurate for the hybrid MC estimate to overcome.
-If we continue training for too long, we will not have any samples left for the hybrid estimation procedure.
-As we saw in Figure 3, in all but the simplest case, a MC estimate without any hybridization is suboptimal.
-Figure 4: Left: Surrogate MC failure probabilty estimate for the Ishigami function, updated every 10th acquisition with true $\alpha$ in dashed red.
-Center/Right: Difference in successive estimates with ±$\pm\hat{\sigma}_{\alpha}$ shown in dashed green.
-The blue square marks the second update within ±$\pm\hat{\sigma}_{\alpha}$.
-To monitor the surrogate’s learning, we propose tracking the progress of the surrogate MC estimate, $\hat{\alpha}_{\textrm{SURR}}$ (Eq.
-4), which we will now denote as simply $\hat{\alpha}$.
-This estimator relies only on the trained surrogate and does not require additional simulator evaluations.
-As a demonstration, the left panel of Figure 4 reports the progress in $\hat{\alpha}$ for the 3d Ishigami function following the setup described in Section 3.2.
-To avoid excessive calculations, we evaluate the estimate after every 10th acquisision.
-It is apparent that the surrogate’s learning of the failure contour saturates well before it reaches $N=200$, which we used in our previous experiments.
-The center panel of Figure 4 shows the relative difference between each successive $\hat{\alpha}$.
-As expected, the relative differences converge to zero as the surrogate hones in on the true contour, but they never settle at exactly zero (this is more apparent in the right panel which is zoomed in).
-How small of an update then is small enough to indicate learning has plateaued?
-The significance of an update to $\hat{\alpha}$ is only interpretable within the context of $\alpha$ and $M$.
-Small differences might be inconsequential for large $\alpha$ with small $M$, but could be significant for small $\alpha$ with large $M$.
-To this end, we propose the use of the standard error $\sigma_{\alpha}$ to quantify the relative magnitude of each update.
-This standard error intrinsically accounts for the magnitude of $\alpha$ and $M$.
-Since the true $\alpha$ is unknown, we estimate $\hat{\sigma}_{\alpha}=\sqrt{\frac{\hat{\alpha}(1-\hat{\alpha})}{M}}$ using the current surrogate MC estimate.
-The green dashed lines in the center and right panels of Figure 5 show ±$\pm\hat{\sigma}_{\alpha}$ for the estimated failure probabilities.
-At the start, updates to $\hat{\alpha}$ are far beyond one standard error.
-But after a while, the updates sit comfortably within these bounds.
-We posit that, for a fixed sample $X_{M}$, any update to $\hat{\alpha}$ within ±$\pm\hat{\sigma}$ is “small enough.” Accordingly, we halt surrogate training once we have observed two successive updates within ±$\pm\hat{\sigma}$.
-Our budgeting scheme is summarized as follows.
-Initialize a surrogate, begin a contour locating sequential design, and collect one sample $X_{M}$.
-At the start, it is likely that there will be no failures observed and that $\hat{\alpha}$ will be equal to zero (as is the case in Figure 4).
-To safeguard against halting surrogate training prematurely, we build in a “common sense” check by setting a minimum number of failures that must be observed before the contour location could be halted.
-If the contour location design is succeeding, it will be placing some points in the failure region.
-We recommend a default of 10 failures, but suggest increasing this for higher dimensional problems.
-In our synthetic experiments, we used a minimum of 10 failures for the Herbie, Ishigami, and plateau functions but bumped this up to 40 for the larger dimension of the Hartmann function.
-Once the surrogate has surpassed these minimum requirements, estimate $\hat{\alpha}$ after every 10th acquisition.
-Calculate $\hat{\sigma}_{\alpha}$ and check if $|\hat{\alpha}^{\textrm{(current)}}-\hat{\alpha}^{\textrm{(previous)}}|<\hat{\sigma}_{\alpha}$.
-Once this condition has been satisfied twice in a row, halt the sequential design and proceed to the hybrid Monte Carlo estimation following Section 3.1.
-This stopping criterion requires the same $X_{M}$ sample be used in every $\hat{\alpha}$ calculation.
-If a new $X_{M}$ is generated for each estimate, there will naturally be more variability among the $\hat{\alpha}$’s, and the ±$\pm\hat{\sigma}_{\alpha}$ threshold would need to be increased.
-Removing this extraneous source of variability allows successive estimates to pinpoint differences stemming solely from the surrogate.
-Figure 5: Distribution of the chosen $N$ for each repetition of the synthetic examples from Figure 3. Red star indicates the value that was used previously.
-Revisiting synthetic experiments.
-We implemented this stopping criteria with each of the four functions described in Section 3.2.
-We used the same total budgets ($T=\{160,400,800,240\}$, respectively), but chose $N$ based on this decision rule.
-Once $N$ was determined, we set $B=T-N$.
-All other simulation settings were the same.
-Figure 5 shows the distribution of chosen $N$ values.
-The red star indicates the value used previously.
-Most repetitions halted surrogate training before the preset value, allotting more runs for estimation, but some proceeded longer.
-We view this variability as a strength – some random initializations need more data to effectively locate the contour while some get lucky from the start.
-Allowing the surrogate to dictate its own budget mitigates some of the risks of having an unlucky starting design.
-The resulting failure probability estimates are shown as the *hybMC ENT method (yellow) in Figure 3.
-The original hybMC ENT methods already performed well, and we were able to match this performance with the budgeted *hybMC ENT method.
-We acknowledge the benefit of the budgeting method is not in beating the un-budgeted performance on synthetic problems, but in informing budget allocations in real-world scenarios when there is no clear choice for $N$ and $B$.
+Across the board, both importance sampling based methods (“SIIS” and “SIIS UCB”) suffered from high variability in their failure probability estimates.
+Although their median performance was generally on target (indicative of their unbiasedness), these methods consistently offered the poorest performance.
+The more conservative “SIIS UCB” had no apparant advantage over the traditional “SIIS.” We believe importance sampling is simply ill-suited for our low data settings.
+In all scenarios aside from the simple two-dimensional Herbie function, our two-stage design (“Two-stage Entropy”) outperformed the exhaustive contour location approach (“All CL”).
+Particularly as dimension grows and the response surface becomes more complex, there is a clear advantage to reserving some simulator evaluations for the second stage.
+Most of the time there was agreement between the two criteria for allocating samples to $\mathcal{U}$ (“Entropy” and “Proximity”), but when there was disagreement the entropy criterion offered higher accuracy, with uniformly superior worst-case performance.
+As suspected, the entropy criterion which incorporates uncertainty is more effective at identifying samples which the surrogate would misclassify.
 5 RAE-2822 airfoil computer experiment
@@
 The simulation suite is solved via SU2, and the software is publicly available (Economon et al.,, 2016).
-The experiment accepts seven inputs: four shape parameters ($S_{i}$ for $i=1,\dots,4$) and 3 environmental parameters, namely angle of attack, Reynolds number, and Mach number.
+We consider seven inputs: four shape parameters ($S_{i}$ for $i=1,\dots,4$) and 3 environmental parameters, namely angle of attack, Reynolds number, and Mach number.
 We are interested in the lift over drag ($L/D$) ratio.
-$L/D$ values that are too low represent inefficient flight conditions.
+$L/D$ values that are too low represent inefficient aircraft performance.
 We define a failure threshold at $t=3$ such that $L/D<3$ indicates failure.
@@
 Our objective is to estimate the probability of failure given the following input distribution:
-$S_{i}$ $\sim\mathrm{Unif}(-1e-4,1e-4)\;\;\textrm{for}\;\;i=1,\dots,4$
+$S_{i}$ $\sim\mathrm{Unif}(-1\times 10^{-4},1\times 10^{-4})\;\;\textrm{for}\;\;i=1,\dots,4$
 Angle of Attack $\sim\mathcal{N}\left(\mu=5,\sigma=1\right)\;\;\textrm{truncated to}\;\;[0,10]$
-Reynolds number $\sim\mathcal{N}\left(\mu=1e7,\sigma=1e6\right)\;\;\textrm{truncated to}\;\;[5e6,1.5e7]$
+Reynolds number $\sim\mathcal{N}\left(\mu=1\times 10^{7},\sigma=1\times 10^{6}\right)\;\;\textrm{truncated to}\;\;[5\times 10^{6},1.5\times 10^{7}]$
 Mach number $\sim\mathcal{N}\left(\mu=0.8,\sigma=0.02\right)\;\;\textrm{truncated to}\;\;[0.7,0.9]$
-We fix our total budget at $T=500$ simulator evaluations.
+We fix our total budget at $N=500$ simulator evaluations and use $M=2.5\times 10^{6}$ for the size of $X_{M}$.
 Booth et al., (2024b) found that DGP surrogates outperformed GP surrogates for contour location of this experiment.
@@
 We proceed with contour location following Booth et al., (2024b), stopping after every 10th acquisition to estimate $\hat{\alpha}$ and check whether it is with in the ±$\pm\hat{\sigma}_{\alpha}$ bounds.
-Given the input dimension, we require atleast 40 observed failures before we consider halting surrogate training.
-Since the true failure probability is unknown, we repeat this entire exercise with a newly randomized starting design for some additional context.
+Once we halt contour location (Section 3.1), we use entropy to inform our hybrid MC estimator (Section 3.2).
+Given the larger input dimension, we require at least 40 observed failures before we consider halting surrogate training.
+Since the true failure probability is unknown, we opt to repeat this entire exercise twice, with a newly randomized starting design.
+We consider the agreement between the two randomized repetitions as an indication that our procedure is well-calibrated.
+Figure 7: Progress in $\hat{\alpha}$ across two contour locating designs for the airfoil computer experiment.
+Proposed two-stage failure probability estimate shown in dashed red.
+Grey shading indicates $\hat{\alpha}\pm 2\hat{\sigma}_{\alpha}$ for these final estimates.
+Left/right panels show two repetitions with newly randomized starting designs.
 Figure 7 reports the progress in $\hat{\alpha}$ over the course of each design.
-Figure 6: Two surrogate MC failure probabilty estimates for the airfoil computer experiment, updated every 10th acquisition with final estimates in dashed red.
-Grey shading indicates $\hat{\alpha}\pm\hat{\sigma}_{\alpha}$ for the final estimates.
-Our stopping criterion halted the sequential designs at $N=180$ and $N=160$.
-For each repetition, we collect a separate random sample of size $M=2.5\times 10^{6}$, then identify the remaining points with the highest entropy, and evaluate the true simulator at these locations.
-After updating the DGP surrogate, we obtain the final hybrid Monte Carlo estimates of $\hat{\alpha}_{\textrm{HYB}}=\{0.0003368,0.0003432\}$, respectively.
+Our stopping criterion halted the sequential designs at $N=180$ and $N=160$, leaving $B=320$ and $B=340$ for the second design stage, respectively.
+Our two hybrid Monte Carlo estimates were $\hat{\alpha}_{\textrm{HYB}}=\{0.0003368,0.0003432\}$.
 These values are indicated by the red dashed lines in Figure 7.
-The grey shading shows $\hat{\alpha}\pm\hat{\sigma}_{\alpha}$, calculated from these final hybrid estimates.
+The grey shading shows $\hat{\alpha}\pm 2\hat{\sigma}_{\alpha}$, calculated from these final estimates.
 Although the two estimates vary slightly, they are both within the bounds of reasonable variability for our chosen $M$.
-Additionally, notice how the final hybrid estimate of the second exercise (right panel) brought the failure probability lower than what the surrogate had estimated, closer to the estimate of our first attempt (left panel).
+Additionally, notice how the final hybrid MC estimate of the second exercise (right panel) brought the failure probability lower than what the surrogate had originally estimated, closer to the estimate of our first attempt (left panel).
 We take this as an indication that the hybrid estimator is effectively adjusting for surrogate inaccuracies.
 6 Discussion
-We presented a hybrid Monte Carlo method for estimating failure probabilities of expensive computer experiments.
-Our hybrid method leverages the uncertainty quantification provided by a surrogate to reserve expensive simulator evaluations for samples with high classification uncertainty.
-We limited our experiments to GP and DGP surrogates, but our proposed estimator is applicable to any surrogate which provides posterior predictive distributions (required for calculation of entropy).
+We presented a unique two-stage surrogate design for estimating failure probabilities of expensive computer experiments.
+Starting with a contour locating sequential design, we halt contour location once surrogate training has plateaued, as indicated by practical convergence of a surrogate Monte Carlo estimator.
+Then we use the remaining budget of simulator evaluations on the Monte Carlo samples with highest entropy, to ultimately return a hybrid MC estimate of $\alpha$.
+Our design strategically leverages the uncertainty quantification provided by the surrogate to reserve expensive simulator evaluations for samples with high classification uncertainty.
+While we limited our experiments to GP and DGP surrogates, our proposed scheme is applicable to any surrogate which provides posterior predictive distributions (which are required for entropy calculations).
 Extensions to other surrogates, such a treed Gaussian processes (Gramacy and Lee,, 2008) or Bayesian additive regression trees (Chipman et al.,, 2010), is straightforward.
-Our hybrid MC estimator is computationally efficient in the number of evaluations of the expensive simulator, but it requires many predictive evaluations from the surrogate.
-Posterior predictions must be gathered at all $M$ sample points.
-While surrogate predictions are very cheap, collecting millions of them is not necessarily trivial.
+Our two-stage design is computationally efficient in the number of evaluations of the expensive simulator, but it requires many predictive evaluations from the surrogate.
+Posterior predictions must be computed at all $M$ samples multiple times.
+While surrogate predictions are very cheap, collecting billions of them is not necessarily trivial.
 Our experiments were limited to failure probabilities above $1\times 10^{-5}$ for this very reason.
-We could tackle smaller failure probabilities, but it would require larger $M$, which would eventually become impractical.
-Our approach may simply not be suited to very rare failure probabilities.
-The efficacy of any surrogate-informed failure probability estimation relies on the surrogate’s ability to accurately classify failures.
-Although our hybrid estimator is generally applicable, we strongly recommend the use of contour locating sequential designs to train the surrogate.
-We proposed a budgeting scheme which informs allocation of simulator evaluations between surrogate training and failure probability estimation based on progress in the surrogate Monte Carlo estimate.
-If a budget is severly restricted, the budget could be exhausted training the surrogate before our proposed stopping criterion is met.
-While this is not ideal, we contend that fully-training a surrogate and using a surrogate MC estimate is the best option in this scenario.
-Particularly in light of the data hungry nature of IS-based methods, which perform poorly in small data settings.
+We could tackle smaller probabilities, but it would require larger $M$, which would eventually become impractical.
+Although we are focused on expensive simulators whose computational costs overshadow the cost of these surrogate predictions, our approach may simply not be suited to extremely rare failure probabilities.
+In all of our examples, our stopping criterion terminated contour location before the complete budget ($N$) was exhausted.
+If a simulation budget is severely restricted, the budget could be exhausted before our proposed stopping criterion is met.
+While we prefer to have additional evaluations reserved for the estimation stage, we contend that using the entire budget for contour location to provide a surrogate MC estimate is the best option in this scenario.
+As we’ve shown, importance sampling based methods perform poorly when data is limited.
+We have exclusively focused on deterministic computer experiments.
+In fact, the noise-free nature of the simulator is what enabled us to frame our contribution as a two-stage design.
+Once $f(\mathbf{x}_{i})$ has been observed and incorporated into the surrogate, surrogate predictions should interpolate these observations, i.e., $\mu_{n}(\mathbf{x}_{i})=f(\mathbf{x}_{i})$.
+Thus our hybrid MC estimator (7) is really a surrogate MC estimator (4), with a strategically trained surrogate.
+Different design choices could be warranted for stochastic simulators (Baker et al.,, 2022).
+In particular, the hybrid estimator of Eq.
+(7) should not make use of noisy $f(\mathbf{x})$.
+Replicates in the design may be crucial (Binois et al.,, 2019).
 When the input distribution is not uniform over $\mathcal{X}$, it could be beneficial to leverage information about $p(\mathbf{x})$ in the sequential training of the surrogate (Abdelmalek-Lomenech et al.,, 2024).
-There is arguably no need to train a surrogate in a region where inputs will not plausibly occur, such as in the outer corners of the Herbie function surface (Figure 1, center).
-Incorporating this information in the contour location stage could improve failure probability estimation downstream.
-But in the case of rare probabilities, it may be hard to draw the line between regions that are plausible and ones that are not.
-This avenue is currently underexplored.
+There is arguably no need to train a surrogate in a region where inputs will never occur.
+Incorporating this information in the contour location stage could improve failure probability estimation downstream, but we have seen positive impacts from having “anchor points” in regions of very low density (such as the black open circles around the edges of Figure 3).
+We suspect there is a delicate balance that would work well here, but this avenue is currently underexplored.
+Estimation of failure probabilities is often an intermediate step within a larger objective, such as a stochastically constrained optimization.
+For example, in the context of our airfoil experiment, we may wish to find the optimal wing configuration given that the probability of failure remains below a specified threshold.
+The ability to effectively estimate failure probabilites from limited data, which we have tackled here, is integral to the success of this larger enterprise.
 SUPPLEMENTARY MATERIAL
-Here we provide the details of the functions, thresholds, and input distributions used in the benchmark exercises of Section 3.2. In each case, failures are defined as $f(\mathbf{x})>t$.
+Here we provide the details of the functions, thresholds, and input distributions used in the benchmark exercises of Section 4. In each case, failures are defined as $f(\mathbf{x})>t$.
 Herbie.
@@
 $x_{2}$ $\sim\mathcal{N}(1.5,1.5)\;\;\textrm{truncated to}\;\;[-\pi,\pi]$
-$x_{3}$ $\sim\mathrm{Uniform}(-pi,pi).$
+$x_{3}$ $\sim\mathrm{Uniform}(-\pi,\pi).$
 Hartmann.
@@
 We use $d=4$ and set the failure threshold at $t=0$.
-We define the input distribution as $x_{i}\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathcal{N}(0.5,0.085)$ for $i=1,\dots,4$, truncated to $[0,1]$.
+We define the input distribution as $x_{i}\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathcal{N}(0.6,0.11)$ for $i=1,\dots,4$, truncated to $[0,1]$.

### v3

@@
 Abstract
-We tackle the problem of quantifying failure probabilities for expensive deterministic computer experiments with stochastic inputs.
-The computational cost of the computer simulation prohibits direct Monte Carlo (MC) and necessitates a surrogate model, turning the problem into a two-stage enterprise (surrogate training followed by probability estimation).
-Limited budgets create a design problem: how should expensive evaluations be allocated between and within the training and estimation stages?
-One may use the entire evaluation budget to sequentially train the surrogate through contour location (CL), with failure probabilities then estimated solely from the surrogate (we call it “surrogate MC”).
-But extended CL offers diminishing returns, and surrogate MC relies too stringently on surrogate accuracy.
-Alternatively, a partially trained surrogate may inform importance sampling, but this can provide erroneous results when budgets are limited.
-Instead we propose a two-stage design: starting with sequential CL, halting CL once learning has plateaued, then greedily allocating the remaining budget to MC samples with high classification entropy.
-Ultimately, we employ a “hybrid MC” estimator which leverages the trained surrogate in conjunction with the true responses observed in this second stage.
-Our unique two-stage design strikes an appropriate balance between exploring and exploiting, and outperforms alternatives, including both of the aforementioned approaches, on a variety of benchmark exercises.
+We tackle the problem of quantifying failure probabilities for expensive deterministic computer experiments with stochastic inputs under a fixed budget.
+The computational cost of the computer simulation prohibits direct Monte Carlo (MC) and necessitates a surrogate model, which may facilitate either a “surrogate MC” estimator or a surrogate-informed importance sampling estimator.
+We embrace the former, finding importance sampling too variable when budgets are limited, and propose a novel design strategy to effectively train a surrogate for the purpose of failure probability estimation.
+Existing works exhaust the entire evaluation budget on active learning through sequential contour location (CL), attempting to balance exploration with exploitation of the failure contour throughout the design, but we find exhaustive CL to be suboptimal.
+Instead we propose a novel two-stage surrogate design.
+In Stage 1, we conduct sequential CL to locate the failure contour.
+In Stage 2, once surrogate learning has saturated, we use a solely exploitative strategy – allocating the remaining evaluation budget to MC samples with the highest classification entropy to ensure they are classified correctly.
+We propose a stopping criterion to determine the transition between stages without any tuning.
+Our two-stage design outperforms alternatives, including exhaustive CL and surrogate-informed importance sampling, on a variety of benchmark exercises.
 With these tools, we are able to effectively estimate small failure probabilities with only hundreds of simulator evaluations, showcasing functionality with both shallow and deep Gaussian process surrogates, and deploying our method on a simulation of fluid flow around an airfoil.
-Keywords: computer experiment, contour location, deep Gaussian process, emulator, entropy, importance sampling, Monte Carlo
+Keywords: active learning, contour location, deep Gaussian process, entropy, importance sampling
 1 Introduction
-Quantifying the probability of failure in complex dynamic systems is of great interest but is challenging due to the computational expense of the so-called “computer experiment.” Consider a deterministic black-box computer simulation, $f:\mathcal{X}\rightarrow\mathbb{R}$ with inputs $\mathcal{X}\subset\mathbb{R}^{d}$ governed by known measure $\mathbb{P}_{\mathbf{x}}$ (with density $p(\mathbf{x})$).
+Quantifying the probability of failure in complex systems is of great interest but is challenging due to the computational expense of the so-called “computer experiment.” Consider a deterministic black-box computer simulation, $f:\mathcal{X}\rightarrow\mathbb{R}$ with inputs $\mathcal{X}\subset\mathbb{R}^{d}$ governed by known measure $\mathbb{P}_{\mathbf{x}}$ (with density $p(\mathbf{x})$).
 The response is thresholded such that $f(\mathbf{x})>t$ indicates failure.
@@
 […]
+from a small, fixed budget of simulator evaluations.
 We are particularly motivated by aeronautic applications where system failures may incur immense financial or environmental costs, or even loss of life.
@@
 A rudimentary approach to solve Eq.
-(1) involves brute force Monte Carlo (MC): simply sample an exhaustively large set of inputs, evaluate the computer simulation at each of these configurations, and report the proportion that resulted in failure.
-But this approach is prohibitively expensive if the computer simulation is costly (incurring long wall clock times), the probability of failure is rare (requiring enormous sample sizes), or both!
-In this work, we propose a new strategy for quantifying failure probabilities of complex computer simulations when the computation budget restricts evaluations of the simulator to only a few hundred, rendering traditional MC impractical.
+(1) involves brute force Monte Carlo (MC): sample a large number of inputs, evaluate the computer simulation at each of these configurations, and report the proportion that resulted in failure.
+But this approach is prohibitively expensive if the computer simulation is costly and $\alpha$ is small, as in our target application.
 In the face of expensive data, statistical models ($\hat{f}$, called “surrogates” or “emulators”) are essential.
 Surrogates are trained on a limited set of simulator evaluations in order to provide predictions with appropriate uncertainty quantification (UQ) at unobserved inputs.
-Gaussian processes (GPs) are the canonical choice (Santner et al.,, 2018; Gramacy,, 2020), but deep Gaussian processes (DGPs; Damianou and Lawrence,, 2013) are becoming increasingly popular, particularly for nonstationary computer experiments (Sauer,, 2023).
+Gaussian processes (GPs) are the canonical choice (Santner et al.,, 2018; Gramacy,, 2020), but deep Gaussian processes (DGPs; Damianou and Lawrence,, 2013) are becoming increasingly popular, particularly for nonstationary computer experiments (Booth et al.,, 2024a).
 Notably, our contribution is agnostic to surrogate choice as long as effective UQ is provided, and is thus applicable to traditional GPs and Bayesian DGPs (Sauer et al.,, 2023b) and, potentially, Bayesian neural networks (MacKay,, 1995).
-Choosing a surrogate model, like a GP or DGP, is only a first step.
-The utility of a surrogate boils down to two things: how it is trained and how it is used.
-For our objective (1), surrogate training should prioritize accurate classification of failures and surrogate deployment should focus on effective estimation of $\alpha$.
-The prevailing question is then: how can a limited budget of expensive simulator evaluations be optimally allocated between and within the training and estimation stages?
 We briefly digress to acknowledge some early methods of failure probability estimation, namely the first and second order reliability methods (FORM/SORM; Haldar and Mahadevan,, 1995), which leveraged simplistic linear approximations of the complex model and placed all failure probability estimation in the hands of the “most probable point” (MPP).
 Although advancements to FORM and SORM have been made with more flexible GP models (e.g., Zhang et al.,, 2015; Su et al.,, 2017), sole reliance on a MPP is too restricting for our target applications.
+The utility of a surrogate boils down to two things: how it is trained and how it is used.
 To solve Eq.
@@
 Space-filling designs (Joseph,, 2016) may miss failure regions altogether, particularly when training budgets are limited.
-Instead, it is common to train a surrogate on a small initial design then pair it with a relevant acquisition function to select subsequent inputs for evaluation through the simulator.
-This iterative process of acquisition and surrogate updating can strategically hone in on the contour, all while minimizing evaluations of the expensive simulator.
-So-called “contour location” (CL) naturally outperforms space-filling counterparts.
-Acquisition functions seek to balance exploitation and exploration by leveraging both predicted distance from the contour and posterior uncertainty.
+Instead, it is common to train a surrogate on a small initial design then pair it with a relevant acquisition function to select subsequent inputs for evaluation through the simulator – a process called “active learning” (AL).
+The goal is to maximize learning from limited evaluations of the expensive simulator.
+Contour location (CL) is an AL variant which targets learning of a failure contour $\{\mathbf{x}\in\mathcal{X}\mid f(\mathbf{x})=t\}$.
+CL acquisition functions seek to balance exploitation and exploration by leveraging both predicted distance from the contour and posterior uncertainty.
 Popular variations fall under expected improvement frameworks (Bichon et al.,, 2008; Ranjan et al.,, 2008), stepwise uncertainty reduction frameworks (Bect et al.,, 2012; Chevalier et al.,, 2014; Azzimonti et al.,, 2021), or both (Duhamel et al.,, 2023).
 Others utilize pass/fail classification entropy (Marques et al.,, 2018; Cole et al.,, 2023).
-Booth et al., (2024b) recently extended CL sequential designs to DGP surrogates.
+Booth et al., (2024c) recently proposed a CL sequential design for DGP surrogates.
 For small failure regions, contour location may be replaced with subset simulation (Au and Beck,, 2001) which initially targets a conservative threshold ($t^{\star}<t$) and iteratively refines it until the true threshold is found ($t^{\star}\rightarrow t$) (Dubourg et al.,, 2011; Huang et al.,, 2016; Hristov et al.,, 2019).
 Yet these approaches require Markov chain Monte Carlo (MCMC) sampling of cascading input distributions, which we intentionally avoid here.
-A well trained surrogate may be leveraged directly within an MC estimator, with surrogate predictions (which are of neglible cost) in place of true simulator evaluations.
-One could reasonably use the entire budget of simulator evaluations for contour location, then deploy the trained surrogate in such a “surrogate MC” estimator.
-Although this is common (e.g., Zhu and Du,, 2016; Cheng and Lu,, 2020; Li et al.,, 2021; Lu et al.,, 2023), we find that surrogate learning of failure contours offers quickly diminishing returns.
-Furthermore, surrogate MC estimators place too much stock in the accuracy of the surrogate and funnel surrogate inaccuracies directly into the failure probability estimate.
-Another common approach splits the budget of simulator evaluations into two disjoint stages: some evaluations are used to train the surrogate and some are used for importance sampling (IS; Srinivasan,, 2002; Tokdar and Kass,, 2010).
+A well-trained surrogate may be leveraged directly within an MC estimator with surrogate predictions (which are of negligible cost) in place of true simulator evaluations, which we term “surrogate MC.” One could reasonably use the entire budget of simulator evaluations for contour location, then deploy the trained surrogate in such a “surrogate MC” estimator.
+Although this is common (e.g., Zhu and Du,, 2016; Cheng and Lu,, 2020; Li et al.,, 2021; Lu et al.,, 2023), we find that surrogate learning of failure contours quickly plateaus, and that common surrogates fail to achieve sufficient classification accuracy under tightly restricted evaluation budgets.
+Another common approach splits the budget of simulator evaluations into two disjoint stages: some evaluations are used to train a surrogate and some are used for importance sampling (IS; Srinivasan,, 2002; Tokdar and Kass,, 2010).
 IS hinges on the construction of a bias distribution that intentionally focuses density in the failure region.
@@
 In SIIS, a trained surrogate is used to identify predicted failures, and a bias distribution is fit to these locations (more on this in Section 2.3).
-SIIS has been deployed with GP surrogates in a variety of subsequent works (e.g., Cole et al.,, 2023; Renganathan et al.,, 2023; Booth et al.,, 2024a; Renganathan,, 2024), and we view it as our primary competitor.
+SIIS has been deployed with GP surrogates in a variety of subsequent works (e.g., Cole et al.,, 2023; Renganathan et al.,, 2023; Booth et al.,, 2024b; Renganathan,, 2024), and we view it as our primary competitor.
 Yet SIIS has several crucial drawbacks.
 We will delay a thorough discussion to Section 2.3, but the highlights are: (i) SIIS relies too heavily on an arbitrary bias distribution, (ii) it largely ignores the carefully trained surrogate, and (iii) it offers no guidance for allocating the budget between surrogate training and IS estimation.
+Figure 1: Overview of proposed two-stage design for failure probability estimation.
+Stage 1 uses contour location to balance exploitation and exploration of $\{\mathbf{x}\in\mathcal{X}\mid f(\mathbf{x})=t\}$.
+Stage 2 solely exploits high uncertainty Monte Carlo samples from $p(\mathbf{x})$.
+For a fixed budget of $n+B$, our approach determines how to effectively allocate $n$ (for contour location) and $B$ (for Monte Carlo estimation).
 We prefer to avoid a bias distribution altogether, instead focusing on how we could optimally leverage a surrogate and an MC estimator in tandem.
-The answer lies in a two-stage design.
-First, we leverage sequential contour location to pin down the failure contour.
-At the start, it is essential to target inputs which provide the most insight into the entire failure contour (getting the most “bang for our buck” from early observations).
-As training progresses, successive acquisitions will offer less additional insight, indicated by apparent convergence of the surrogate MC estimator.
-Although we can not be sure of theoretical convergence with our limited budget, we propose a decision rule to recognize practical convergence (when surrogate learning from CL has saturated).
-We then spend the remaining simulation budget precisely on the Monte Carlo samples with the highest classification entropy, ultimately training the surrogate on observations from both stages.
-Our final failure probability estimate is a “hybrid MC” estimator which combines the observations from this second stage with surrogate predictions at unobserved samples.
+The answer lies in a two-stage surrogate design, represented in Figure 1 with details reserved for Section 3.
+First, in Stage 1 we conduct traditional contour location, balancing exploration and exploitation while locating the failure contour within domain $\mathcal{X}$.
+As training progresses, successive acquisitions will offer less additional insight, indicated by apparent convergence of the “surrogate MC” estimator.
+Although we can not be sure of theoretical convergence with our limited budget, we propose a stopping criterion to recognize practical convergence (when surrogate learning from CL has saturated).
+Once this stopping criterion is met, instead of continuing contour location or turning to importance sampling, we propose a novel “Stage 2” in which the remaining simulation budget is spent precisely on the Monte Carlo samples from $p(\mathbf{x})$ with the highest classification entropy.
+By observing the true $f(\mathbf{x})$ for samples where our surrogate is uncertain, we ensure these locations are classified correctly within our final MC estimate.
+Ultimately, surrogates trained on the observations acquired in these two unique stages provide more accurate and precise MC estimates of $\alpha$.
+While we prefer to view our contribution through the lens of surrogate design, our final failure probability estimate may be considered a “hybrid MC” estimator which combines the true simulator observations from Stage 2 with surrogate predictions at unobserved samples.
 This manuscript is laid out as follows.
 Section 2, reviews GP/DGP surrogates, contour location, and importance sampling.
-Section 3.2 presents our two-stage design and failure probability estimator.
+Section 3 presents our two-stage design and failure probability estimator.
 Section 4 presents a variety of synthetic benchmarks.
@@
 2 Preliminaries
-Denote single inputs as the row vector $\mathbf{x}$ of size $1\times d$ with response $y=f(\mathbf{x})$.
+Denote single inputs as the row vector $\mathbf{x}$ of size $1\times d$ with response $y=f(\mathbf{x})$, where $f$ represents the black-box computer simulation.
 Let $X_{n}$ represent the row-combined matrix of $n$-many inputs with corresponding response vector $\mathbf{y}_{n}$.
@@
 (1), namely $\mathrm{P}(f(\mathbf{x})>t)$ for $\mathbf{x}\sim p(\mathbf{x})$.
-Typical Monte Carlo with a budget of size $M$ provides the estimator
+Brute force Monte Carlo with a budget of size $M$ provides the estimator
 […]
-This estimator is ineffective unless a sufficient number of failures are observed, necessitating very large $M$ particularly for small $\alpha$.
-With an expensive computer model, we simply cannot evaluate $f(\mathbf{x})$ thousands/millions/billions of times.
-Throughout, let $N$ denote the total budget of expensive simulator evaluations alotted.
-We will describe our sample efficient method for estimating failure probabilities in Section 3.2, targeting $N$ in the hundreds, after a quick review of preliminaries.
+but this estimator is infeasible when $f$ is expensive.
+Let $N$ denote the total budget of expensive simulator evaluations allotted.
+We will describe our sample efficient method for estimating failure probabilities in Section 3, targeting $N$ in the hundreds, after a quick review of preliminaries.
 2.1 Gaussian process surrogates
-A surrogate model $\hat{f}$, trained on a limited budget of simulations, may provide predictions at unobserved inputs in place of the true $f$.
-To enable our proposed methodology, the surrogate must also provide posterior predictive distributions.
+A surrogate model $\hat{f}$, trained on a limited budget of simulations, may provide predictions for unobserved inputs in place of the true $f$.
+To enable our proposed methodology, the surrogate must also provide uncertainty quantification through posterior predictive distributions.
 Gaussian processes (GPs) are the predominant choice; they assume a multivariate normal distribution over the response.
-For any $n$ observed locations, our GP prior assumption is $\mathbf{y}_{n}\sim\mathcal{N}(\mu,\Sigma(X_{n}))$.
-The prior mean $\mu$ is often simplified to the zero vector after centering (which we will use moving forward), and the prior covariance $\Sigma(X_{n})$ is typically a function of inverse euclidean distances, e.g., $\Sigma(X_{n})^{ij}=\Sigma(\mathbf{x}_{i},\mathbf{x}_{j})=k(||\mathbf{x}_{i}-\mathbf{x}_{j}||^{2})$ with kernel $k(\cdot)$.
+For any $n$ observed locations, our GP prior assumption is $\mathbf{y}_{n}\sim\mathcal{N}_{n}(\boldsymbol{\mu},\Sigma(X_{n}))$.
+The prior mean $\boldsymbol{\mu}$ is often simplified to the zero vector after centering (which we will use moving forward), and the prior covariance $\Sigma(X_{n})$ is typically a function of Euclidean distances, e.g., $\Sigma(X_{n})^{ij}=\Sigma(\mathbf{x}_{i},\mathbf{x}_{j})=k(||\mathbf{x}_{i}-\mathbf{x}_{j}||^{2})$ with kernel $k(\cdot)$.
 See Santner et al., (2018); Rasmussen et al., (2006); Gramacy, (2020) for thorough reviews.
 Conditioned on training data $\left\{X_{n},\mathbf{y}_{n}\right\}$, posterior predictions at singleton input $\mathbf{x}$ follow
-$\hat{f}_{n}(\mathbf{x})\sim\mathcal{N}\left(\mu_{n}(\mathbf{x}),\sigma_{n}^{2}(\mathbf{x})\right)\quad\textrm{for}\quad\begin{array}[]{rl}\mu_{n}(\mathbf{x})&=\Sigma(\mathbf{x},X_{n})\Sigma(X_{n})^{-1}\mathbf{y}_{n}\\ \sigma_{n}^{2}(\mathbf{x})&=\Sigma(\mathbf{x},\mathbf{x})-\Sigma(\mathbf{x},X_{n})\Sigma(X_{n})^{-1}\Sigma(\mathbf{x},X_{n})^{\top},\end{array}$ (3)
+$\hat{f}_{n}(\mathbf{x})\sim\mathcal{N}_{1}\left(\mu_{n}(\mathbf{x}),\sigma_{n}^{2}(\mathbf{x})\right)\quad\textrm{for}\quad\begin{array}[]{rl}\mu_{n}(\mathbf{x})&=\Sigma(\mathbf{x},X_{n})\Sigma(X_{n})^{-1}\mathbf{y}_{n}\\ \sigma_{n}^{2}(\mathbf{x})&=\Sigma(\mathbf{x},\mathbf{x})-\Sigma(\mathbf{x},X_{n})\Sigma(X_{n})^{-1}\Sigma(\mathbf{x},X_{n})^{\top},\end{array}$ (3)
 where $\Sigma(\mathbf{x},X_{n})$ denotes the vector formed from applying the kernel between $\mathbf{x}$ and every row of $X_{n}$.
 Joint predictions are straightforward but not necessary for our contribution.
-A natural alternative to true MC (2) is “surrogate MC” which leverages surrogate predictions in place of the true simulator,
+A natural alternative to brute force MC (2) is “surrogate MC” which leverages surrogate predictions in place of the true simulator,
 […]
-Here we use the posterior mean of our GP in place of $f$. If the surrogate is not accurate near the contour, it will misclassify inputs which will severely hinder the effectiveness of this estimator.
-Figure 1: Left/Center: Herbie function with failure contour (red) and input distribution (white). Right: Entropy surface (yellow/high, purple/low) for a GP trained on a 100-point LHS (white circles).
-Figure 2: Slice of Herbie function along $x_{2}=1.1$. GP was trained on 100-point LHS in 2 dimensions.
-For example, consider the two-dimensional “Herbie” function (Lee et al.,, 2011) shown in the left and center panels of Figure 1 (we will revisit the right panel momentarily).
+Here we use the posterior mean of our surrogate $\hat{f}$ in place of the true $f$.
+This estimator, however, inherits the bias of the surrogate – if the surrogate is not accurate near the contour, it will misclassify inputs which will severely hinder the effectiveness of this estimator.
+Figure 2: Left: Herbie function (yellow/high, purple/low) with true failure contour (red) and input distribution (white).
+Center/Right: Posterior predicted mean and entropy of a GP trained on an initial 20-point LHS (white circles) followed by 30 contour locating acquisitions (white triangles).
+Predicted contour in dashed red.
+Dashed line at $x_{2}=1.05$ indicates the slice shown in Figure 3.
+Figure 3: Slice of Herbie function along $x_{2}=1.05$ (solid black) with failure threshold (dotted red).
+GP surrogate from Figure 2 shown in blue, with entropy in green along the $x$-axis.
+Yellow shaded regions indicate inputs which the surrogate would misclassify.
+For example, consider the two-dimensional “Herbie” function (Lee et al.,, 2011) shown in the left panel of Figure 2 (we will revisit the other panels momentarily).
 We define a failure threshold of $t=1.065$ which demarcates four disjoint failure regions (red contours).
@@
 [^1] Full details of all simulated functions, thresholds, and distributions are provided in the Supplementary Material.
-We trained a GP surrogate using a random Latin hypercube sample (LHS; McKay et al.,, 2000) of size $n=100$ in two dimensions.
-Figure 2 shows the GP posterior prediction along the slice $x_{2}=1.1$.
-The posterior mean (dashed blue) and 95% confidence interval (blue shading) capture the true nonlinear surface (solid black) fairly well.
-Even so, there are inputs (highlighted by the yellow shading) which the surrogate would misclassify as “passes” instead of “failures.” Thus any estimate reliant solely on this surrogate would underestimate the failure probability.
+We trained a GP surrogate using an initial random Latin hypercube sample (LHS; McKay et al.,, 2000) of size $n_{0}=20$ (white circles), followed by 30 contour locating acquisitions (white triangles, discussed in Section 2.2).
+Figure 3 shows the GP posterior along the slice $x_{2}=1.05$.
+The posterior mean (dashed blue) and 99% credible interval (blue shading) capture the true nonlinear surface (solid black) fairly well.
+Even so, there are inputs (highlighted by the yellow shading) which the surrogate would misclassify as “passes” instead of “failures” or vice versa.
+While the predictions in these regions are reasonably accurate, their proximity to the contour makes precise classification difficult.
 Deep Gaussian processes.
@@
 The computer simulation is evaluated at the selected input(s), the surrogate is updated, and the process is repeated until the budget is exhausted or a stopping criterion is met.
+Stage 1 of Figure 1 visualizes this iterative loop.
+While acquisitions may be made in batches, we only consider single acquisitions here.
 Identifying acquisition criteria which effectively target failure contours is a hot topic, particularly in conjunction with GP surrogates, (e.g., Bichon et al.,, 2008; Ranjan et al.,, 2008; Bect et al.,, 2012; Chevalier et al.,, 2014; Marques et al.,, 2018; Cole et al.,, 2023).
@@
 $H(\mathbf{x})=-p_{\mathbf{x}}\log(p_{\mathbf{x}})-(1-p_{\mathbf{x}})\log(1-p_{\mathbf{x}}),$ (5)
-where $p_{\mathbf{x}}$ represents the probability of failure at input $\mathbf{x}$. With a GP surrogate, this probability boils down to a Gaussian CDF computation:
-$p_{\mathbf{x}}=\mathrm{P}\left(\hat{f}_{n}(\mathbf{x})>t\right)=1-\Phi\left(\frac{t-\mu_{n}(\mathbf{x})}{\sigma_{n}(\mathbf{x})}\right).$
-Returning to the Herbie example, the entropy surface for the trained GP is shown in the right panel of Figure 1 and along the slice in Figure 2 (green dotted/dashed).
-Notice the high entropy regions mostly agree with the locations of the true failure contour (we could improve upon this with a CL design instead of an LHS).
-A reasonable acquisition could fall in any of these local optima.
-Naturally, implementations of entropy as an aqcuisition function vary on the finer details.
+where $p_{\mathbf{x}}$ represents the predicted probability of failure at input $\mathbf{x}$, i.e., $\mathrm{P}\left(\hat{f}_{n}(\mathbf{x})>t\right)$.
+With a GP surrogate, this predicted probability boils down to a Gaussian CDF computation,
+$p_{\mathbf{x}}=\mathrm{P}\left(\hat{f}_{n}(\mathbf{x})>t\right)=1-\Phi\left(\frac{t-\mu_{n}(\mathbf{x})}{\sigma_{n}(\mathbf{x})}\right),$
+with $\mu_{n}(\mathbf{x})$ and $\sigma_{n}(\mathbf{x})$ following Eq.
+(3).
+Revisiting the Herbie example of Section 2.1, we offer two visuals of the entropy surface for the trained GP.
+First, the right panel of Figure 2 shows entropy over the entire 2d space.
+Notice high entropy regions are predominantly near the failure contour and/or away from training data locations.
+Second, Figure 3 shows the entropy surface (green) along the slice $x_{2}=1.05$.
+Entropy is high where the credible interval (blue shading) captures the failure threshold (dotted red).
+It is influenced by how much of the credible interval is split above/below $t$, regardless of the interval’s width.
+For the highest $x_{1}$ values, surrogate uncertainty is high but entropy is low because the predicted values are far from the failure threshold.
+As an acquisition criterion, entropy effectively identifies regions where the surrogate is uncertain of its pass/fail prediction.
+A reasonable acquisition could fall in any of the local optima of the entropy surface.
+Naturally, implementations of entropy as an acquisition function vary on the finer details.
 For GP surrogates we will leverage the optimization scheme of Cole et al., (2023), which first deploys a small LHS, identifies the LHS sample with the highest entropy, then runs a numerical optimization of entropy initialized at that point.
 By avoiding multi-starts, Cole et al., target local optima in the entropy surface.
-For DGP surrogates, we will follow Booth et al., (2024b).
+For DGP surrogates, we will follow Booth et al., (2024c).
 To avoid cumbersome numerical optimizations with the heftier MCMC-based DGP, Booth et al., use triangulation candidates (Gramacy et al.,, 2022), then select the candidate on the Pareto front of entropy and uncertainty.
 Henceforth, all GPs and DGPs we consider will be trained with these two contour location methods, respectively.
-Figure 3: Various designs for estimating the failure probability of the Herbie function with a GP surrogate and a total budget of $N=150$.
-All methods start with the same CL design of size $n=100$ (black circles), leaving a remaining budget of $B=50$ (blue triangles) for either continued surrogate training or importance sampling.
-Left: Entire budget is spent on continued contour location, with acquisitions exploring the entire space.
-Center: Remaining budget is sampled from a bias distribution for importance sampling (black circles omitted since they do not feature in the SIIS estimate).
-Right: Remaining budget is used for MC samples with highest entropy, solely exploiting near the predicted contour.
-Comparison of failure probability estimates obtained from these is provided later in Section 4.
-Although there is no clear distinction in performace between the left/right designs for this simple illustrative example, the continued exploration of the “All CL” design will hinder performance on larger, harder problems (Figure 6).
-Continuing the Herbie function example, the left panel of Figure 3 shows the design points chosen from entropy-based contour location with a GP surrogate (Cole et al.,, 2023), indicated by both the black circles and blue triangles (the distinction between the two will become clear when we discuss the other panels).
-Notice, the chosen points attempt to balance exploration (points spread out over the entire domain) and exploitation (points hovering around the red failure contour).
-The resulting surrogate may inform an $\hat{\alpha}_{\textrm{SURR}}$ estimate (4); we will include this as a competitor in our later exercises.
+The right and center panels of Figure 2 showed a contour locating design for the Herbie function.
+Starting with a GP surrogate trained on an LHS design of size $n_{0}=20$ (white circles), we used the entropy-based scheme of Cole et al., (2023) to acquire 30 more inputs, one-at-a-time (white triangles).
+The acquired points attempt to balance exploring the space and exploiting learning of the failure contour.
 2.3 Importance sampling
@@
 […]
-Monte Carlo approximation of this integral provides the estimator
+Monte Carlo approximation of this integral provides the estimator,
 […]
+This estimator requires $B$-many evaluations of the expensive simulator $f$.
 The “bias weights” $w_{i}$ account for the difference between the input density and the bias density.
-This estimator is theoretically unbiased and requires a smaller budget of simulator evaluations ($B\ll M$) if the bias distribution appropriately covers and targets the failure region (Srinivasan,, 2002).
+This estimator is theoretically unbiased for any $B>0$ and requires a smaller budget of simulator evaluations ($B\ll M$) if the bias distribution appropriately covers and targets the failure region (Srinivasan,, 2002).
 But estimation of an effective bias distribution is tricky.
@@
 In “surrogate-informed importance sampling,” Peherstorfer et al., (2016) leverage a surrogate to identify predicted failures from a large Monte Carlo sample of the input distribution, then train a GMM bias distribution to these predicted failures.
+Recent works have leveraged contour location to train surrogates for this very purpose (e.g., Cole et al.,, 2023; Booth et al.,, 2024b).
 If the surrogate effectively identifies the failure region, then these samples should mimic the ideal $q(\mathbf{x})^{*}$.
@@
 (6) to provide an estimate of the failure probability.
-To illustrate, consider our running Herbie function example.
-Using the GP surrogate trained on the 100-point contour locating design from Figure 3 (black circles) and a large MC sample from the input distribution, we trained a GMM bias distribution to the MC samples which the surrogate predicted to be failures.
-The blue triangles in the center panel show 50 points sampled from this bias distribution.
-These samples may be evaluated through the true model to estimate $\hat{\alpha}_{\textrm{IS}}$ (6), thereby exhausting our entire budget of 150 evaluations.
-While the blue triangles target the failure region, they do not provide enough coverage to ensure a precise estimate.
-In our later experiments (Section 4), SIIS performance is highly variable.
-While SIIS has been widely embraced for its unbiasedness, it has several key weaknesses.
+SIIS is a two-stage enterprise (first training a surrogate then using the surrogate to estimate a bias distribution for importance sampling), but it is not a “two-stage design” since the final estimator makes no use of the surrogate nor the observations that had been used to train the surrogate.
+Although SIIS has been widely embraced for its unbiasedness, it has several key weaknesses.
 First, the bias distribution, which features heavily in the final estimate of Eq.
-(6), is a completely auxilary quantity.
+(6), is a completely auxiliary quantity.
 Hypothetically, two different bias distributions could produce the same samples $\left\{\mathbf{x}_{i}\right\}_{i=1}^{B}$, thus producing the same $\left\{f(\mathbf{x}_{i})\right\}_{i=1}^{B}$.
@@
 Furthermore, there is no clear avenue for determining the budget allocation between surrogate training and importance sampling estimation.
-[Throughout, we will borrow from our proposed method to inform budget allocation for the SIIS competitor, which is how we chose to separate the budget of Figure 3 into 100 for training and 50 for estimation.] Finally, the SIIS method does not leverage the full potential of the surrogate model.
-Many expensive simulator evaluations go into training the surrogate and, in the end, the surrogate is only used to identify predicted failures, with no mention of its UQ.
-[Notice, as portrayed in the center panel of Figure 3, the black circles are abandoned after the bias distribution is trained.]
+[In our benchmark exercises, we will borrow from our proposed method to inform budget allocation for the SIIS competitor.] Finally, the SIIS method does not leverage the full potential of the surrogate model.
+Many expensive simulator evaluations go into training the surrogate, but the surrogate is abandoned after it has identified potential bias locations for estimation of $q(\mathbf{x})$.
 3 Two-stage design for failure probability estimation
-All in all, we are unsatisfied with the two state-of-the-art surrogate-based methods for estimating $\alpha$ which we just described: surrogate MC and SIIS.
-While the drawbacks to SIIS are apparent, the limitations of surrogate MC deserve more attention.
-To investigate further, we conducted the same CL sequential design of the Herbie function (Figure 3), this time storing the surrogate MC estimate $\hat{\alpha}_{\textrm{SURR}}$ (4), which we now denote as simply $\hat{\alpha}$, after every 10th acquisition.
-Progress is shown in the left panel of Figure 4.
+Ultimately, we are unsatisfied with the two aforementioned surrogate-based methods for estimating $\alpha$: exhaustive CL with surrogate MC and surrogate-informed importance sampling.
+While the drawbacks to SIIS are apparent, the limitations of exhaustive CL deserve more attention.
+To investigate further, we revisit the CL design of the Herbie function (Figure 2), this time continuing contour location until an entire budget of $N=150$ has been exhausted.
+As the design progresses, we evaluate and store the surrogate MC estimate $\hat{\alpha}_{\textrm{SURR}}$ (4), which we now denote as simply $\hat{\alpha}$, after every 10th CL acquisition.
+Progress in this estimate is shown in the left panel of Figure 4.
 While there is significant learning occurring with early acquisitions, it quickly plateaus.
 By the time the surrogate has been trained on 100 points, it knows just as much about $\alpha$ as it does when it has been trained on the full budget of $N=150$ points.
-In this case, there is bound to be a more effective way to use the remaining simulation budget.
-We propose a solution: first developing a stopping criterion to decide when to halt contour location, then proposing a hybrid Monte Carlo estimator that greedily leverages the remaining simulation budget precisely for failure probability estimation.
-Figure 4: Left: Surrogate MC failure probability estimate for the Herbie function, updated after every 10th contour locating acquisition.
+In this case, continued contour location is effectively wasting the remaining simulation budget.
+We propose a solution: first developing a stopping criterion to decide when to halt contour location (Figure 1, Stage 1), then proposing a second design stage (Figure 1, Stage 2) which greedily leverages the remaining simulation budget for Monte Carlo samples with the highest classification uncertainty.
+The second stage, in which observations are acquired directly from $p(\mathbf{x})$ enables a“hybrid Monte Carlo” estimator, using both true simulator observations and surrogate predictions.
+Figure 4: Left: Failure probability estimate (4) from a GP surrogate of the Herbie function (as shown in Figure 2), updated after every 10th acquisition of the CL design.
 True $\alpha$ in dashed red.
-Right: Difference in successive estimates with ±$\pm\hat{\sigma}_{\alpha}$ shown in dashed green.
-Blue square marks the second update within ±$\pm\hat{\sigma}_{\alpha}$.
+Right: Difference in successive $\hat{\alpha}$ estimates with ±$\pm\hat{\sigma}_{\alpha}$ shown in dashed green.
+Narrow teal line indicates true $\sigma_{\alpha}$.
+Blue square marks the second update within ±$\pm\hat{\sigma}_{\alpha}$, which satisfies our proposed stopping criterion.
 3.1 A stopping criterion for contour location
-Let $n$ denote the budget used on contour location and $B$ denote the remaining budget such that the total budget is $N=n+B$.
-We seek an objective, automatable decision rule for selecting $n$.
-The right panel of Figure 4 shows the relative difference between each successive $\hat{\alpha}$ estimate from the left panel.
-As expected, the relative differences converge to zero as the surrogate hones in on the true contour, but they never settle at exactly zero.
-How small of an update then is small enough to indicate learning has plateaued?
-The significance of an update to $\hat{\alpha}$ is only interpretable within the context of $\alpha$, the true failure probability, and $M$, the size of the MC sample used.
-Small differences might be inconsequential for large $\alpha$ with small $M$, but could be significant for small $\alpha$ with large $M$.
-To this end, we propose the use of the standard error $\sigma_{\alpha}$ to quantify the relative magnitude of each update.
-Notice in the estimator of Eq.
-(2), the indicator $\mathbbm{1}_{\left\{f(\mathbf{x})>t\right\}}$ is a Bernoulli random variable with expected value $\alpha$ and variance $\alpha(1-\alpha)$.
+Let $n$ denote the budget of expensive simulator evaluations used on contour location (Stage 1).
+Let $B$ denote the remaining evaluation budget, such that the total budget is $N=n+B$.
+We seek an objective, automatable decision rule for selecting $n$, halting CL once surrogate learning has plateaued.
+Consider the progress in $\hat{\alpha}$ shown in Figure 4.
+The right panel shows the relative difference between each successive $\hat{\alpha}$ estimate from the left panel.
+As expected, the relative differences converge to zero as the surrogate hones in on the true contour, but with this finite budget they never settle at exactly zero.
+We need an objective method to decide when these updates to $\hat{\alpha}$ are no longer significant.
+The challenge is that $\hat{\alpha}$ exhibits variability from two distinct sources: surrogate error and the natural variability that accompanies any Monte Carlo estimator with finite $M$.
+Hypothetically, presume we could evaluate the true $f$ an unlimited number of times to obtain the brute force MC estimator of Eq.
+(2).
+In this estimator, the indicator $\mathbbm{1}_{\left\{f(\mathbf{x})>t\right\}}$ is a Bernoulli random variable with expected value $\alpha$ and variance $\alpha(1-\alpha)$.
 Thus, the MC estimator has variance inversely proportional to $M$:
 $\mathbb{V}[\hat{\alpha}_{\textrm{MC}}]=\frac{1}{M^{2}}\sum_{i=1}^{M}\mathbb{V}\left[\mathbbm{1}_{\left\{f(\mathbf{x})>t\right\}}\right]=\frac{\alpha(1-\alpha)}{M}.$
-Let $\sigma_{\alpha}=\sqrt{\frac{\alpha(1-\alpha)}{M}}$ denote the standard error of this estimator.
-This standard error intrinsically accounts for the magnitude of $\alpha$ and $M$.
+Let $\sigma_{\alpha}=\sqrt{\frac{\alpha(1-\alpha)}{M}}$ denote the standard error of this estimator; it quantifies the degree of variability that we would expect even if we used the true $f$ (i.e., zero surrogate error).
+We consider successive updates to $\hat{\alpha}$ within this realm of intrinsic MC variability to indicate stabilization of the surrogate.
 Since the true $\alpha$ is unknown, we estimate $\hat{\sigma}_{\alpha}=\sqrt{\frac{\hat{\alpha}(1-\hat{\alpha})}{M}}$ using the current surrogate MC estimate.
-The green dashed lines in the right panel of Figure 4 show ±$\pm\hat{\sigma}_{\alpha}$, updated along with $\hat{\alpha}$.
-At the start, updates to $\hat{\alpha}$ are far beyond one standard error.
+While this does intertwine surrogate error and MC error, it still provides a useful and feasible estimate.
+To demonstrate, the green dashed lines in the right panel of Figure 4 show ±$\pm\hat{\sigma}_{\alpha}$, updated with $\hat{\alpha}$ after every 10th acquisition.
+The estimated $\hat{\sigma}_{\alpha}$ (dashed green) effectively reflects the true $\sigma_{\alpha}$ (narrow solid teal) after very few acquisitions.
+At the start of the design, updates to $\hat{\alpha}$ are far beyond this estimated standard error.
 But after a while, the updates sit comfortably within these bounds.
 Let $X_{M}$ denote the row-combined matrix of $M$-many samples from $p(\mathbf{x})$.
-We posit that, for fixed $X_{M}$, any update to $\hat{\alpha}$ within ±$\pm\hat{\sigma}$ is “small enough.” Accordingly, we propose halting CL once we have observed two successive updates within ±$\pm\hat{\sigma}$.
-The blue square in Figure 4 marks this point, stopping CL for this example at $n=100$, effectively identifying a point where surrogate learning has saturated and leaving a remaining budget of $B=50$ (this informed the budget allocation of Figure 3).
+We posit that, for fixed $X_{M}$, any update to $\hat{\alpha}$ within ±$\pm\hat{\sigma}$ is insignificant, as it is within the realm of expected MC error.
+[To avoid excessive computation, we recommend updating estimates after every 10 acquisitions, which we will do in all later exercises.] Accordingly, we propose halting CL once we have observed two successive updates within ±$\pm\hat{\sigma}$.
+The blue square in Figure 4 marks this point, stopping CL for this example at $n=100$, effectively identifying a point where surrogate learning has saturated and leaving a remaining budget of $B=50$.
 Note, this stopping criterion requires the same $X_{M}$ be used in every $\hat{\alpha}$ calculation.
-If a new $X_{M}$ is generated for each estimate, there will be additional variability among the $\hat{\alpha}$’s, and the proposed ±$\pm\hat{\sigma}_{\alpha}$ threshold would need to be increased.
-Removing this extraneous source of variability allows successive estimates to pinpoint differences stemming solely from the surrogate.
+If a new $X_{M}$ is generated for each estimate, there would be additional sampling variability reflected among the $\hat{\alpha}$’s (on top of the surrogate error and MC variability).
 Our budgeting scheme is summarized as follows.
@@
 At the start, it is likely that there will be no failures observed and that $\hat{\alpha}$ will be equal to zero (as seen in Figure 4).
-To safeguard against halting surrogate training prematurely, we build in a “common sense” check by setting a minimum number of failures that must be observed before the contour location could be halted.
-If the contour location design is succeeding, it will be placing some points in the failure region.
+To safeguard against halting surrogate training prematurely, we build in a “common sense” check by setting a minimum number of failures that must be observed before contour location could be halted.
+If the contour locating design is succeeding, it will be placing some points in the failure region.
 We recommend a default of 10 failures, but suggest increasing this for more complex, higher dimensional problems.
@@
 Calculate $\hat{\sigma}_{\alpha}$ and check if $|\hat{\alpha}^{\textrm{(current)}}-\hat{\alpha}^{\textrm{(previous)}}|<\hat{\sigma}_{\alpha}$.
-When this condition has been satisfied twice in a row, halt the sequential design and proceed to the next stage (the topic of Section 3.2).
+When this condition has been satisfied twice in a row, halt the CL design of Stage 1 and proceed to Stage 2 (the topic of Section 3.2).
 Choosing $M$.
@@
 3) for every sample in $X_{M}$.
-Although this computation is significantly cheaper than evaluation of the expensive simulator, it may still be significant for extremely large $M$, particularly with a heftier surrogate like a Bayesian DGP.
-If we knew $\alpha$, we could determine $M$ by thresholding the standard error $\sigma_{\alpha}$.
-For example, to achieve a standard error less than 10% of $\alpha$ requires:
+Although this computation is significantly cheaper than evaluation of the expensive simulator, it could be significant for extremely large $M$, particularly with a heftier surrogate like a Bayesian DGP.
+If we knew $\alpha$, we could determine $M$ by thresholding the standard error $\sigma_{\alpha}$. For example, to achieve a standard error less than 10% of $\alpha$ requires:
 $\sigma_{\alpha}<\frac{\alpha}{10}\quad\longrightarrow\quad\frac{100(1-\alpha)}{\alpha}<M.$
@@
 Gather surrogate predictions along this incremental process, and use $\hat{\alpha}$ to inform $\hat{\sigma}_{\alpha}$.
-3.2 Hybrid Monte Carlo for failure probability estimation
-Once $n$ is determined, how should we spend the remaining budget of simulator evaluations ($B=N-n$), in conjunction with our trained surrogate, to most effectively estimate $\alpha$?
-We propose an upgrade to the surrogate MC estimator of Eq.
+3.2 Stage 2: Hybrid Monte Carlo for failure probability estimation
+Once $n$ is determined and the contour location of Stage 1 is complete, how should we spend the remaining budget of simulator evaluations ($B=N-n$) in Stage 2?
+Recall, our objective is to leverage the trained surrogate to estimate $\alpha$ (1).
+We propose a new design for Stage 2, which enables an upgrade to the surrogate MC estimator of Eq.
 (4).
-Let $\mathcal{U}\subset\mathcal{X}$ represent some subset of the input domain (boundary inclusive).
-Then, Eq.
-(1) may be partitioned as
+To set the stage, let $\mathcal{U}\subset\mathcal{X}$ represent some subset of the input domain (boundary inclusive). Then, Eq. (1) may be partitioned as
 […]
@@
 […]
-Now for $\mathbf{x}_{i}\in\mathcal{U}$, upgrade surrogate prediction $\mu_{n}(\mathbf{x}_{i})$ to expensive evaluation of the true $f(\mathbf{x}_{i})$, resulting in the following “hybrid MC” estimator:
+Now for $\mathbf{x}_{i}\in\mathcal{U}$, we upgrade surrogate prediction $\mu_{n}(\mathbf{x}_{i})$ to expensive evaluation of the true $f(\mathbf{x}_{i})$, resulting in the following “hybrid MC” estimator:
 […]
@@
 The accuracy of this estimator is contingent upon the classification accuracy of the surrogate and the choice of $\mathcal{U}$.
-As long as we restrict the size of $\mathcal{U}$ so only $B$-many samples from $X_{M}$ fall within it, i.e., $\sum_{i=1}^{M}\mathbbm{1}_{\left\{\mathbf{x}_{i}\in\mathcal{U}\right\}}=B$, then we will meet our budgetary constraints.
+If we restrict the size of $\mathcal{U}$ so only $B$-many samples from $X_{M}$ fall within it, i.e., $\sum_{i=1}^{M}\mathbbm{1}_{\left\{\mathbf{x}_{i}\in\mathcal{U}\right\}}=B$, then we can use this hybrid estimator while staying within our budgetary constraints.
+The evaluation of the expensive simulator at the Monte Carlo samples within $\mathcal{U}$ constitutes the second stage of our design.
 Rather than thinking of $\mathcal{U}$ as a continuum, we find it easier to conceptualize in the context of the discrete set $X_{M}$.
@@
 Denote these as $X_{M}^{B}$.
-Revisiting Figure 2, notice the high entropy regions (green dotted/dashed) align well with the misclassification regions (yellow shading).
-The entropy criterion is effectively identifying regions of high classification uncertainty where we could most utilize expensive true simulator evaluations to correct inaccurate surrogate predictions.
-Ultimately, after evaluating $f(\mathbf{x})$ for $\mathbf{x}\in X_{M}^{B}$, we update the surrogate with these additional observations, thereby updating $\mu_{n}(\mathbf{x})$, before returning $\hat{\alpha}_{\textrm{HYB}}$ (7).
-Since the $X_{M}^{B}$ points are likely near a failure contour, they could significantly improve the accuracy of the surrogate predictions for other samples near the contour.
-For reference, consider the right panel of Figure 3.
-The black circles showing the contour locating design for the GP surrogate are unchanged, but here the additional budget of 50 evaluations is spent on the MC samples from $p(\mathbf{x})$ with the highest classification entropy (blue triangles).
-Compared to the left panel which continues to explore the entire domain, these highest entropy samples exclusively exploit the predicted failure contour.
-Compared to the bias samples portrayed in the middle panel, where the original surrogate design is abandoned, our two-stage design is able to leverage all simulator evaluations in its final estimate.
+Revisiting Figure 3, notice the high entropy regions (green) align well with the misclassification regions (yellow shading).
+The entropy criterion is effectively identifying regions of high classification uncertainty where we could most utilize expensive true simulator evaluations to correct inaccurate surrogate classifications.
+Ultimately, after completing the second design stage by evaluating $f(\mathbf{x})$ for $\mathbf{x}\in X_{M}^{B}$, we update the surrogate with these additional observations, thereby updating $\mu_{n}(\mathbf{x})$, before returning $\hat{\alpha}_{\textrm{HYB}}$ (7).
+The process of observing the true simulator at the high entropy locations and updating the surrogate accordingly could be done in small batches within a feedback loop as demonstrated in Figure 1.
+Yet in our exercises, we have not seen smaller batch sizes to improve performance (we provide empirical evidence of this in Supplementary Material).
+Accordingly, we prefer to collect a single batch of all $B$-many high entropy samples at once.
 3.3 Summary and related work
@@
 Start with a very large Monte Carlo sample from $p(\mathbf{x})$, denoted $X_{M}$.
-Train a surrogate model through a contour locating sequential design, monitoring the progress of the $\hat{\alpha}_{\textrm{SURR}}$ estimate for given $X_{M}$, and halting contour location once two consecutive updates are within ±$\pm\hat{\sigma}_{\alpha}$.
+Train a surrogate model through a contour locating sequential design, monitoring the progress of the $\hat{\alpha}_{\textrm{SURR}}$ estimate (4) for given $X_{M}$, and halting contour location once two consecutive updates are within ±$\pm\hat{\sigma}_{\alpha}$.
 Denote the design points chosen in this stage as $X_{n}$.
@@
 (7).
-Given the deterministic nature of $f(\mathbf{x})$, effective surrogates should interpolate observed points.
+Given the deterministic nature of $f(\mathbf{x})$, effective surrogates should interpolate observed points, i.e., $\mu_{n}(\mathbf{x}_{i})=f(\mathbf{x}_{i})$ when the surrogate has been trained on $\{\mathbf{x}_{i},y_{i}\}$.
 In this light, the hybrid MC estimator (7) is akin to a surrogate MC estimator (4) when the surrogate has been trained on all $\mathbf{x}_{i}\in\mathcal{U}$.
@@
 The final design of size $N$ is then $X_{N}=[X_{n}^{\top},{X_{M}^{B}}^{\top}]^{\top}$.
-There are some noteworthy distinctions between these two design stages.
-To illustrate, refer back to the black circles ($X_{n}$) and blue triangles ($X_{M}^{B}$) in the right panel of Figure 3.
+There are some noteworthy distinctions between these two design stages, as depicted in Figure 1.
 In the first stage, the $n$ points selected for contour location may be acquired anywhere in the domain $\mathcal{X}$, seeking to balance both exploration of the response surface and exploitation of the suspected contour.
-In fact, the contour locating sequential design strategies we consider in Section 4 (Cole et al.,, 2023; Booth et al.,, 2024b) intentionally circumvent acquisitions based on maximal entropy in order to avoid clustered acquisitions and to promote more exploration of the surface.
+In fact, the contour locating sequential design strategies we consider in Section 4 (Cole et al.,, 2023; Booth et al.,, 2024c) intentionally circumvent acquisitions based on maximal entropy in order to avoid clustered acquisitions and to promote more exploration of the surface.
 Alternatively, the $B$ points chosen in the second stage must originate from $p(\mathbf{x})$ and are solely exploitative.
-By nature, they will be clustered around the predicted contour.
+By nature, they will be clustered around the predicted contour, where pass/fail classifications are hardest to pin down.
 We contend that having a mix of data from these two realms is superior to either extreme.
-Some of the works we referenced in Section 1 have investigated stopping criteria for surrogate training in reliability settings.
-The most similar to our proposed criterion is that of Zhu and Du, (2016) who recommend a convergence criterion akin to $\frac{\hat{\sigma}_{\alpha}}{\hat{\alpha}}\leq\epsilon$, where $\epsilon$ is a user specified threshold.
+Some of the previous works referenced in Section 1 have investigated stopping criteria for surrogate training in reliability settings.
+The most similar to our proposed criterion is that of Zhu and Du, (2016) who recommend a convergence criterion akin to $\frac{\hat{\sigma}_{\alpha}}{\hat{\alpha}}\leq\epsilon$, where $\epsilon$ is a user-specified threshold.
 While similar to our stopping criterion, the choice of $\epsilon$ here is make-or-break, and little guidance is provided.
@@
 7) have been entertained in other settings.
-Li and Xiu, (2010) use polynomial chaos models for a hybrid MC estimator with samples delegated to $\mathcal{U}$ based on predicted proximity to the contour, i.e., $|\mu_{n}(\mathbf{x}_{i})-t|$.
-This work, however, does not consider surrogate design and does not leverage surrogate UQ.
+Li and Xiu, (2010) use polynomial chaos models for a hybrid MC estimator with samples allocated to $\mathcal{U}$ based on predicted proximity to the contour, i.e., $|\mu_{n}(\mathbf{x}_{i})-t|$.
+Their work, however, does not consider surrogate design and does not leverage surrogate UQ.
 Fuhrländer and Schöps, (2020) use hybrid MC, but they only consider training the surrogate on samples from $p(\mathbf{x})$.
@@
 But a GP surrogate would benefit from some “anchor points” in these outer regions (Gramacy and Apley,, 2015).
-This distinction will be less prominent if $p(\mathbf{x})$ is uniformly distributed over $\mathcal{X}$, but the trade-off between exploring in the training phase and exploiting in the estimation phase will still be impactful.
-Even with uniform $p(\mathbf{x})$, training a surrogate only on maximal entropy points is suboptimal (Cole et al.,, 2023; Booth et al.,, 2024b).
+This distinction will be less prominent if $p(\mathbf{x})$ is uniformly distributed over $\mathcal{X}$, but the trade-off between exploring in Stage 1 and exploiting in Stage 2 will still be impactful.
+Even with uniform $p(\mathbf{x})$, training a surrogate only on maximal entropy points is suboptimal (Cole et al.,, 2023; Booth et al.,, 2024c).
 4 Synthetic experiments
 In this section we benchmark our method against state-of-the-art competitors on a variety of synthetic experiments.
-We consider four functions: the 2d Herbie function (Lee et al.,, 2011) with a total budget of 150, the 3d Ishigami function (Ishigami and Homma,, 1990) with a total budget of 300, the 6d Hartmann function (Picheny et al.,, 2013) with a total budget of 600, and the 4d Plateau function (Booth et al.,, 2024b) with a total budget of 200.
+We consider four functions: the 2d Herbie function (Lee et al.,, 2011) with a total budget of 150, the 3d Ishigami function (Ishigami and Homma,, 1990) with a total budget of 300, the 6d Hartmann function (Picheny et al.,, 2013) with a total budget of 600, and the 4d Plateau function (Booth et al.,, 2024c) with a total budget of 200.
 The first three of these are plausibly stationary, so we use a standard GP surrogate, fit with the Scikit-learn package in python (Pedregosa et al.,, 2011), and the entropy contour location scheme of Cole et al., (2023).
 The Plateau function is known to be nonstationary; it is characterized by flat regions with a steep sloping drop between them.
-To address the nonstationarity, we use a Bayesian DGP surrogate (Sauer et al.,, 2023b), fit with the deepgp package in R (Booth,, 2023), with the contour location scheme of Booth et al., (2024b).
+To address the nonstationarity, we use a Bayesian DGP surrogate (Sauer et al.,, 2023b), fit with the deepgp package in R (Booth,, 2024), and the contour location scheme of Booth et al., (2024c).
 With the right surrogate, the failure contour of the Plateau function can be well estimated from a rather small amount of data.
@@
 Hartmann $1.001\times 10^{-5}$ 6 100 600 $1.0\times 10^{8}$ GP Cole et al., (2023)
-Plateau $4.308\times 10^{-4}$ 4 30 200 $3.5\times 10^{6}$ DGP Booth et al., (2024b)
+Plateau $4.308\times 10^{-4}$ 4 30 200 $3.5\times 10^{6}$ DGP Booth et al., (2024c)
 Table 1: Simulation settings. True $\alpha$ values are estimated from 10 billion samples.
@@
 Figure 5: Distribution of the chosen $n$ across 30 repetitions.
-First, we deploy our stopping criterion (Section 3.1) for each of the four scenarios, starting with an initial LHS sample of size $n_{0}$ (Table 1) and repeating for 30 re-randomized Monte Carlo repetitions.
+First, we deploy our stopping criterion (Section 3.1) for each of the four scenarios (Table 1).
+Starting with an initial LHS sample of size $n_{0}$, we proceed with Stage 1, conducting CL with the specified method and determining the $n$ at which to stop.
+We conduct 30 re-randomized Monte Carlo repetitions.
 To avoid premature termination, we require a minimum of 10 observed failures and $n\geq 2n_{0}$ (at least as many contour locating acquisitions as original space-filling observations).
-Figure 5 shows the distribution of chosen $n$ values.
-Notice, all trials terminated before the allotted $N$, reserving some evaluations for additional estimation, but there was considerable variability across trials.
+Figure 5 shows the distribution of chosen $n$ values for each scenario.
+Notice, all trials terminated before the allotted $N$, reserving some evaluations for Stage 2, but there was considerable variability across trials.
 We view this variability as a strength – some random initializations need more data to effectively locate the contour while some get lucky from the start.
-Allowing the surrogate to dictate its own budget mitigates some of the risks of having an unlucky starting design.
-For each trained surrogate with chosen $n$ (which varies across repetitions), we collect a large Monte Carlo sample ($X_{M}$, randomized across repetitions), then deploy the following estimation methods:
-  • All CL: Ignoring the chosen $n$, continue contour location until the entire budget is exhausted and return the surrogate MC estimate (4).
-  • Hybrid Entropy: Our two-stage approach using the hybrid MC estimate (7) where $\mathcal{U}$ is chosen to contain the $B$ points from $X_{M}$ with highest entropy.
-  • Hybrid Proximity: Same as “Hybrid Entropy,” but $\mathcal{U}$ is chosen to contain the $B$ points from $X_{M}$ which are predicted to be closest to the failure threshold (i.e., the samples with the smallest $|\mu_{n}(\mathbf{x}_{i})-t|$).
+Allowing the surrogate to dictate its own budget allocation mitigates some of the risks of having an unlucky starting design.
+For each trained surrogate with chosen $n$ (which varies across repetitions), we collect a large Monte Carlo sample ($X_{M}$, constant across methods, randomized across repetitions), then compare the following estimation methods.
+All variations use the same Stage 1 contour location scheme up to point $n$; they differ only in their use of the remaining $B=N-n$ observations.
+  • Exhaustive CL: Ignoring the chosen $n$, continue contour location until the entire budget is exhausted and return the surrogate MC estimate (4).
+  • Two-stage Entropy: Our proposed two-stage approach (Figure 1) using the hybrid MC estimate (7) where $\mathcal{U}$ is chosen to contain the $B$ points from $X_{M}$ with highest entropy.
+  • Two-stage Proximity: Same as “Two-stage Entropy,” but $\mathcal{U}$ is chosen to contain the $B$ points from $X_{M}$ which are predicted to be closest to the failure threshold (i.e., the samples with the smallest $|\mu_{n}(\mathbf{x}_{i})-t|$).
 This is inspired by the work of Li and Xiu, (2010).
-  • SIIS: Surrogate-informed importance sampling (6) with a GMM bias distribution (Peherstorfer et al.,, 2016).
+  • SIIS: Surrogate-informed importance sampling (6), with $B$-many samples observed from an estimated GMM bias distribution (Peherstorfer et al.,, 2016).
 The bias distribution is trained on the predicted failures, i.e., $\left\{\mathbf{x}_{i}\in X_{M}\;|\;\mu_{n}(\mathbf{x}_{i})>t\right\}$.
@@
   • SIIS UCB: Same as “SIIS,” but the bias distribution is trained on the samples whose upper 95% confidence bound exceeds the threshold, i.e., $\left\{\mathbf{x}_{i}\in X_{M}\;|\;\mu_{n}(\mathbf{x}_{i})+1.645\sigma_{n}(\mathbf{x}_{i})>t\right\}$.
-Each method uses the same $X_{M}$ and the same surrogate trained up to point $n$.
-The distinction is in what is done with the remaining $B$ samples.
-As implemented, the latter three are to some degree an internal comparison since they use our proposed stopping criterion to determine the budget allocation between surrogate training and estimation (Figure 5).
+As implemented, the latter three are to some degree an internal comparison since they use our proposed stopping criterion to determine the budget allocation between surrogate training ($n$) and estimation ($B$).
 Without borrowing from our proposed approach, there would be no guidance on the allocation of $n$ and $B$ for these methods.
-The “Hybrid Proximity” approach is truly an internal comparison as it borrows everything from our proposed two-stage design aside from the specific allocation of samples to $\mathcal{U}$.
-It is included to show the impact of the entropy selection criterion, which uses surrogate UQ, over a selection solely based on the predicted mean.
+The “Two-stage Proximity” approach is truly an internal comparison as it borrows everything from our proposed two-stage design aside from the specific allocation of samples to $\mathcal{U}$.
+It is included to show the impact of the entropy selection criterion, which uses surrogate UQ, over a selection solely based on the posterior mean.
 Reproducible code for all experiments, including both the python and R implementations, is available in our public git repository.
@@
 Although their median performance was generally on target (indicative of their unbiasedness), these methods consistently offered the poorest performance.
-The more conservative “SIIS UCB” had no apparant advantage over the traditional “SIIS.” We believe importance sampling is simply ill-suited for our low data settings.
-In all scenarios aside from the simple two-dimensional Herbie function, our two-stage design (“Two-stage Entropy”) outperformed the exhaustive contour location approach (“All CL”).
-Particularly as dimension grows and the response surface becomes more complex, there is a clear advantage to reserving some simulator evaluations for the second stage.
-Most of the time there was agreement between the two criteria for allocating samples to $\mathcal{U}$ (“Entropy” and “Proximity”), but when there was disagreement the entropy criterion offered higher accuracy, with uniformly superior worst-case performance.
+The more conservative “SIIS UCB” had no apparent advantage over the traditional “SIIS.” We believe importance sampling is simply ill-suited for our low data settings.
+In all scenarios aside from the simplest two-dimensional Herbie function (where performance was comparable), our two-stage design (“Two-stage Entropy”) outperformed the exhaustive contour location approach (“Exhaustive CL”).
+Particularly as dimension grows and the response surface becomes more complex, there is a clear advantage to our two-stage design which transitions from CL to exploitation of MC samples.
+Within the two-stage designs, there was generally agreement between the two variations we considered for allocating samples to $\mathcal{U}$ (“Entropy” and “Proximity”), but when there was disagreement the entropy criterion offered higher accuracy, with uniformly superior worst-case performance.
 As suspected, the entropy criterion which incorporates uncertainty is more effective at identifying samples which the surrogate would misclassify.
@@
 Reynolds number $\sim\mathcal{N}\left(\mu=1\times 10^{7},\sigma=1\times 10^{6}\right)\;\;\textrm{truncated to}\;\;[5\times 10^{6},1.5\times 10^{7}]$
-Mach number $\sim\mathcal{N}\left(\mu=0.8,\sigma=0.02\right)\;\;\textrm{truncated to}\;\;[0.7,0.9]$
+Mach number $\sim\mathcal{N}\left(\mu=0.8,\sigma=0.02\right)\;\;\textrm{truncated to}\;\;[0.7,0.9].$
 We fix our total budget at $N=500$ simulator evaluations and use $M=2.5\times 10^{6}$ for the size of $X_{M}$.
-Booth et al., (2024b) found that DGP surrogates outperformed GP surrogates for contour location of this experiment.
+Booth et al., (2024c) found that DGP surrogates outperformed GP surrogates for contour location of this experiment.
 We follow the settings therein, starting with an LHS sample of size $n_{0}=100$ and fitting a Bayesian DGP surrogate with the deepgp package.
-We proceed with contour location following Booth et al., (2024b), stopping after every 10th acquisition to estimate $\hat{\alpha}$ and check whether it is with in the ±$\pm\hat{\sigma}_{\alpha}$ bounds.
-Once we halt contour location (Section 3.1), we use entropy to inform our hybrid MC estimator (Section 3.2).
+We proceed with contour location (Stage 1) following Booth et al., (2024c), stopping after every 10th acquisition to estimate $\hat{\alpha}$ and check whether it is within the ±$\pm\hat{\sigma}_{\alpha}$ bounds.
+Once we halt contour location (Section 3.1), we proceed to Stage 2, using entropy to inform our hybrid MC estimator (Section 3.2).
 Given the larger input dimension, we require at least 40 observed failures before we consider halting surrogate training.
@@
 Additionally, notice how the final hybrid MC estimate of the second exercise (right panel) brought the failure probability lower than what the surrogate had originally estimated, closer to the estimate of our first attempt (left panel).
-We take this as an indication that the hybrid estimator is effectively adjusting for surrogate inaccuracies.
+We take this as an indication that the hybrid estimator of Stage 2 is effectively adjusting for surrogate inaccuracies remaining after Stage 1.
 6 Discussion
-We presented a unique two-stage surrogate design for estimating failure probabilities of expensive computer experiments.
+We presented a unique two-stage surrogate design for estimating failure probabilities of expensive computer experiments, which outperforms both exhaustive contour location and surrogate-informed importance sampling.
 Starting with a contour locating sequential design, we halt contour location once surrogate training has plateaued, as indicated by practical convergence of a surrogate Monte Carlo estimator.
@@
 Our two-stage design is computationally efficient in the number of evaluations of the expensive simulator, but it requires many predictive evaluations from the surrogate.
-Posterior predictions must be computed at all $M$ samples multiple times.
 While surrogate predictions are very cheap, collecting billions of them is not necessarily trivial.
-Our experiments were limited to failure probabilities above $1\times 10^{-5}$ for this very reason.
-We could tackle smaller probabilities, but it would require larger $M$, which would eventually become impractical.
-Although we are focused on expensive simulators whose computational costs overshadow the cost of these surrogate predictions, our approach may simply not be suited to extremely rare failure probabilities.
+The computational cost of obtaining these predictions increases with $n/N$, $d$, and $M$, but it also depends on the surrogate and computing environment.
+For context, the most expensive predictions we needed in our exercises were those of our motivating airfoil simulator, with the fully-Bayesian DGP trained on all $N=500$ points in 7-dimensions.
+With parallelization over 10 CPUs, it took just over 10 minutes to obtain one million predictions.
+For smaller $\alpha$ requiring $M$ in the billions, the cost of these surrogate predictions may present an obstacle, but it is still far cheaper than comparative evaluations of the true expensive computer simulator.
+There may be avenues to circumvent surrogate predictions at all $X_{M}$ locations.
+We hypothesize that a classification rule based on nearest neighbors could work well: for a particular $\mathbf{x}_{i}$, if the surrogate predicts that all of its $k$-many nearest neighbors are “passes,” then we could presume $\mathbbm{1}_{\{\mu_{n}(\mathbf{x}_{i})>t\}}=0$ without needing to obtain $\mu_{n}(\mathbf{x}_{i})$ directly.
+This requires effective tuning of $k$, and would only speed-up computations if the cost of obtaining the nearest neighbor sets is less than that of obtaining the single surrogate prediction.
+For these reasons, we did not entertain such workarounds in our current work.
 In all of our examples, our stopping criterion terminated contour location before the complete budget ($N$) was exhausted.
 If a simulation budget is severely restricted, the budget could be exhausted before our proposed stopping criterion is met.
-While we prefer to have additional evaluations reserved for the estimation stage, we contend that using the entire budget for contour location to provide a surrogate MC estimate is the best option in this scenario.
+While we prefer to have additional evaluations reserved for Stage 2, we contend that using the entire budget for contour location in Stage 1 to inform a traditional surrogate MC estimate is the best option in this scenario.
 As we’ve shown, importance sampling based methods perform poorly when data is limited.
@@
 In fact, the noise-free nature of the simulator is what enabled us to frame our contribution as a two-stage design.
-Once $f(\mathbf{x}_{i})$ has been observed and incorporated into the surrogate, surrogate predictions should interpolate these observations, i.e., $\mu_{n}(\mathbf{x}_{i})=f(\mathbf{x}_{i})$.
+Once $f(\mathbf{x}_{i})$ has been observed and incorporated into the surrogate, surrogate predictions should interpolate these observations.
 Thus our hybrid MC estimator (7) is really a surrogate MC estimator (4), with a strategically trained surrogate.
-Different design choices could be warranted for stochastic simulators (Baker et al.,, 2022).
-In particular, the hybrid estimator of Eq.
-(7) should not make use of noisy $f(\mathbf{x})$.
-Replicates in the design may be crucial (Binois et al.,, 2019).
+Stochastic simulators (e.g., Baker et al.,, 2022) present a unique challenge as repeated observations of noisy $f(\mathbf{x})$ may result in a mix of passes and failures for the same $\mathbf{x}$.
+To accommodate this, we could have broadened the formulation of Eq.
+(1) by replacing the indicator function $\mathbbm{1}_{\{f(\mathbf{x})>t\}}$ with the unknown probability $\mathrm{P}(f(\mathbf{x})>t)$, which recovers the indicator when $f$ is deterministic.
+The key advantage of our hybrid estimator – that observing $f(\mathbf{x}_{i})$ ensures correct classification of $\mathbbm{1}_{\{f(\mathbf{x}_{i})>t\}}$ – is not applicable to noisy simulators.
+Instead, new active learning strategies that strategically allocate replicates (Binois et al.,, 2019) and inform a MC estimator leveraging surrogate posterior predictive probabilities are warranted.
+We leave this as an avenue for future work.
 When the input distribution is not uniform over $\mathcal{X}$, it could be beneficial to leverage information about $p(\mathbf{x})$ in the sequential training of the surrogate (Abdelmalek-Lomenech et al.,, 2024).
 There is arguably no need to train a surrogate in a region where inputs will never occur.
-Incorporating this information in the contour location stage could improve failure probability estimation downstream, but we have seen positive impacts from having “anchor points” in regions of very low density (such as the black open circles around the edges of Figure 3).
+Incorporating this information in the contour location stage could improve failure probability estimation downstream, but we have seen positive impacts from having “anchor points” in regions of very low density (such as those around the edges in Figure 2).
 We suspect there is a delicate balance that would work well here, but this avenue is currently underexplored.
-Estimation of failure probabilities is often an intermediate step within a larger objective, such as a stochastically constrained optimization.
+Furthermore, while we focused on stochastic inputs, it is possible to extend our approach to incorporate controllable/design variables by using point masses in $p(\mathbf{x})$.
+Finally, estimation of failure probabilities is often an intermediate step within a larger objective, such as a stochastically constrained optimization.
 For example, in the context of our airfoil experiment, we may wish to find the optimal wing configuration given that the probability of failure remains below a specified threshold.
-The ability to effectively estimate failure probabilites from limited data, which we have tackled here, is integral to the success of this larger enterprise.
+The ability to effectively estimate failure probabilities from limited data, which we have tackled here, is integral to the success of this larger enterprise.
+While we focused on allocating a predetermined budget of evaluations, it could be useful to develop a decision rule for halting Stage 2, potentially reserving some evaluations, particularly if this design is within a larger optimization loop.
 SUPPLEMENTARY MATERIAL
+Appendix A Test functions
 Here we provide the details of the functions, thresholds, and input distributions used in the benchmark exercises of Section 4. In each case, failures are defined as $f(\mathbf{x})>t$.
@@
 Plateau.
-The Plateau function (Booth et al.,, 2024b) is defined in arbitrary dimension over $[0,1]^{d}$ as
+The Plateau function (Booth et al.,, 2024c) is defined in arbitrary dimension over $[0,1]^{d}$ as
 $f(\mathbf{x})=2*\Phi\left[\sqrt{2}\left(-4-3\sum_{i=1}^{3}(4x_{i}-2)\right)\right]-1$
@@
 We define the input distribution as $x_{i}\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathcal{N}(0.6,0.11)$ for $i=1,\dots,4$, truncated to $[0,1]$.
+Appendix B Batch size comparison
+Here we consider variations of the batch sizes used within the second stage of our design.
+When Stage 1 is complete and $n$-many evaluations have been spent on contour location, the remaining budget ($B=N-n$) is to be spent on the highest entropy samples from $X_{M}$.
+Since the surrogate is used to select these samples, there is an option to select them in smaller batches, updating the surrogate with the new observations between each batch (indicated by the dotted arrow in Figure 1).
+But smaller batches require more computation as surrogate predictions must be obtained at all $X_{M}$ locations after every batch update.
+To investigate, we consider three different batch sizes for the Herbie function example of Section 4.
+First, we choose all high entropy points in a single batch of size $B$ (this is the method used throughout the paper, thus these results are duplicated from Figure 6).
+Second, we consider selecting one high entropy point at a time, updating the surrogate after every high entropy point is observed to better inform the selection of subsequent high entropy points.
+Third, we consider selecting high entropy points in batches of size 10.
+Note, changing the batch size used within Stage 2 has no effect on Stage 1; each of these batch variations starts in the same place after the completion of contour location and uses the same $X_{M}$.
+Figure 8 shows the resulting failure probability estimates across 30 Monte Carlo repetitions.
+There is no significant difference in performance among the three batch sizes.
+The grey shaded region marks $\{\alpha\pm 2\sigma_{\alpha}\}$; it indicates the level of variability that we would expect from any Monte Carlo estimator of specified $M$, even if every sample was classified accurately.
+All methods consistently fall within this region, indicating effective estimation (note the narrower y-axis range compared to Figure 6).
+Given the computational expense of repeatedly updating the surrogate, we embrace the single batch approach.
+Figure 8: Estimated failure probabilities shown over the true failure probability (dashed red) ±$\pm 2$ standard errors (grey shading) for proposed two-stage designs of the Herbie function.
+Variations use different batch sizes for selecting the highest entropy samples within Stage 2.